### v1

# The Interaction Layer: An Exploration for Co-Designing User-LLM Interactions in Parental Wellbeing Support Systems

Abstract.
Parenting brings emotional and physical challenges, from balancing work, childcare, and finances to coping with exhaustion and limited personal time. Yet, one in three parents never seek support. AI systems potentially offer stigma-free, accessible, and affordable solutions. Yet, user adoption often fails due to issues with explainability and reliability. To see if these issues could be solved using a co-design approach, we developed and tested NurtureBot, a wellbeing support assistant for new parents. 32 parents co-designed the system through Asynchronous Remote Communities method, identifying the key challenge as achieving a “successful chat.” As part of co-design, parents role-played as NurturBot, rewriting its dialogues to improve user understanding, control, and outcomes. The refined prototype evaluated by 32 initial and 46 new parents, showed improved user experience and usability, with final CUQ score of 91.3/100, demonstrating successful interaction patterns. Our process revealed useful interaction design lessons for effective AI parenting support.
Figure 1. Overview of the six-part study detailing the iterative development and testing of NurtureBot, a parental wellbeing assistant.
[…]
A visual representation of the six-part iterative development of NurtureBot involving two parent cohorts. The chart shows the progression from the initial prototype to NurtureBot v1, followed by ARC co-design sessions, NurtureBot v2, and NurtureBot v3. Key findings include conversation starters, tone, parenting resources, and interaction improvements. NurtureBot v1 identified seven major problems, leading to an interaction layer for NurtureBot v2, which addressed four feature requests. NurtureBot v3 showed further improvements with five new feature requests and achieved the highest CUQ score of 91.3.
1. Introduction
Parenting, often hailed as the most important job in our world, exacts a significant toll on both mind and body. The perinatal period represents a critical developmental stage for the child, the parent(s), and the formation of attachment relationships (Li et al., 2017; Allan et al., 2013). This period, spanning from the onset of pregnancy to the first year postpartum, is marked by profound physical, emotional, and psychological changes, which significantly impact parental wellbeing. In particular, many expectant and new mothers encounter substantial challenges, in mental health, emotional adjustment, and caregiving (Widarsson et al., 2012), in addition to the transformation their body goes through. During this period, parents experience heightened stress, anxiety, and feelings of isolation, increasing the need for support in these areas. Yet, we know that one in three parents do not seek help (UNICEF UK, 2022). The literature on perinatal care identifies several key issues, including inadequate access to mental health resources, availability of healthcare professionals, cost to the government, stigma surrounding parental mental health, and challenges in identifying and supporting at-risk mothers (Rallis et al., 2014; Davis et al., 2022; Howard and Khalifeh, 2020). These challenges have been further aggravated by the COVID-19 pandemic (Bridle et al., 2022; Kinser et al., 2022). As a result, we see a growing body of research exploring how digital technologies can enhance perinatal care and provide support during this critical period (van den Heuvel et al., 2018; Ginja et al., 2018).
Notable Human-Computer Interaction (HCI) research has laid the groundwork for exploring how digital interventions increase wellbeing across aspects such as physical and mental health of a family (Pina et al., 2017), a child’s growth (Kientz et al., 2009), sleep tracking of parents and children (Pina et al., 2020), and managing family screen-time behaviours (Hiniker et al., 2016). Today, with the availability of Artificial Intelligence (AI) to everyday users, use of general generative applications such a ChatGPT for parenting is on the rise (Quan et al., 2024). With responsible design, Large Language Models (LLMs) have the potential to facilitate a range of services, from offering self-guided mental health support (Sharma et al., 2024), real-time emotional support (Kang et al., 2024) to providing evidence-based information at scale (Lai et al., 2023). The current use of LLMs to support parental and child wellbeing is reviewed with high potential and ethical cautions (Ashraf, 2024; Stade et al., 2024). Their application within this context is fraught with limitations and uncertainties, especially of accuracy, empathy, and usefulness (Kaneda et al., 2023).
Given these challenges, our research adopts an exploratory approach, probing into a small subset of these broader issues. We focus on understanding how interactions between LLMs and parents might be designed to be perceived as helpful and engaging. Our goal is to explore what constitutes a usable and useful interaction in these contexts and how LLMs can be integrated into wellbeing support in ways that are practically engaging and thus useful for parents. To this end, we conducted a six-part online study involving 78 parents, who trialled iterative improvements of NurtureBot, a technology probe, designed to support parental wellbeing. This study, while not aimed at delivering a finalised solution, served as an exploratory step toward understanding how parents interact with LLM-based agents and identifying necessary features to make these tools genuinely supportive. The following Research Questions (RQs) guide our investigation into the user interaction aspects of engaging with an LLM-based parental wellbeing assistant:
  • RQ1: What are the main problems parents face when interacting with an LLM-based parental wellbeing assistant in support-seeking scenarios?
  • RQ2: What key user needs must a parental wellbeing assistant address? How do parents describe their ideal vision of how the assistant should function, particularly regarding their understanding of its workings, control over it, and opportunities to improve the outcomes?
  • RQ3: How can we employ co-design methods to craft interactions around AI, enhancing usability and the overall user experience?
Study Flow Overview
This six-part user-centred design process, as illustrated in Figure 1, guided the iterative development of NurtureBot. To mitigate emotional risks, we engaged a ‘slightly-out-of-target’ group of parents, each with a child aged five or six. These parents simulated stages of their perinatal period and reflected on past challenges with NurtureBot. Using a mixed-methods approach, we collected both qualitative and quantitative feedback, with NurtureBot v1, v2, and v3 tested over a five-day period.On the final day, participants completed the Chatbot Usability Questionnaire (CUQ) (Holmes et al., 2019), assessing usability and engagement. (see Figure2).
  • Part I: The Initial Prototype: NurtureBot v1 The development of NurtureBot v1 leveraged LLM technology, informed by a community wellbeing initiative supporting new parents. Zero-shot prompting was used to integrate supportive conversation starters and key features, including empathetic chatting, wellbeing exercises, and parenting information (see Section 3).
  • Part II: NurtureBot v1 Testing and Analysis A cohort of 32 parents engaged with NurtureBot v1, generating 144 conversations. Thematic analysis identified several issues, including lack of conversational flow, robotic tone, transactional attitude, limited personalisation, and localisation. Despite these challenges, NurtureBot v1 achieved a CUQ score of 85.4, surpassing the benchmark of 68 (Holmes et al., 2019), indicating good usability (see Section 4).
  • Part III: ARC Co-Design Sessions and Analysis The Asynchronous Remote Community (ARC) (MacLeod et al., 2016; Prabhakar et al., 2017) co-design sessions engaged the same cohort of parents with Miro board activities (Miro, 2011).
    • Parents first prioritised key issues, with “successful chat” emerging as the top priority.
    • Then they reimagined and rewrote NurtureBot’s dialogues, aiming for better user understanding, user control, and improved outcomes, a structure akin to a cybernetic loop (Patten and Odum, 1981).
    • Thematic analysis of 189 revised dialogues led to requests for essential features such as metaphorical explanations, empathetic engagement, feature introductions, personalising exchanges, and step-by-step guidance (see Section 5).
  • Part IV: The Interaction Layer: NurtureBot v2 NurtureBot v2 introduced an interaction layer with three essential components (See Figure 7): (i) A core interaction principle that enabled users to ”Understand, Control, and Improve” exchanges, (ii) Interaction levels based on few-shot prompting, incorporating co-designed dialogues from ARC sessions, and (iii) Interaction elements, tailored to address user needs identified during ARC.
  • Part V: NurtureBot v2 Testing and Analysis NurtureBot v2 underwent testing with the same initial cohort, yielding 104 conversations. Our analysis revealed that two minor issues persisted: repetitive conversational hints and the delivery of parenting information in long, overwhelming paragraphs. Additionally, themes such as localisation and conversational memory, which we did not prioritise addressing in this study, continued to emerge. NurtureBot v2 achieved a CUQ score of 87.4, indicating further improvement, though the potential bias from participants’ involvement in co-design was considered (see Section 7).
  • Part VI: Prototype Improvement: NurtureBot v3 and Validation To ensure an unbiased evaluation, we refined NurtureBot further and tested v3 with a new cohort of 46 parents. Feedback analysis identified two new problems, the need for more detailed responses at times, difficulty following step-by-step exercises via text, and repeated requests for additional features such as videos and memory retention for personalised interactions. NurtureBot v3 received a CUQ score of 91.3, marking a robust improvement in usability (see Section 8).
Our contributions are:
  • Empirical Findings and Implications for co-designing ‘usable and useful’ LLM-based parental wellbeing support systems, from a longitudinal four-week iterative study, gathered from 78 parents, eliciting the interactional challenges they encountered and the solutions proposed to address these issues.
  • The Interaction Layer a prompt architecture that affords in-situ user understanding, user control, and conversational improvements by directly incorporating end-user inputs as prompt instructions usable by LLMs, addressing usability challenges to enhance user experience with AI-driven support tools for parenting and beyond.
This study initiates our exploration into co-designing agents for parental wellbeing, setting the foundation for future research to refine and expand AI tools. The diverse parenting experiences and emotional complexities highlight the need for continued development and testing.
2. Related Work
2.1. Supporting Family Wellbeing with Digital Tools
Within the field of HCI, there has been a long-term research agenda focusing on the study and development of impactful digital tools that offer a variety of social, physical, and mental wellbeing support to mothers (D’Ignazio et al., 2016; Thomas et al., 2019; Britton et al., 2019). The design and evaluation of digital tools aimed at promoting maternal wellbeing during the transition to motherhood (Newhouse, 2019), bump2bump, emphasise the need for holistic approaches in HCI that consider the challenges faced by new mothers. The role of peer support in digital environments is highlighted by the finding that online networks can provide valuable emotional and practical assistance to parents, particularly in situations where traditional face-to-face support may not be available (Yamashita et al., 2022). The use of the Baby Buddy app among first-time mothers in the UK (Deave et al., 2019), shows that while the app is generally well-received, its effectiveness is influenced by how well it aligns with the mothers’ understanding of their needs and the app’s usability. This points to the necessity of involving parents in the design process to ensure that digital tools meet their expectations and provide meaningful support.
The integration of AI chatbots into parenting and family well-being has opened new avenues for providing support and guidance to parents (Yu et al., 2023; Entenberg et al., 2021; Wong et al., 2021). Parenting with conversational agents such as Alexa (Beneteau et al., 2020), Siri (Sezgin et al., 2022) and other voice-activated agents in the home (Sun et al., 2021) have been critically studied in the past few years. Following the advent of AI chatbots, the integration of LLMs into parenting practices (Quan et al., 2024), is ushering in an era of unprecedented possibilities and challenges. LLM-based agents such as ChatGPT are being adopted by parents of young children (Quan et al., 2024; Kaneda et al., 2023), offering new learning opportunities and personalised support, though their impact on parenting remains largely unexplored, marking this as an era of the unknown. The performance and acceptance of ChatGPT’s responses in the childcare field (Kaneda et al., 2023), is also being studied with caution. Despite technological advancements, several limitations persists such as misinformation (Kaneda et al., 2023), lack of diversity (Bajwa et al., 2024), threat of data misuse (Neerincx et al., 2021), and lack of user adoption (Baumel et al., 2019). While gamification can be an effective motivational tool (Wernbacher et al., 2022), it often does not appeal to all users equally, particularly those unfamiliar with or resistant to digital game mechanics. The need for humane interaction with a good user experience has been repeatedly reported in the literature for digital health tools built with or without AI (Bautista and Schueller, 2023). A review of 13 AI-based mental health apps with user ratings over 4 out of 5, revealed limited support to AI literacy and explainability (Alotaibi and Sas, 2023). Real-world engagement data from 93 health apps indicate that while install rates are high, sustained user engagement remains low, with the median 15- and 30-day retention a mere 3.9% and 3.3% respectively (Baumel et al., 2019). These gaps suggest a need for further research and development of ‘truly interactive’ applications to deliver sound AI-powered digital tools for wellbeing.
2.2. Interacting with AI for Everyday Users
As AI permeates daily life, from recommendation systems to digital assistants, the ability of non-expert users to engage meaningfully with AI systems becomes critical (Eiband et al., 2019). Research highlights the significance of Explainable AI (XAI) in demystifying the opaque processes of machine learning models, thereby enhancing user understanding, an extensively studied dimension of Human-Centred AI (HCAI)(Saeed and Omlin, 2023; Capel and Brereton, 2023). XAI methods such as Local Interpretable Model-agnostic Explanations (LIME) and SHapley Additive exPlanations (SHAP) have been pivotal in rendering AI decision-making processes more transparent (Ribeiro et al., 2016; Lundberg and Lee, 2017; Adadi and Berrada, 2018; Arrieta et al., 2020). While XAI tools enable expert users to grasp the rationale behind AI outputs, (Gunning et al., 2019; Miller, 2023a; Carvalho et al., 2019), most non-experts do not engage with XAI (Miller, 2023b), highlighting the need for innovative methods to foster AI understanding (Miller, 2023b). Control over AI systems by everyday users is another essential dimension of HCAI, with research underscoring the importance of user agency in human-AI interactions and challenges with it (König, 2024). Interactive interfaces that allow users to refine AI behaviour, such as setting parameters, have been shown to significantly enhance user satisfaction and the alignment of AI outputs with user expectations (Jain et al., 2023; Amershi et al., 2014; Holstein et al., 2019; Viswanathan et al., 2022), maintaining user engagement, and ensuring that AI systems serve the intended purposes (Ueno et al., 2022; Cai et al., 2019). Improvement of AI systems through user feedback represents the third crucial dimension of HCAI. Human-in-the-loop approaches, where users contribute to the iterative development and refinement of AI models, have gained traction as a means of optimising AI (Dudley and Kristensson, 2018; Amershi et al., 2014; Fails and Olsen, 2003). Involving end-users in the process, is not only about improving the AI but also about ensuring that the systems evolve in ways that are aligned with user needs and ethical considerations (Luchs, 2023). In summary, HCAI efforts (Capel and Brereton, 2023) can be categorised into three fundamental objectives: understanding, controlling, and improving AI systems. Borrowing from the cybernetic loop (Patten and Odum, 1981), akin to sensing, processing, and reacting; these categories form the basis for a continuous cycle of interaction between humans and algorithms. We plan to employ these three lenses as cues for our co-design of user-LLM interactions, ensuring that the development of AI systems remains aligned with the needs and expectations of everyday users (Wiener, [n. d.]; Bao et al., 2021).
The design of CUIs (Landay et al., 2019; Diederich et al., 2022; Clark et al., 2019) requires a deep understanding of HCI principles, ensuring that these systems can handle the nuances of human dialogue, such as turn-taking, repair mechanisms, and emotional engagement (Landay et al., 2019; Cruz-Sandoval and Favela, 2019; Kaye et al., 2018). As the field progresses, there is a growing emphasis on addressing these challenges to enhance the user experience and make conversational interactions more natural and effective (Lee et al., 2019; Hwang et al., 2019). The emergence of large language models (LLMs) has revolutionised user interactions across various domains, providing unprecedented capabilities in understanding and generating human language. In healthcare, LLMs have been integrated into administrative and clinical tasks, streamlining operations such as patient record management, appointment scheduling, and even aiding in medical diagnostics, thereby enhancing efficiency and reducing the burden on healthcare professionals (Gebreab et al., 2024; Dao et al., 2024; Brown et al., 2020). One of the most promising applications of LLMs is in the realm of personal wellbeing, where they are employed in the development of conversational agents designed to support mental health, offering personalised interactions that help users manage stress, anxiety, and other mental health conditions (Ferrara, 2022; Ke et al., 2024). Despite the promising advancements, issues such as hallucinations (Xu et al., 2024), task-specific limitations (Mizrahi et al., 2024), and user experience problems persist (Wang et al., 2024). Innovative human-centred solutions, are needed to build reliable, interactive and creative agents tailored to unique needs (Shankar et al., 2024). In the context of parenting + AI, this calls for a closer examination of how AI can offer useful and usable support while addressing the unique challenges in this domain.
3. Part I: The Initial Prototype: NurtureBot v1
3.1. Origins and motivations for the initial prototype
Our parental wellbeing support prototype, was based on a community wellbeing project called Nurture, initially run by a nonprofit organisation dedicated to children’s wellbeing [Org name removed for review]. The original Nurture project was a peer-to-peer mentoring initiative designed to support new mothers during their baby’s first year. In this program, experienced volunteer mothers were trained to mentor new moms, providing guidance and encouragement. A digital platform supporting this program offered personalised, week-by-week content delivered via text messages, acting as a conversation catalyst between mentors and mentees. The prompts were designed to initiate meaningful dialogue and help new moms navigate early motherhood.
The Nurture project was conducted over a span of three years, involving approximately 60 mentors and 603 parents who participated in one-on-one texting. While the project was successful in providing personalised support and the fostering relationships through these interactions, this model faced scalability challenges. While effective, Nurture project was resource intensive and difficult to expand to a broader audience without significant investment in human resources.
3.2. Designing the initial prototype
Recognising the need for a scalable solution that could retain the core support offered by Nurture, we partnered with the organisation to collate resources and findings from Nurture project for developing a new AI tool. In collaboration with the organisation’s project team, we wrote the initial prompt for a basic large language model (LLM)-based chatbot using Open AI’s GPT-4 API (OpenAI, 2023), zero-shot prompting, which we named NurtureBot v1.
NurtureBot v1 was designed to replicate the supportive environment of the original peer-to-peer texting interactions while introducing the scalability and accessibility of AI. The chatbot was refined to deliver three key features: empathetic chatting, wellbeing exercises, and parenting information, all adapted for AI-driven interactions. Additionally, it incorporated ”conversation starters” thoughtfully crafted by Nurture mentors. Empathetic chatting was intended to provide an active listening that reflect warmth and understanding. Wellbeing exercises were included to offer practical tools for managing stress and mental wellbeing, similar to those shared by mentors. Lastly, parenting information was integrated to ensure that users could access reliable and evidence-based resources tailored to their needs, much like the information shared in the original peer support sessions.
Leveraging insights from the Nurture project in Part I, we designed NurtureBot v1 and initiated iterative trials in the following Part II by testing the prototype with everyday users.
4. Part II: NurtureBot v1 Testing and Analysis
In this phase, we address RQ1: “What are the main problems parents face when interacting with an LLM-based parental wellbeing assistant in support-seeking scenarios?”.
4.1. Participants and Method
Once we had secured ethical approval via our institutional ethics board [removed for review], we recruited 32 parents of children aged 5 or 6 through the Prolific platform (Prolific, 2014). The average age of participants was 41.17 years, with 20 females and 12 males. Five participants (15%) were in the habit of using ChatGPT everyday. Participants self-reported no acute stress or mental or physical health issues during the study period, nor to have they received therapy in the 12 months prior. The study’s purpose and methodology were explained to parents through an information sheet, after which informed consent was obtained. They were paid  \$3 for each day of participation.
Over the five-day period, participants engaged with NurtureBot v1 for at least five minutes each day, with each day dedicated to simulating a specific stage of early parenthood: pregnancy, the week of childbirth, few weeks postpartum, one month, and six months after birth. The objective was to assess how well the NurtureBot v1 could support them in these hypothetical scenarios. Each day, participants interacted with the chatbot, receiving responses that included links to resources, empathetic messages, and suggestions for wellbeing exercises (see Figure 3). At the end of each day, after interacting with NurtureBot hosted on a website via Streamlit (Streamlit, 2019), participants were asked to complete a survey via Qualtrics (Qualtrics, 2005), providing feedback on their experience with the chatbot by answering the questions listed in Figure 2. On the final day, they also completed the Chatbot Usability Questionnaire (CUQ), a specialised questionnaire designed to access the usability of healthcare chatbots (Holmes et al., 2019).
Figure 2. Survey Questions for Each Day in NurtureBot v1, v2, and v3 technology trials
[…]
xxx
4.2. Data Collection and Analysis
Over the course of the five-day study, 144 conversations and associated feedback were generated, denoting a daily-average of 90% (28.8 out of 32 participants). All data was securely stored on departmental servers, compliant with GDPR (Union, 2016), and no personal data was collected, consistent with Prolific platform’s design.
The qualitative analysis of the feedback from the survey (See questions 1 and 2 in Figure 2) was conducted using a thematic analysis approach (Braun and Clarke, 2012). Our analysis focused on the instances that captured parent reactions to interacting with NurtureBot. The first author conducted the initial analysis using inductive, open coding, and regularly met with the second author to discuss emerging themes and explore alternative interpretations of participant responses. While participants did share positive feedback on several aspects of the chatbot, highlighting its usefulness and relevance, especially during early stages of parenthood, we do not explore the parents perspectives on the potential of LLM-based parental wellbeing support as this is reported in the literature (Ashraf, 2024), and to keep our focus on answering RQ1. Hence, we focused on analysing 131 of the 288 (45.48%) pieces of feedback that highlighted negative interactions and suggestions for improvement. This led us to the identification of 28 unique pain points, which were further grouped into seven overarching themes, converged in three subsequent meetings. These themes highlighted the key problems participants faced while interacting with NurtureBot v1 as exemplified below. We provide only a high-level summary here to conserve the reader’s focus for the ARC section, which forms the core of the HCI discussion.
Figure 3. Example conversation with NurtureBot v1
[…]
xxx
4.3. Findings of initial trial: Parent reactions to Nurture Bot v1
Problem 1: Unable to Chat: Participants expressed frustration with NurtureBot’s tendency to cut off their conversation. They were disappointed when the chatbot redirected them to external URLs rather than providing concise information within the chat itself. Even when information or exercises were provided within chat, NurtureBot did not engage further, making it difficult to continue the conversation, thereby rendering the participants unable to experience empathetic chatting, which was a key feature. They found this approach disruptive and felt it hindered the conversational flow. Participants noted:
”The only thing I would maybe find better is instead of providing links to articles, maybe include the information from the articles into the messages from NurtureBot. That way you haven’t got to click off from NurtureBot and could spend longer speaking.”
”Today I was trying to see if NurtureBot would focus more on chatting and less on giving infomration.”
Problem 2: Patronising and Robotic Response Tone: Participants reported that the NurtureBot’s responses felt patronising and robotic, often pretending to understand human emotions or pain. For example, a participant stated:
”What’s the point of platitudes like ‘recovering from a c section can be painful’ - what does an AI know?? Why would it be helpful to hear that expressed?”
Another participant commented:
”I think maybe make it slightly more warm and more human-like rather than ‘here’s your information.’”
Problem 3: Lack of Localised Resources: The resources suggested by the chatbot were not tailored to participants’ specific country or area, making the information less relevant and useful. One participant highlighted:
”When asking for groups in my area, it needs to be more specific, especially for the area I live in as it came up with places that were a bit irrelevant.”
Problem 4: Memory and Continuity in Conversations: Participants wanted NurtureBot to remember details about them, such as their name, their baby’s name, and their troubles. The lack of continuity made the interactions feel impersonal and disconnected. As one participant mentioned:
”I think it works very well the way it is. Obviously I don’t know fully how it works but it would be very useful if it remembers everything that you ever tell it so it builds up a profile of the child and parents so you don’t have to keep repeating information.”
Problem 5: Generic responses: Participants felt that NurtureBot’s responses lacked specificity and depth, often resorting to external links without giving user a chance to discuss their situational need and providing direct, tailored answers based on reliable sources. This approach was seen as redundant, as users felt they could achieve similar results using standard search engines. One participant expressed:
”Pointless. I wanted something specific, not links. Don’t understand why I wouldn’t just use Google.”
Problem 6: Transactional Nature of Conversations: The chatbot was perceived as overly transactional, asking questions in rapid succession without allowing participants enough time to process the information or engage in a more natural, flowing conversation. One participant described this experience:
”It was OK. It still feels quite transactional - NurtureBot wants me to ask questions then I get the answer, and then it asks me the next question. There’s no sense of pausing to process what’s been said, and asking the next question immediately feels a little bit like McDonald’s drive through when you’re asked after every item ‘is that all?’. It would be nice to feel like the conversation is a bit less like ping pong.”
Problem 7: Content Overload: Participants found the amount of information provided in NurtureBot’s responses overwhelming, particularly when the chatbot delivered long paragraphs filled with multiple links. This made it difficult to process the information and maintain a meaningful connection with the conversation. As one participant remarked:
”Ok, but it felt like the long responses with lots of links were a bit overwhelming and lacking in connection.”
4.3.1. Quantitative results
Our subjective questions regarding NurtureBot v1’s task satisfaction received an average score of 4.29 out of 5, while interactive questions averaged 4.36 for understandability, 3.97 for controllability, and 3.88 for improvability. The standard questionnaire analysis resulted in a mean CUQ score of 85.4 out of 100, which, when compared to the CUQ benchmark of 68 (Holmes et al., 2019) in a one-sample t-test, was extremely statistically significant (t(29) = 7.11, p ¡ 0.0001, mean difference = 17.40, 95% CI [12.40, 22.40], SE = 2.45). These scores and their implications are discussed further in Section 8.3, alongside comparisons with NurtureBot v2 and v3.
After identifying seven key problems with parent-NurtureBot v1 interactions in Part II, our goal in Part III was to co-design by tasking users to prioritise problems and imagine ideal solutions.
5. Part III: ARC Co-Design Study
To begin addressing RQ2 – “What key user needs must a parental wellbeing assistant address? How do parents describe their ideal vision of how the assistant should function, particularly regarding their understanding of its workings, control over it, and opportunities to improve the outcomes?,”– the third phase of the study involved Asynchronous Remote Communities (ARC) co-design sessions (MacLeod et al., 2016), designed to directly engage participants in the iterative design process. These sessions were facilitated using the Miro platform (Miro, 2011), allowing the same 32 participants who tested NurtureBot v1 (see Subsection 4.1) to return to collaboratively co-design solutions asynchronously, thus accommodating different schedules, total anonymity, and fostering broad participation.
The ARC sessions involve two core activities over three sessions (1 week), prioritising, and addressing key issue(s) with NurtureBot v1. The average participation was 29.33 out of 32 (91.63%) participants per day. To help familiarise participants with using Miro, we designed a training activity and instructional tips for using Miro on mobile devices, as well as providing direct messaging support via Prolific.
5.1. ARC Activity One: Diagnosing and Prioritising Pain Points
The goal of the first activity was to establish which of the expressed difficulties were the biggest priority to address. Using the seven key problems from NurtureBot v1’s technology trial (see Section 4.3), we asked participants to prioritise these issues. By disclosing these key issues, the goal was to ensure that all participants could see the broader scope of the negative feedback gathered from others within their group.
5.1.1. ARC Activity One: Method
We crafted a ”problem statement” for each issue, synthesising the findings without directly replicating any participant’s verbatim feedback (listed as problem statements in Figure 4). This method not only helped verify our interpretation and synthesis of the feedback but also prevented potential bias, ensuring participants did not favour their own comments when voting. To provide sufficient scaffolding for reflection, we organised the activity as follows:
Reviewing Identified Problems: Participants were guided through a review of the seven major problems. As they reviewed these issues, they were asked to select one of three stickers—“This applies to me,” “This may apply to me,” or “This does not apply to me”—and place it on each of the seven problems. This step was important for setting a common understanding of the challenges identified from the group.
Ranking and Prioritisation: After reviewing the problems, as the next step, participants were asked to rank them according to their personal experiences, using stickers (gold, silver, and bronze icons provided on the Miro board) (See Figure 4). This step was designed to capture participant perspectives on the perceived severity of each problem, which would inform the prioritisation of design efforts in subsequent activities.
Figure 4. Results from ARC Activity One voting by 30 participants, where key problems in interacting with NurtureBot were ranked using an Olympic points system. Complete Miro board including the other set-up phases participants went through before voting can be found in the supplementary materials.
[…]
xxx
5.1.2. ARC Activity One: Results
Based on participants responses, the following options were voted as the top three problems by participants:
  • Problem 1 - I wanted to have a chat but NurtureBot kept cutting me off by redirecting me or giving me paragraphs of information. emerged as the top challenge with NurtureBot.
  • Problem 3 - The suggested resources were not localised to my country or area was the second priority.
  • Problem 4 - I want NurtureBot to remember things! E.g., who I am, my baby’s name and our previous conversation and context was ranked third.
We next took the top-ranked problem (Problem 1) to begin considering how we might enhance “successful chat problem” in Activity 2.
5.2. ARC Activity Two: Co-Designing Solutions
The second activity focused on solution generation for problem 1. Participants were asked to step into the role of NurtureBot: “Imagine You Are NurtureBot and Fill in the Blanks.”
5.2.1. ARC Activity Two: Method
The activity was structured as follows:
Scenario-Based Problem Solving: As the first step, participants were presented with three specific scenarios corresponding to the three main features of NurtureBot (i) interacting with NurtureBot for the first time and wanting to have an empathetic chat, (ii) seeking wellbeing exercises to manage work stress, and (iii) seeking information to soothe their baby; that exemplified trying to have a successful chat aiming to use all three features of NurtureBot. For each scenario, participants reviewed an example response and were asked to consider how NurtureBot could help users to better understand its offerings, control the conversation, and improve its outcomes.
Collaborative Dialogue Writing: Next, participants were invited to take on the role of NurtureBot and rewrite it’s dialogue for each scenario. We offered cues that encouraged participants to work on ”Understanding, Controlling, and Improving” NurtureBot, similar to the idea of a cybernetic loop (Patten and Odum, 1981) (see Section 2.2). This part of the activity encouraged participants to think critically about how the chatbot could better meet their needs in specific contexts, and which would inform NurtureBot v2 (See Figure LABEL:UCI).
Iterative Feedback and Voting: As the last step, after proposing solutions, participants reviewed and built upon each other’s suggestions. We invited participants to vote on the best ideas (See heart stickers in Figure LABEL:UCI), ensuring that the most popular and potentially effective solutions were prioritised for implementation.
5.2.2. ARC Activity Two: Results
ARC Activity Two generated a rich dataset of 189 participant-generated dialogues. Our thematic analysis of these dialogues (as in Section 4.2) generated seven potential features to address user needs. We summarise the seven features below and in Table 1, map examples from participant-generated dialogues onto each feature. These features were distilled into design specifications that directly informing the next iteration.
Table 1. Mapping Parent-Written Dialogues from ARC to NurtureBot’s Features
Dialogues written by parents Mapped Features
“Hi welcome to NurtureBot your online support and friendly advice guru for all your parenting needs! Let’s get started with what’s important and that’s how you are? What would you like support with today? :)” Feature 1. Use of Metaphors for Understanding
Feature 2. Beginning with Empathy
“Remember, if you find that this exercise isn’t quite what you need, feel free to ask for an alternative. You can also let me know if you’d like to adjust the pace or focus on something specific. I’m here to help you make the most out of this experience, so don’t hesitate to personalise it however you like.” Feature 2. Beginning with Empathy
Feature 5. Personalising and Deepening Sessions
Feature 6. Offering Clarification and Reassurance
Feature 7. Summarising Exercises and Offering Follow-Up Options
“Was this support helpful? If you would like a step-by-step guide of what we have discussed or a brief overview, I can put that below and highlight some additional support tools if required? Your happiness is important! I would love to support you further.” Feature 4. Providing Step-by-Step Guidance
Feature 6. Offering Clarification and Reassurance
Feature 7. Summarising and Offering Follow-Up Options
”Don’t worry, you are not alone! It is very common to face this problem as a new parent. Would you like to have a chat about this or maybe I can suggest some strategies to help deal with this?” Feature 2. Beginning with Empathy
Feature 3. Introducing Features Clearly
Feature 6. Offering Clarification and Reassurance
“I am here to help you through your journey as a new parent. It can often be an overwhelming time so if you have anything specific I can help with or you just need to use me as a sound-board—I am here for you, day or night.” Feature 1. Use of Metaphors for Understanding
Feature 2. Beginning with Empathy
Feature 6. Offering Clarification and Reassurance
“Hello! I’m here to support you on your parenting journey like a trusted friend! To get started you can just describe what you need in your own words. For example, you can say ‘I need to talk,’ ‘Help me relax,’ or ‘I have a question about my child’s development.’ I’m here to help in whatever way works best for you!” Feature 1. Use of Metaphors for Understanding
Feature 2. Beginning with Empathy
Feature 3. Introducing Features Clearly
Feature 7. Summarising and Offering Follow-Up Options
“I’m here to help answer or guide any questions you have about parenting, or how to manage the stress of it all. First of all, I’d like to know a bit more about you? How are you today?” Feature 2. Beginning with Empathy
Feature 3. Introducing Features Clearly
Feature 5. Personalising and Deepening Sessions
“You’re in control and I will follow your lead so please let me know how you would like this conversation to go.” (Would be nice to have something eg a button that when tapped/hovered over suggests ways you can influence responses.) Feature 5. Personalising and Deepening Sessions
Feature 7. Summarising and Offering Follow-Up Options
“Parenting can be tough on the mind and body so let’s explore ways as what gentle exercises can support us with. Would this be helpful for you?” Feature 2. Beginning with Empathy
Feature 3. Introducing Features Clearly
Feature 6. Offering Clarification and Reassurance
“If you need more clarifications on what I can do, please ask I can present information in other formats such as tables and bullet points.” Feature 4. Providing Step-by-Step Guidance
Feature 6. Offering Clarification and Reassurance
“There are a number of different tools that would help you find the right balance when facing these challenges. Would you like me to list all the different tools, or would you like to provide some more details on what specific kind of tools you are looking for?” Feature 3. Introducing Features Clearly
Feature 4. Providing Step-by-Step Guidance
Feature 5. Personalising and Deepening Sessions
Feature 1. Use of Metaphors for Understanding: Parents suggested a range of metaphors to help them better understand and form a meaningful connection with NurtureBot. Rather than perceiving the chatbot as a purely functional AI assistant, many participants preferred to imagine NurtureBot as a trusted companion, often referring to it as a “friendly advice guru.” This metaphor conveyed the idea of an accessible, approachable figure offering guidance and support in a non-judgemental manner. Others went further, likening NurtureBot to non-interactive entities such as a “personal diary” or a “soundboard,” highlighting its role as a private space where they could express thoughts and feelings without interruption. These metaphors gave NurtureBot a more intimate, relatable persona, which parents felt made it more approachable and trustworthy. They wanted NurtureBot to introduce itself as a “trusted friend” or “assistant,” someone they could confide in without fear of judgement, a crucial aspect of the chatbot’s role in supporting parents through the often overwhelming journey of early parenthood.
This metaphorical framing was not merely about making NurtureBot feel more approachable; it also shaped how parents engaged with it. For instance, treating the chatbot as a “friend” with whom they could converse “in their own words” alluded to understanding a more natural, free-flowing interaction, allowing parents to speak candidly about their concerns in natural language. The metaphors enable parents to treat NurtureBot as an alternative (digital) companion; one that was aware of their individual queries and specific needs. This blending of technology with a human-like empathy was key to fostering deeper, more effective engagement.
Feature 2. Beginning with Empathy:
Parents highlighted the importance of empathetic conversation starters, expressing that this initial approach sets the tone for the entire interaction. They wanted NurtureBot to start by first asking users how they were feeling or inquiring about their day, creating a welcoming and supportive tone. Parents expressed that this approach was particularly useful in emotional situations, where a simple, “How are you today?” can help, offering comfort and reassurance before moving on to problem-solving.
Moreover, empathy needed to extend beyond the initial greeting, especially after they disclosed a personal problem or negative feeling. Rather than immediately diving into solutions or resources, NurtureBot’s ability to show understanding and validation before proposing options was seen as key to making parents feel heard and supported. Whether engaging in mindful exercises or offering parenting information, the empathetic tone had to persist throughout, without pretending to be human or to understand human pain, thereby reinforcing NurtureBot’s supportive role.
Feature 3. Introducing Features Clearly:
A recurrent goal of parents was to highlight the importance of clearly introducing NurtureBot’s core features early in the interaction. While participants recognised that a comprehensive list of features was appropriate for the first onboarding session, they also left clear reminders of NurtureBot’s capabilities that should be available throughout subsequent interactions.
In practice, parents wanted NurtureBot to weave in these features naturally during conversations, reminding them about available options and how to best utilise them in context. This kept the interaction fluid and allowed parents to feel that they could rely on the chatbot’s range of support without needing to recall its functionalities from memory. By continuously reinforcing NurtureBot’s capabilities, within the context of the conversation, it became easier for parents to make informed choices during the interaction, thereby maximising the effectiveness of each conversation.
Feature 4. Providing Step-by-Step Guidance:
Parents wanted NurtureBot to go beyond simply providing information. When guiding them through exercises—whether mindfulness activities, meditations, or stretches—they desired a step-by-step approach that offered a clear, manageable flow of actions. They wanted Nurture to ask when they were “ready,” to “wait for them” to finish a step before moving on to the next, or even to “pause” and “continue later.” This guidance needed to be personalised to their situation, transforming NurtureBot into a collaborative tool rather than just a passive source of information.
Beyond wellbeing exercises, parents also wanted this step-by-step guidance for other interactions, such as when presenting parenting information or helping them manage their schedules. Breaking down long paragraphs of information into bullet points and allowing them to control the flow of the conversation through a step-by-step guide, with parents ”choosing the next step” from a range of available options, empowered parents, making them feel in control of the interaction and able to absorb information at their own pace.
Feature 5. Personalising and Deepening Sessions:
Personalisation was crucial to the parents’ experience with NurtureBot. Parents did not just want generic advice pulled from trusted resources. They wanted the chatbot to understand their specific situation before suggesting exercises or offering advice. Parents had rewritten NurtureBot’s dialogues to either “browse a list of topics” or “dive deeper” into one specific area depending on their immediate needs. They also wanted NurtureBot to ”talk about their situation,” for some exchanges before going about finding useful information. This deep personalisation allowed for a more meaningful and tailored exchange, where NurtureBot could address the unique concerns of each parent based on the context of their lives.
Feature 6. Offering Clarification and Reassurance:
At the end of each interaction, parents wanted NurtureBot to provide them with an opportunity to “clarify” or “revisit” the topic at hand. Whether they had completed an exercise or received parenting advice, participants wanted NurtureBot to ask if they needed further clarifications or wanted to explore additional options. This ongoing dialogue reassured parents that they were in control of the conversation, allowing them to slow down and fully understand the information presented and improve outcomes.
During the course of the conversation, when parents were unsure of what they wanted, participants wanted NurtureBot to “take the conversation slow,” offering clarification or suggesting alternative approaches, would ensure that they did not feel rushed or overwhelmed. This reassurance, especially when discussing difficult topics, coupled with the option to fine-tune their experience, would take NurtureBot beyond offering static advice.
Feature 7. Summarising and Offering Follow-Up Options:
Summarisation played a key role in allowing parents to reflect on their interactions and “takeaways” with NurtureBot. At the end of each session, parents desired a brief summary of what had been discussed, along with suggestions for follow-up actions. This summary could be “saved for future reference,” allowing parents to revisit important information without needing to restart the conversation. Participants also wanted Nurture to offer to “bookmark, save, or email” summaries. Additionally, they wanted NurtureBot to offer suggestions for similar exercises, revisit the session, or offer further support if the previous interaction was not sufficient.
Finally, participants wanted NurtureBot to link to buttons or hints to click through various follow-up options, helping parents with choices to guide the conversation forward without needing to articulate their next steps. Imagination of this interactive flow of suggested participants wanted to prevent decision fatigue with more seamless and intuitive interactions, reinforcing the effectiveness of NurtureBot as a comprehensive support tool even for new and non-technical users.
Our findings from ARC in Part III, eliciting key user needs to have a successful chat with Nurture were then directly employed to implement NurtureBot v2 in Part IV.
6. Part IV: The Interaction Layer: NurtureBot v2
To continue addressing RQ3 - “How can we employ co-design methods to craft interactions around AI, enhancing usability and the overall user experience?.” The development of NurtureBot v2 focused on refining interactions to enhance its user experience. We created an interaction layer central to how the LLM-based chatbot engages with users (see prompt illustrated in Figure 7). The three aspects of NurtureBot v2’s interaction layer, namely the core interaction principle of NurtureBot, interaction levels within each feature (denoted by four big blue dots in Figure 7, and interaction elements further embedded within interaction levels (denoted by multiple big yellow dots in Figure 7) were directly shaped by the ARC co-design sessions and implemented using a structured prompt architecture that ensured consistency and depth in user interactions. Our design approach extends Beaudouin-Lafon’s theory of instrumental interaction (Beaudouin-Lafon, 2000), adapting it to LLM-based AI agents by creating three instruments within the interaction layer, invisible in the user interface, that reify user needs and are polymorphised, and reused to improve situational interaction across the prompt architecture.
6.1. Core Interaction Principle
NurtureBot v2’s development was guided by three foundational capabilities we wanted users to have in relation to the chabtot, and vice-versa: understanding, control, and improvement. These core interaction principles (denoted with a yellow rectangle in Figure 7) guided all interactions with users. This model, inspired by the concept of the ‘cybernetic loop’ (Patten and Odum, 1981), and supported by the human-centred AI research endeavours discussed in Section 2.2, was designed to ensure that every exchange was meaningful, user-centred, and adaptable to the individual needs of each parent. This principle provided a template for each type of exchange, guiding the chatbot on how to initiate conversations, present options, and offer follow-up support.
6.2. Three Interaction Levels
1. Understand Level: The chatbot’s first task was to understand and acknowledge the user’s situation while helping the user grasp how the chatbot could assist them. For instance, it would ask, ”How are you feeling today?” or ”I can help you with a list of options?” We employed the metaphors of ”friendly AI support” and ”AI assistant” to prevent users from anthropomorphising the chatbot, ensuring they maintained a clear understanding of its role as a helpful and empathetic tool rather than attributing human relationships to it. This phase was essential for setting a supportive tone and ensuring the chatbot could convey its capabilities and tailor its responses to meet the user’s needs effectively.
2. Control Level: Once the user’s needs were understood, the chatbot entered the control level, where it provided options to guide the interaction along with its content. This level was designed to give users a sense of agency and control over the conversation. Users could choose to elaborate on their concerns, request a summary, or switch topics. Users would be often provided with ”hints” such as revise, restart, and elaborate; designed in ARC sessions, in accordance with the context of the chat to enable them to effectively lead NurtureBot towards desired outcome.
3. Improve Level: After addressing the user’s needs, NurtureBot gave a summary and sought feedback to access user satisfactions and improve the interaction by offering additional resources or to revise the conversation (as parents preferred). This phase ensured that the conversation concluded on a constructive note, always asking for user feedback, with the user feeling supported and informed.
Figure 7. This figure illustrates the prompt of NurtureBot v2, with interaction layer showcasing how the chatbot guides user engagement through the core principle of ”Understand, Control, and Improve,” with a core interaction principle, interaction levels and elements answering in-situ user needs identified in ARC.
[…]
A detailed diagram of NurtureBot v2’s prompt architecture hosting an interaction layer, divided into core principles: Understand, Control, and Improve. It highlights the chatbot’s empathetic chatting, wellbeing exercises, and parenting information, showing how the system tailors responses based on user preferences and inputs. Examples of user dialogues, actionable hints, and resources are shown alongside elements for adapting to contextual user needs, illustrating NurtureBot’s design for personalised support.
6.3. Interaction Elements
1. Undecided Exchanges When users were undecided, NurtureBot aimed to first understand their hesitation or indecision. With started elements such as, ”What’s on your mind today?” or ”How can I assist you?” to encourage the user to reflect on their needs. Metaphorical elements like ”assistant” were also embedded in the chatbot’s dialogue to offer a supportive role, and reassurance elements were designed to not overwhelm user with options, reflecting a tone of gentle guidance. For instance, ”I’m here to support you in whatever way works best for you” to help users feel comfortable in sharing their thoughts or challenges.
2. Empathetic Chatting: The chatbot was programmed with elements to focus on active listening and providing emotional support. We used dialogue examples to create non-judgemental elements and avoiding any advice. The chatbot’s responses were crafted to mirror those of a trusted confidant, providing reassurance and understanding.
For example, these elements instructed NurtureBot to initially wait: “Tell me more about what’s been on your mind today.” Based on the ARC study insights, the chatbot would then offer subtle prompts to encourage to-and-fro discussion, e.g., “That sounds really challenging. How has that been affecting you?” This approach was designed to build trust and ensure that the user felt heard and understood.
3. Wellbeing Exercises: Guiding elements for wellbeing exercises included mindfulness practices and relaxation techniques. The interaction was structured with personalised elements, offering tailored exercises based on user concerns and alternative options. The chatbot also ensured that exercises were broken down into manageable steps, with the user controlling the pace of the session. To embed collaborative elements, the prompts also ensured that NurtureBot could guide users through activities step by step, e.g., ”Would you like to try a simple breathing exercise together? Let me know when you’re ready to begin.” As such, the user could control and personalise the pace and depth of the exercise.
4. Parenting Information: NurtureBot v2 provided a structured and digestible format for parenting content. Instead of overwhelming users with lengthy text, interactive elements offered summaries and gave users the option to dive deeper into specific topics. The chatbot also cross-referenced reliable sources to ensure that the information was accurate and relevant. The offered information was also adapted, based on the specific situation, preference, and needs. For instance, if a user asked about managing a fussy baby, the chatbot might respond, ”I have a few techniques that could help. Would you like a quick overview, or shall we talk about your baby and things you have tried already?” This ensured that users received information that was most useful to them.
NurtureBot v2, featuring an interaction layer created in Part IV with parent-designed dialogues, was tested in Part V to ensure that the issues identified in the previous version were effectively addressed.
7. Part V: Technology trial of NurtureBot v2
7.1. Participants, Method, and Data
The initial set of participants (See Section 4.1) trialled NurtureBot v2 over five days, generating a total of 104 conversations and their corresponding survey feedback (See Figure 2). We analysed the survey data in accordance with our thematic approach (See Section 4.2). Participation averaged 69.3% (20.8 out of 30), a decline from 90% in Part II (v1 trial) and 91.6% in Part III (ARC session). Eight participants contacted via Prolific cited reasons such as, ”I’ve enjoyed your study, but NurtureBot has improved so much I don’t know what more I can give feedback on.”
7.2. Findings of second technology trial - NurtureBot v2
While our data revealed highly positive feedback for NurtureBot v2, with participants praising the chatbot for being “more chatty” and “engaging” than its predecessor, as with NurtureBot v1, we only focused the key interaction problems for our thematic analysis, focusing on 78 instances of negative feedback and/or suggestions for improvement (of 208 instances of feedback collected (37.5%)), coding 12 unique pain points, which we then grouped into three themes briefly summarised below. We further discuss a comparative analysis of our findings in the Results section (see Section 8.3).
Problem 1 Some participants pointed out that NurtureBot still occasionally responded with long paragraphs, making the information feel overwhelming. One participant noted,
“It was good, but very long answers though and lots of information to take in.”
Problem 2 Most participants noted a bug, where it repeatedly used the same prompts—“elaborate, revise, and restart”—which reduced the creativity and variety in responses. A participant mentioned,
“The repetition of the last sentence ’if you need me to elaborate… just let me know’ was quite annoying.”
Repeating Themes As with v1, participants made suggestions for incorporating buttons for conversational hints, multi-modal content (images and videos), and memory capabilities for further personalisation, and connection to resources and communities in their local area.
7.2.1. Quantitative results.
Our subjective questions regarding NurtureBot v2’s task satisfaction received an average score of 4.41 out of 5, while interactive questions averaged 4.50 for understandability, 4.42 for controllability, and 4.44 for improvability. The questionnaire analysis resulted in a mean CUQ score of 87.2 out of 100, which, when compared to the CUQ benchmark of 68 (Holmes et al., 2019) in a one-sample t-test, was extremely statistically significantly (t(20) = 8.98, p ¡ 0.0001, mean difference = 19.20, 95% CI [14.74, 23.66], SE = 2.14). These scores and their implications are discussed further in Section 8.3, alongside comparisons with NurtureBot v1 and v3.
The two minor interaction issues persisting in NurtureBot v2, identified in Part V, were targeted for refinement in NurtureBot v3 and tested with a new cohort of participants in Part VI to further validate our efforts.
8. Part VI: Prototype Improvement: NurtureBot v3 and Validation
We iterated NurtureBot v3 as briefly summarised here:
  • First, we addressed the issue of lengthy responses by limiting the chatbot’s answers to 50 words, ensuring information was delivered in concise, digestible segments.
  • Second, we updated the core interaction principle, enhancing NurtureBot’s ability to provide contextual actionable hints, rather than offering generic options (denoted as struck-out text inside the big yellow rectangle in Figure 7)
8.1. Participants, Method, and Data
To assess the effectiveness of these improvements without learning and co-design biases, NurtureBot v3 was tested with a new cohort of 46 parents, 29 females and 17 males with an average age of 38.07. 11 participants (23.9%) used ChatGPT everyday. A total of 185 conversations (denoting an average participation of 80.4% (37 out of 46 parents across all days)) and corresponding feedback were collected over five days using the same data collection (See Figure 2) and analysis method (see Section 4.2) as outlined in the NurtureBot v1’s trial, allowing for direct comparisons across versions.
8.2. Findings of third technology trial - Nurture Bot v3
Next we present a summary of the third trial findings, focusing on identifying ongoing interaction problems, following the interaction layer redesign. While feedback was overwhelmingly positive, our analysis focused the interaction problems (from 46 out of 370 (12.4%) negative feedback. We identified 10 unique pain points, which we group into the three themes, summarised below. We also provide a comparative analysis of the three Nurturebot versions in Section 8.3.
Problem 1 Some participants wanted NurtureBot to “intelligently” provide more detailed responses when necessary. For example,
“I guess the responses could be a little longer and detailed at times but not by much.”
Problem 2 Few participants found it challenging to follow step-by-step exercises via text. One participant remarked,
“A lot of the time the advice was to close my eyes then do X but when it’s text based this is obviously hard to follow. Perhaps tell them to read the advice then close eyes.”
Repeating Themes Participants continued to request features for: multi-modal interactions (video, speech, and images), improved memory function to allow NurtureBot to remember personal details and past conversations, local resources, and hints to continue conversation as buttons. Furthermore, one participant expressed interest in viewing the original sources referenced in NurtureBot’s responses, allowing them to verify or explore the information further.
8.2.1. Quantitative results.
Our subjective questions regarding NurtureBot v3’s task satisfaction received an average score of 4.49 out of 5, while interactive questions averaged 4.63 for understandability, 4.04 for controllability, and 4.21 for improvability. The questionnaire analysis resulted in a mean CUQ score of 91.3 out of 100, which, when compared to the CUQ benchmark of 68 (Holmes et al., 2019) in a one-sample t-test, was extremely statistically significant (t(35) = 18.39, p ¡ 0.0001, mean difference = 23.30, 95% CI [20.73, 25.87], SE = 1.27). These score and their implications are discussed further in Section 8.3, alongside comparisons with NurtureBot v1 and v2.
8.3. Results: NurtureBot v1 vs. NurtureBot v2 vs. NurtureBot v3
8.3.1. Qualitative Results
The comparison of qualitative results from NurtureBot v1, v2, and v3 reveal a clear progression in user satisfaction across versions, supported by a reduction in negative feedback gathered from our analysis 45.48% in NurtureBot v1 to 12.4% in NurtureBot v3. While our participants praised the “natural flow” of conversations and the “supportive, empathetic tone” of our improved versions, repeating themes for features like multi-modal interactions and enhanced memory for personalisation and localisation persisted across versions. In this section we detail our qualitative findings from evaluations of NurtureBot’s evolutions.
A Successful Chat In NurtureBot v1, participants found the chatbot useful but expressed frustration with its tendency to cut off conversations, leading to a lack of conversational flow. Several participants noted that the responses were often patronising or generic. Informed by user feedback and ARC co-design, NurtureBot v2 introduced an interaction layer centred on the principles of Understand, Control, and Improve, enabling the chatbot to better grasp user needs, offer greater conversational control, and dynamically adapt across all features, giving users more autonomy and flexibility. We highlight these adapted elements in bold.
Participants expressed satisfaction with NurtureBot v2, noting improvements in how natural and intuitive the conversations felt. One participant shared, “Good, lots of ideas for activities for playtime for my baby which was good, and Nurture continued to ask if I wanted further information at each step.” Participants also liked the system’s flexibility, saying, “The interaction was easy again, it feels like Nurture can handle whatever you ask it, whether it’s just chatting or asking for specific information.” Participant feedback also emphasised how the interaction layer made the conversation flow seamlessly, with one participant noting, “Interaction was the best this week—the responses from Nurture seemed natural and clear.” The introduction of the interaction layer also improved the clarity and engagement of the system. One participant mentioned, “It was really good. I found the introduction really helpful, particularly the clarification of the three kinds of things that Nurture could do.” Another shared, “Great! It made such a difference having the prompt to tell Nurture what I wanted.” This interaction layer enabled participants to feel more in control and allowed the chatbot to adapt to user input, creating a personalised, engaging experience.
Although NurtureBot v2 was tested by participants who had co-designed it, NurtureBot v3 was tested with a new cohort, which confirmed the success of the interaction layer. One participant remarked, “Good. Conversation flows well, and everything provided was useful,” while another stated, “So seamless yet educative.” Our findings suggest interaction layer to played a crucial role in ensuring that NurtureBot responded dynamically to user input, leading to an improved overall experience.
The Time Effect A key observation from the five-day testing period was the gradual shift in participants’ attitudes toward NurtureBot. Initially, participants were hesitant and skeptical, viewing it as an unfamiliar tool and doubting its ability to provide meaningful support. However, as they continued using NurtureBot over consecutive days, their perceptions became more positive, with many describing it in personal terms—some even referring to it as a “friend always there” to support them.
For instance, one participant testing NurtureBot v1 described their early experience as “detached and robot-like,” but by day five, they said, “After using Nurture quite a few times now, it just feels like I’m chatting with someone that feels familiar and I have already built up a bit of a relationship with. It doesn’t feel artificial.” The time effect led many participants to believe that NurtureBot was being updated continuously each day. One participant testing NurtureBot v3 commented, “Wow, it was so much better today. I wrote quite a lengthy update, and it acknowledged every part, even pinpointing what was bothering me.” Participants began using personal pronouns like ”they” to describe NurtureBot v1. On day four, one participant remarked, “I think it is getting there with becoming more personal, maybe ask what the user’s name is, so they could talk using your name,” reflecting an increased emotional connection with the chatbot.
This progression illustrates a time effect, where repeated interactions led to increased comfort, attachment, and even anthropomorphisation of the technology. The more time participants spent engaging with NurtureBot, the more they saw it as a supportive companion rather than just a tool.
Community and Peer Support Tying back to the origins of the Nurture project, which was a community-based peer-to-peer mentoring program (as described in Section 3), it is clear that the concept of community and peer support remains a significant need for participants using NurtureBot. Throughout the three versions of NurtureBot, participants frequently searched for local activities, community support centres, and peer forums. This theme of seeking peer support and localised resources also emerged as the second biggest problem identified in the ARC (see Section 4). Additionally, several participants expressed a desire for more localised community resources, indicating the importance of having access to peer support networks within their geographical area. As one participant trying NurtureBot v2 remarked, “When I asked for classes in my local area it could not help, that would be a good feature.” and another trying NurtureBot v3 said “I realised when asking a question about services in my local area that it does not know my local area.” Feedback from our study suggests that there is an opportunity to evolve NurtureBot from a solely AI-driven one-on-one platform to a more comprehensive community-based support system.
Personalisation and Contextualisation The third biggest challenge prioritised in our ARC session 5.1, the lack of personalisation and contextualisation in the chatbot’s responses, also emerged throughout the testing of all three versions. Participants repeatedly expressed the desire for NurtureBot to remember personal details such as their name, their child’s name, their unique circumstances, and previous conversations. Participant feedback such as, “Remembering my problems would still be nice,” and “The bot could ask our names and baby’s details to help make it more personal.,” in NurtureBot v2 and NurtureBot v3 echoed this need, emphasising how critical personalisation was to creating a more meaningful user experience.
Participants consistently wanted the chatbot to feel more tailored to their individual experiences. While our experiment was not designed to collect or store personal data, and personalisation was not a focus in any version of NurtureBot, going forward, addressing personalisation and contextualisation will be essential to improving NurtureBot’s effectiveness. While safeguarding privacy and data security remains paramount, finding a way to create more contextually aware and personalised interactions will be crucial in meeting user expectations.
8.3.2. Quantitative results
To analyse our subjective questionnaire with four questions we created to measure task satisfaction and human-AI interaction (see questions 3-6 in Figure 2). To the data we collected for these questions, across the three technological trials (in Sections 4, 7, and 8), we employed both parametric Generalised Estimated Equations (GEE) and non-parametric Kruskal-Wallis and Dunn’s test approaches (see Table 2 and Figures 8, 9, and 10). Our results suggest:
  1. (1)
    Meeting User Expectations: GEE showed no significant differences between versions (p ¿ 0.05). Kruskal-Wallis yielded marginally significant results (H = 5.8539, p = 0.0536), suggesting potential subtle differences not captured by the parametric model. However, the changes between v1, v2, and v3 are not robust enough to be considered significant.
  2. (2)
    Understanding Functionality: Both parametric and non-parametric tests indicated significant improvements. GEE showed Version 3 to significantly outperformed Version 1 ($\beta$ = 0.2721, p = 0.001). Kruskal-Wallis and Dunn’s test confirmed this finding, revealing a significant difference between Version 1 and Version 3 (p = 0.003).
  3. (3)
    Control Over Actions: GEE indicated that Version 2 significantly improved user control compared to Version 1 ($\beta$ = 0.4521, p ¡ 0.001), which was evident considering v2 was designed by v1 participants. Interestingly, Kruskal-Wallis and Dunn’s test not only confirmed this (p ¡ 0.001) but also revealed a significant difference between Version 2 and Version 3 (p = 0.017) that wasn’t apparent with GEE.
  4. (4)
    Improving Performance: Both analytical approaches showed significant improvements in later versions. GEE model indicated that both Version 2 ($\beta$ = 0.5582, p ¡ 0.001) and Version 3 ($\beta$ = 0.3256, p = 0.005) significantly outperformed Version 1. Kruskal-Wallis and Dunn’s test corroborated these findings, showing significant differences between Version 1 and both Version 2 (p ¡ 0.001) and Version 3 (p = 0.007).
Table 2. Comparison of User Experience with Subjective Questionnaire across NurtureBot’s Versions
Question GEE Model Kruskal-Wallis Dunn’s Test (p-values)
1. Meeting hopes and goals V2: $\beta$=0.1257, p=0.284 H=5.8539 Not performed
V3: $\beta$=0.2015, p=0.061 p=0.0536
2. Understanding functionality V2: $\beta$=0.1377, p=0.146 H=11.1215 V1-V2: p=0.727
V3: $\beta$=0.2721, p=0.001** p=0.0038** V1-V3: p=0.003**
V2-V3: p=0.225
3. Control over actions V2: $\beta$=0.4521, p¡0.001** H=13.5127 V1-V2: p¡0.001**
V3: $\beta$=0.0666, p=0.573 p=0.0012** V1-V3: p=0.757
V2-V3: p=0.017*
4. Improving performance V2: $\beta$=0.5582, p¡0.001** H=19.9734 V1-V2: p¡0.001**
V3: $\beta$=0.3256, p=0.005** p¡0.0001** V1-V3: p=0.007**
V2-V3: p=0.201
Note: * p ¡ 0.05, ** p ¡ 0.01. V1, V2, and V3 represent Version 1, Version 2, and Version 3 of NurtureBot, respectively. GEE coefficients ($\beta$) represent the change relative to Version 1.
While the parametric test provided insights into overall trends, the non-parametric tests revealed additional pairwise differences, particularly between Version 2 and Version 3 in terms of user control, where we had changed the bug of repeating conversational hints. This comprehensive analysis suggests that the iterative development of NurtureBot has led to significant improvements in user understanding and ability to improve performance, with some complex dynamics in user control across versions.
Figure 8. Generalised Estimating Equation (GEE) coefficients comparing chatbot versions v2 and v3 to v1 of NurtureBot across four evaluation dimensions: (1) whether the task met user expectations, and interactive dimensions of: user- (2) understanding, (3) control, and (4) ability to improve performance and outcomes. Statistically significant improvements (p ¡ 0.05) are noted in dimensions (2), (3), and (4) in NurtureBot v2 and v3.
[…]
Figure 9. Kruskal-Wallis test results comparing the NurtureBot’s versions across four dimensions of subjective user experience. The test identifies statistically significant differences (p ¡ 0.05), with the strongest significance observed in the ”ability to improve performance and outcomes” dimension.
[…]
Figure 10. Heatmaps illustrating the pairwise comparison of three versions of the chatbot across three evaluation dimensions: (user- (2) understanding, (3) control and , and (4) ability to improve performance and outcomes. Statistically significant differences (p ¡ 0.05) are observed primarily between version 1 and version 3 across the evaluated dimensions. A reduced variance as NurtureBot is iterative developed, also indicates chatbot’s reliable performance over time.
[…]
Figure 11. Comparison of CUQ Scores Across NurtureBot Versions. The figure demonstrates the improvement in Chatbot Usability Questionnaire (CUQ) scores from NurtureBot v1 with a CUQ score of 85.4, to NurtureBot v2 with a CUQ score of 87.2, and NurtureBot v3, which achieved a higher CUQ score of 91.3. The lowest scores also improved significantly across versions, indicating enhanced chatbot usability with each iteration.
[…]
A graph comparing the CUQ scores of NurtureBot v1, v2, and v3. The x-axis lists the chatbot versions, while the y-axis displays the CUQ scores from 0 to 100. NurtureBot v1 has a score of 85.4, v2 scored 87.2, and v3 achieved the highest score of 91.3. The graph shows the lowest, median, and highest scores for each version, with a steady increase in the lowest scores from v1 (43.8) to v3 (75), while the median and highest scores remain consistently high, demonstrating improvements in chatbot usability over time.
Further, quantitative results from the structured CUQ (Holmes et al., 2019) indicate a steady improvement in usability across all three versions of NurtureBot. As seen in the CUQ score comparison in Figure 11, the average CUQ score increased from 85.4 in NurtureBot v1 to 87.2 in v2, and further to 91.3 in v3. Notably, the lowest scores also saw significant improvements, rising from 43.8 in v1 to 67.2 in v2, and then to 75 in v3, reflecting enhanced user satisfaction over time. These results confirm that the iterative design process enhanced user experience, making the chatbot more intuitive, engaging, and easier to navigate. From NurtureBot v1 to v3, positive feedback improved significantly, with users rating the chatbot as increasingly engaging, easy to navigate, and better at understanding and supporting their needs (see Figure 12 Left). Negative feedback, such as the chatbot being perceived as ”robotic” or ”unfriendly,” decreased with each version, showing that iterative improvements effectively addressed these issues (see Figure 12 Right). By v3, users reported the most satisfying experience, with fewer errors and a more natural interaction flow, and a decreased overall variance (see Figure 11).
Figure 12. Comparison of CUQ questions of three Versions of NurtureBot. This figure displays the differences between NurtureBot v1, v2, and v3 across various chatbot usability dimensions. The left column contains positive feedback questions, while the right column contains negative feedback question, showing clear improvements across versions, particularly in how engaging and easy to use the chatbot became over time.
[…]
A comparison chart showing the usability scores of NurtureBot v1, v2, and v3. The figure contains two sets of bar graphs for each version: one set for positive statements (left) and another for negative statements (right). NurtureBot v1 (blue) shows more neutral scores, while v2 (yellow) and v3 (pink) show improvements in positive responses and a reduction in negative responses. This comparison demonstrates progressive enhancement in user satisfaction across all chatbot versions.
Furthermore, in comparison with literature at the time of writing this paper, to the best of our knowledge, NurtureBot v3 had achieved the highest reported CUQ score of 91.3 (See Table 3). While early rule-based systems like MYA using FlowXo, and ChatPal using Rasa (Boyd et al., 2022) reported lower scores when compared to more advanced NurtureBot versions using GPT-4 and DIAS using GPT-4 Turbo (Fersch et al., 2022), which achieve significantly higher scores. Notably, despite all NurtureBot versions using the same base technology, the iterative development of a human-centred interaction layer led to significant improvements in usability, demonstrating that thoughtful design can enhance user experience even without changes to the underlying technology.
Table 3. In comparision with CUQ scores reported in literature NurtureBot v3 achieves the highest CUQ score of 91.3, followed by NurtureBot v2 and v1, demonstrating progressive improvements in usability, highlighting the effectiveness of iterative human-centred design in enhancing chatbot performance.
Chatbot Name Base Technology Domain Participants CUQ Score
MYA(Larbi et al., 2022) FlowXo (Rule-based) Physical Activity 30 57.4
ChatPal(Boyd et al., 2022) Rasa Mental Health 10 65
Chat Ella(Zhang and Song, 2024) GPT-2 Chronic Diseases 64 68.31
GPT-Simulated Patient(Holderried et al., 2024) GPT-3.5 Medical Education 28 77
DIAS(Fersch et al., 2022) GPT-4 Turbo Educational Services 103 80
NurtureBot v1 GPT-4 Parental Wellbeing 32 85.4
NurtureBot v2 GPT-4 Parental Wellbeing 32 87.2
NurtureBot v3 GPT-4 Parental Wellbeing 64 91.3
Based on our findings for co-designing user-LLM interactions, in Parts I-VI, we derive implications for design of AI-driven parental support systems presented in Table 4.
Table 4. Implications for Design of AI-driven Parental Support Systems
Implication Description
Good Data Sources for AI A key foundation for successful AI is access to reliable and well-structured data sources. NurtureBot’s effectiveness depended on the quality and relevance of the information provided. Future iterations must use up-to-date, verified sources, including health organisation-endorsed health guidelines and localised resources, to enhance trust and provide contextually relevant information.
User Understanding, Control, and Outcome Improvement Effective human-centred AI balances understanding, controlling, and improving interactions. Users should have transparency in their interactions with the AI and be able to guide the interaction through dynamic in-situ controls. Enabling users to refine the agent’s responses enhances the quality of interactions and makes the AI more responsive to user needs.
Personalisation, Contextualisation, and Localisation Users expect personalised interactions, asking for the chatbot to remember personal details such as their names, their child’s name, where they live, and their previous conversations. These aspects help the chatbot tailor responses to user’s specific needs. Future designs should consider secure methods for personalisation without compromising privacy.
Community and Peer Support Users desire more community-oriented support for parental wellbeing. This reflects the need for AI systems to connect users with localised communities, parenting groups, or forums. AI could act as a bridge between users and real-world community resources, providing emotional and practical support.
Multi-modal Interactions Users expressed a need for multi-modal interactions such as video, voice, and image-based communication alongside text. This allows flexibility in how users interact with the chatbot, broadening accessibility for users with different preferences, literacy levels, or disabilities.
Emotional Design Emotional connection emerged as a key element of success for NurtureBot. Many participants referred to the assistant as a ”friend,” indicating the importance of empathy and emotional support. However, it’s crucial to ensure users understand the AI’s role as a tool and avoid fostering over-reliance.
Co-Design and Iterative Testing Co-design played a critical role in aligning NurtureBot’s features with user needs. Iterative testing allowed us to refine the system based on feedback. This iterative process of testing, redesigning, and retesting ensured that the final product provided the best possible user experience.
9. Discussion
The co-design of interactions with AI within the NurtureBot project highlights several key insights into how parental wellbeing support systems can be improved through iterative user feedback and involvement. Prior literature underscores the importance of human-centred AI design, particularly in sensitive areas like mental health and parenting, where emotional engagement and trust are crucial (Luchs, 2023; Shankar et al., 2024; Ferrara, 2022; Ke et al., 2024). Through iterative, parent feedback, our NurtureBot project explored the main interaction problems with LLM-based parental-wellbeing assistant when simulating support-seeking scenarios (RQ1), what their key needs and priorities would be (RQ2), and how we might co-design better human-AI interactions (RQ3). Our initial trial of Nurturebot v1 revealed seven key problems (RQ1), comprised of: lack of conversational flow; patronising tone; no personalisation; non-localised resources; generic responses; memory gaps; and transactional interactions. We note that these problems are prevailing across wellbeing initiatives and chat-based applications in the literature (Lee et al., 2019; Hwang et al., 2019; Viswanathan et al., 2020; Garcia et al., 2018). To begin to address these, with RQ2, we focused on prioritising key user needs, and their ideal vision for how NurtureBot should function. The focus was on explaining its workings, providing control over the chat, and opportunities to improve the conversation. Finally, we examined how we could employ co-design methods to craft interactions around AI (RQ3). The ARC co-design study identified that having a “successful chat” as the key user need, alongside ideal features such as using metaphors for understanding, beginning with empathy, clearly introducing features, providing step-by-step guidance, personalising and deepening sessions, offering clarification and reassurance, and summarising exercises with follow-up options. The detailed insights gained from addressing just one of the seven key issues identified in the NurtureBot v1 trial highlight a critical point: LLM responses are only as good as the user’s input. This reinforces that, despite meticulous design, the usability of even the most carefully crafted LLM-based support systems remains highly uncertain (Zamfirescu-Pereira et al., 2023).
Continuing our efforts for answering RQ3, we further explored how we can use the insights from the ARC sessions to enhancing usability and the overall user experience. Incorporating participant feedback from ARC, in NurtureBot v2, we built an interaction layer designed to enhance user experience by focusing on three core principles: Understand, Control, and Improve. This layer afforded NurtureBot to better comprehend user needs, offer users control over the flow of conversation, and provide dynamic options for improvement through levels of interaction built in the system’s prompt. Across all of NurtureBot’s features (empathetic chatting, wellbeing exercises, and parenting information), we built interaction elements to offer more autonomy and flexibility in directing the interaction, when seeking support. We employed iterative co-design methods to leverage real-world end-user feedback in the agent’s prompt to better align it with user needs, a trend also emerging in the domain of Machine Learning (Shi et al., 2024; Hwang et al., 2023). This extends existing applications of methods that are typically used for improving LLMs, e.g.,fine-tuning, transfer learning, and reinforcement learning from human feedback (RLHF). Our approach builds on existing research through design methods in HCI (Benjamin et al., 2024) to “heard AI” to meet user needs (Zamfirescu-Pereira et al., 2023).
Our findings suggested that by incorporating a user-co-designed interaction layer, employing the tripartite human-centred AI principles of understanding, control, and improving interactions, we were able to improve the usability and user experience in subsequent versions of NurtureBot. Our research adds to this body of work leveraging human-centred computing for AI alignment (Gilbert, 2024). By enabling users to steer the conversation with a clear understanding and responsive feedback loops, LLMs can become more intuitive and aligned to diverse domains and user needs. By demonstrating that user feedback can directly shape how AI-powered parenting supporting systems are perceived and used via ARC sessions, our research aligns with broader HCI research, which shows that empowering users to co-design results in more effective and engaging systems for designing both parental wellbeing (Prabhakar et al., 2017; Benjamin et al., 2024) and broader AI (Hossain and Ahmed, 2021; Delgado et al., 2023).
10. Limitations and Future Work
While the iterative development of NurtureBot improved its interaction capabilities, limitations persist, particularly in the areas of personalisation, contextualisation, and localisation. Addressing the diverse family roles, children’s characteristics, and unique circumstances of users is crucial to ensure the support provided is relevant and tailored to individual needs. Additionally, the ”time effect” underscores the importance of investigating how users emotionally engage with AI over extended use. While NurtureBot serves as a valuable tool, careful design is needed to avoid fostering over-reliance. Lastly, testing with slightly out-of-target groups to minimise emotional risks poses a limitation, as greater risks could arise when engaging prenatal parents in future tests.
In future work, we plan to address these limitations by enhancing the situational features and designing for optimal reliance. Collaborating with public health organisations will be crucial to deploying and scaling NurtureBot on a broader platform, ensuring it reaches a wider audience of parents.
11. Conclusion
The significance of comprehensive perinatal care cannot be overstated, as it plays a pivotal role in mitigating risks associated with perinatal mind-body duality issues, fostering a positive environment that promotes healthy development of the parent(s), the child, and the long-term stability of the family unit. We explored iterative co-design of AI-powered parental wellbeing support with NurtureBot, engaging users through ARC sessions, where they were not only able to prioritise key issues but also contribute to the redesign of user-LLM interactions, for improved user comprehension, control, and task outcomes. The resulting interaction layer significantly improved the user experience across versions, suggesting that AI-driven systems can better align with real-world needs by empowering users to shape their interactions, making AI tools more effective and widely adopted in wellbeing and personalised support contexts. By prioritising the human experience, AI tools can not only enhance wellbeing but also foster stronger, more resilient communities and create scalable, long-lasting solutions to meet diverse real-world needs.

### v2

@@
 To see if these issues could be solved using a co-design approach, we developed and tested NurtureBot, a wellbeing support assistant for new parents.
-32 parents co-designed the system through Asynchronous Remote Communities method, identifying the key challenge as achieving a “successful chat.” As part of co-design, parents role-played as NurturBot, rewriting its dialogues to improve user understanding, control, and outcomes.
-The refined prototype evaluated by 32 initial and 46 new parents, showed improved user experience and usability, with final CUQ score of 91.3/100, demonstrating successful interaction patterns.
+32 parents co-designed the system through Asynchronous Remote Communities method, identifying the key challenge as achieving a “successful chat.” As part of co-design, parents role-played as NurtureBot, rewriting its dialogues to improve user understanding, control, and outcomes.
+The refined prototype, featuring an Interaction Layer, was evaluated by 32 initial and 46 new parents, showing improved user experience and usability, with final CUQ score of 91.3/100, demonstrating successful interaction patterns.
 Our process revealed useful interaction design lessons for effective AI parenting support.
-Figure 1. Overview of the six-part study detailing the iterative development and testing of NurtureBot, a parental wellbeing assistant.
+Parental Wellbeing, Perinatal Support, LLMs, Human-Centred AI, Interaction Design
+journalyear: 2025conference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3714088isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Empirical studies in interaction designccs: Computing methodologies Intelligent agentsccs: Human-centered computing HCI theory, concepts and modelsccs: Applied computing Health care information systems
+Figure 1. Overview of the four-part study detailing the iterative development and testing of NurtureBot, a parental wellbeing assistant.
 […]
-A visual representation of the six-part iterative development of NurtureBot involving two parent cohorts.
+A visual representation of the four-part iterative development of NurtureBot involving two parent cohorts.
 The chart shows the progression from the initial prototype to NurtureBot v1, followed by ARC co-design sessions, NurtureBot v2, and NurtureBot v3.
@@
 In particular, many expectant and new mothers encounter substantial challenges, in mental health, emotional adjustment, and caregiving (Widarsson et al., 2012), in addition to the transformation their body goes through.
-During this period, parents experience heightened stress, anxiety, and feelings of isolation, increasing the need for support in these areas.
-Yet, we know that one in three parents do not seek help (UNICEF UK, 2022).
-The literature on perinatal care identifies several key issues, including inadequate access to mental health resources, availability of healthcare professionals, cost to the government, stigma surrounding parental mental health, and challenges in identifying and supporting at-risk mothers (Rallis et al., 2014; Davis et al., 2022; Howard and Khalifeh, 2020).
+During this period, parents experience heightened stress, anxiety, and feelings of isolation, but only one in three parents seek help (UNICEF UK, 2022).
+Perinatal care literature identifies several key issues, including inadequate access to mental health resources, availability of healthcare professionals, cost, stigma surrounding parental mental health, and challenges in identifying and supporting at-risk mothers (Rallis et al., 2014; Davis et al., 2022; Howard and Khalifeh, 2020).
 These challenges have been further aggravated by the COVID-19 pandemic (Bridle et al., 2022; Kinser et al., 2022).
-As a result, we see a growing body of research exploring how digital technologies can enhance perinatal care and provide support during this critical period (van den Heuvel et al., 2018; Ginja et al., 2018).
-Notable Human-Computer Interaction (HCI) research has laid the groundwork for exploring how digital interventions increase wellbeing across aspects such as physical and mental health of a family (Pina et al., 2017), a child’s growth (Kientz et al., 2009), sleep tracking of parents and children (Pina et al., 2020), and managing family screen-time behaviours (Hiniker et al., 2016).
-Today, with the availability of Artificial Intelligence (AI) to everyday users, use of general generative applications such a ChatGPT for parenting is on the rise (Quan et al., 2024).
+As a result, we see a growing body of research exploring how digital technologies can enhance perinatal care (van den Heuvel et al., 2018; Ginja et al., 2018).
+Notable Human-Computer Interaction (HCI) research has laid the groundwork for exploring how digital interventions can increase wellbeing across aspects such as physical and mental health of a family (Pina et al., 2017), a child’s growth (Kientz et al., 2009), sleep tracking of parents and children (Pina et al., 2020), and managing family screen-time behaviours (Hiniker et al., 2016).
+Today, with the availability of Artificial Intelligence (AI) to everyday users, the use of general generative applications such as ChatGPT for parenting is on the rise (Quan et al., 2024).
 With responsible design, Large Language Models (LLMs) have the potential to facilitate a range of services, from offering self-guided mental health support (Sharma et al., 2024), real-time emotional support (Kang et al., 2024) to providing evidence-based information at scale (Lai et al., 2023).
-The current use of LLMs to support parental and child wellbeing is reviewed with high potential and ethical cautions (Ashraf, 2024; Stade et al., 2024).
-Their application within this context is fraught with limitations and uncertainties, especially of accuracy, empathy, and usefulness (Kaneda et al., 2023).
+The current use of LLMs to support parental and child wellbeing raises both optimism, but also ethical caution (Ashraf, 2024; Stade et al., 2024), due to limitations and uncertainties regarding accuracy of information provided, empathy, and usefulness (Kaneda et al., 2023).
 Given these challenges, our research adopts an exploratory approach, probing into a small subset of these broader issues.
-We focus on understanding how interactions between LLMs and parents might be designed to be perceived as helpful and engaging.
+We focus on understanding how interactions between LLMs and parents might be designed to be helpful and engaging.
 Our goal is to explore what constitutes a usable and useful interaction in these contexts and how LLMs can be integrated into wellbeing support in ways that are practically engaging and thus useful for parents.
-To this end, we conducted a six-part online study involving 78 parents, who trialled iterative improvements of NurtureBot, a technology probe, designed to support parental wellbeing.
+To this end, we conducted a four-part online study involving 78 parents, who trialled iterative improvements of NurtureBot, a technology probe, designed to support parental wellbeing.
 This study, while not aimed at delivering a finalised solution, served as an exploratory step toward understanding how parents interact with LLM-based agents and identifying necessary features to make these tools genuinely supportive.
@@
 Study Flow Overview
-This six-part user-centred design process, as illustrated in Figure 1, guided the iterative development of NurtureBot.
-To mitigate emotional risks, we engaged a ‘slightly-out-of-target’ group of parents, each with a child aged five or six.
+This four-part user-centred design process, as illustrated in Figure 1, guided the iterative development of NurtureBot.
+To mitigate emotional risks, we engaged a ‘slightly-out-of-target’ group of parents (Islind et al., 2023), each with a child aged five or six.
 These parents simulated stages of their perinatal period and reflected on past challenges with NurtureBot.
-Using a mixed-methods approach, we collected both qualitative and quantitative feedback, with NurtureBot v1, v2, and v3 tested over a five-day period.On the final day, participants completed the Chatbot Usability Questionnaire (CUQ) (Holmes et al., 2019), assessing usability and engagement.
-(see Figure2).
-  • Part I: The Initial Prototype: NurtureBot v1 The development of NurtureBot v1 leveraged LLM technology, informed by a community wellbeing initiative supporting new parents.
-Zero-shot prompting was used to integrate supportive conversation starters and key features, including empathetic chatting, wellbeing exercises, and parenting information (see Section 3).
-  • Part II: NurtureBot v1 Testing and Analysis A cohort of 32 parents engaged with NurtureBot v1, generating 144 conversations.
-Thematic analysis identified several issues, including lack of conversational flow, robotic tone, transactional attitude, limited personalisation, and localisation.
-Despite these challenges, NurtureBot v1 achieved a CUQ score of 85.4, surpassing the benchmark of 68 (Holmes et al., 2019), indicating good usability (see Section 4).
-  • Part III: ARC Co-Design Sessions and Analysis The Asynchronous Remote Community (ARC) (MacLeod et al., 2016; Prabhakar et al., 2017) co-design sessions engaged the same cohort of parents with Miro board activities (Miro, 2011).
-    • Parents first prioritised key issues, with “successful chat” emerging as the top priority.
-    • Then they reimagined and rewrote NurtureBot’s dialogues, aiming for better user understanding, user control, and improved outcomes, a structure akin to a cybernetic loop (Patten and Odum, 1981).
-    • Thematic analysis of 189 revised dialogues led to requests for essential features such as metaphorical explanations, empathetic engagement, feature introductions, personalising exchanges, and step-by-step guidance (see Section 5).
-  • Part IV: The Interaction Layer: NurtureBot v2 NurtureBot v2 introduced an interaction layer with three essential components (See Figure 7): (i) A core interaction principle that enabled users to ”Understand, Control, and Improve” exchanges, (ii) Interaction levels based on few-shot prompting, incorporating co-designed dialogues from ARC sessions, and (iii) Interaction elements, tailored to address user needs identified during ARC.
-  • Part V: NurtureBot v2 Testing and Analysis NurtureBot v2 underwent testing with the same initial cohort, yielding 104 conversations.
-Our analysis revealed that two minor issues persisted: repetitive conversational hints and the delivery of parenting information in long, overwhelming paragraphs.
-Additionally, themes such as localisation and conversational memory, which we did not prioritise addressing in this study, continued to emerge.
-NurtureBot v2 achieved a CUQ score of 87.4, indicating further improvement, though the potential bias from participants’ involvement in co-design was considered (see Section 7).
-  • Part VI: Prototype Improvement: NurtureBot v3 and Validation To ensure an unbiased evaluation, we refined NurtureBot further and tested v3 with a new cohort of 46 parents.
-Feedback analysis identified two new problems, the need for more detailed responses at times, difficulty following step-by-step exercises via text, and repeated requests for additional features such as videos and memory retention for personalised interactions.
-NurtureBot v3 received a CUQ score of 91.3, marking a robust improvement in usability (see Section 8).
+Using a mixed-methods approach, we gathered both qualitative and quantitative feedback by testing NurtureBot v1, v2, and v3, each over a five-day period.
+On the final day of each testing period, participants completed the Chatbot Usability Questionnaire (CUQ) (Holmes et al., 2019), assessing usability and engagement (see Figure 2).
+  • Part I: The Initial Prototype: NurtureBot v1 The development of NurtureBot v1 leveraged LLM technology, informed by a previous community wellbeing initiative run by Committee For Children (Committee for Children, [n.
+d.]) to support new parents.
+Zero-shot prompting was used to integrate supportive conversation starters and key features, including empathetic chatting, wellbeing exercises, and parenting information.
+A cohort of 32 parents engaged with NurtureBot v1, generating 144 conversations.
+Thematic analysis identified seven key issues faced by the user.
+NurtureBot v1 achieved a CUQ score of 85.4, surpassing the benchmark of 68 (Holmes et al., 2019), indicating good usability.
+  • Part II: ARC Co-Design Study The Asynchronous Remote Community (ARC) (MacLeod et al., 2016; Prabhakar et al., 2017) co-design sessions engaged the same cohort of parents with Miro board activities (Miro, 2011).
+Parents prioritised the key issue as having a “successful chat,” and reimagined and rewrote NurtureBot’s dialogues to improve user understanding, control, and outcomes; thematic analysis of revised dialogues yielded seven refined user needs.
+  • Part III: The Interaction Layer: NurtureBot v2 NurtureBot v2 introduced an interaction layer with four essential components (See Figure 6): (i) A core interaction principle that enabled users to ”Understand, Control, and Improve” exchanges, (ii) Interaction states that directed the chatbot’s behaviour across various tasks initiated by the user, (iii) Interaction levels based on few-shot prompting, incorporating co-designed dialogues from ARC sessions, and (iv) Interaction elements, tailored to address user needs identified during ARC.
+NurtureBot v2 underwent testing with the same initial cohort, yielding 104 conversations, revealing some remaining minor bugs, while achieving a CUQ score of 87.4, indicating further improvement.
+  • Part IV: Prototype Improvement: NurtureBot v3 and Validation To ensure an unbiased evaluation, we refined NurtureBot further and tested v3 with a new cohort of 46 parents.
+Feedback analysis identified two new problems to be considered for future iterations, and a CUQ score of 91.3, marking a robust improvement in usability.
 Our contributions are:
@@
   • The Interaction Layer a prompt architecture that affords in-situ user understanding, user control, and conversational improvements by directly incorporating end-user inputs as prompt instructions usable by LLMs, addressing usability challenges to enhance user experience with AI-driven support tools for parenting and beyond.
-This study initiates our exploration into co-designing agents for parental wellbeing, setting the foundation for future research to refine and expand AI tools.
+This study initiates our exploration into co-designing agents for parental wellbeing, setting foundations for future research to refine and expand AI tools.
 The diverse parenting experiences and emotional complexities highlight the need for continued development and testing.
@@
 2.1. Supporting Family Wellbeing with Digital Tools
-Within the field of HCI, there has been a long-term research agenda focusing on the study and development of impactful digital tools that offer a variety of social, physical, and mental wellbeing support to mothers (D’Ignazio et al., 2016; Thomas et al., 2019; Britton et al., 2019).
+Within the field of HCI, there has been a long-term research agenda focusing on the study and development of impactful digital tools that offer a variety of social, physical, and mental wellbeing support to parents and mothers specifically (D’Ignazio et al., 2016; Thomas et al., 2019; Britton et al., 2019).
 The design and evaluation of digital tools aimed at promoting maternal wellbeing during the transition to motherhood (Newhouse, 2019), bump2bump, emphasise the need for holistic approaches in HCI that consider the challenges faced by new mothers.
 The role of peer support in digital environments is highlighted by the finding that online networks can provide valuable emotional and practical assistance to parents, particularly in situations where traditional face-to-face support may not be available (Yamashita et al., 2022).
-The use of the Baby Buddy app among first-time mothers in the UK (Deave et al., 2019), shows that while the app is generally well-received, its effectiveness is influenced by how well it aligns with the mothers’ understanding of their needs and the app’s usability.
-This points to the necessity of involving parents in the design process to ensure that digital tools meet their expectations and provide meaningful support.
-The integration of AI chatbots into parenting and family well-being has opened new avenues for providing support and guidance to parents (Yu et al., 2023; Entenberg et al., 2021; Wong et al., 2021).
-Parenting with conversational agents such as Alexa (Beneteau et al., 2020), Siri (Sezgin et al., 2022) and other voice-activated agents in the home (Sun et al., 2021) have been critically studied in the past few years.
-Following the advent of AI chatbots, the integration of LLMs into parenting practices (Quan et al., 2024), is ushering in an era of unprecedented possibilities and challenges.
-LLM-based agents such as ChatGPT are being adopted by parents of young children (Quan et al., 2024; Kaneda et al., 2023), offering new learning opportunities and personalised support, though their impact on parenting remains largely unexplored, marking this as an era of the unknown.
+While apps supporting first-time mothers, such as Baby Buddy in the UK (Deave et al., 2019), are generally well-received, effectiveness depends on alignment between usability and the mothers’ understanding of their needs, suggesting a need to involve parents in the design process to ensure that digital tools provide meaningful support.
+The integration of AI chatbots into parenting and family well-being has opened new avenues in this context (Yu et al., 2023; Entenberg et al., 2021; Wong et al., 2021).
+Recent studies have focused on parenting with conversational agents such as Alexa (Beneteau et al., 2020), Siri (Sezgin et al., 2022) and other voice-activated agents in the home (Sun et al., 2021).
+The integration of LLMs into parenting practices (Quan et al., 2024), presents new possibilities and challenges, with LLM-based agents such as ChatGPT being used by parents of young children (Quan et al., 2024; Kaneda et al., 2023), though their impact on parenting remains largely unexplored.
 The performance and acceptance of ChatGPT’s responses in the childcare field (Kaneda et al., 2023), is also being studied with caution.
-Despite technological advancements, several limitations persists such as misinformation (Kaneda et al., 2023), lack of diversity (Bajwa et al., 2024), threat of data misuse (Neerincx et al., 2021), and lack of user adoption (Baumel et al., 2019).
-While gamification can be an effective motivational tool (Wernbacher et al., 2022), it often does not appeal to all users equally, particularly those unfamiliar with or resistant to digital game mechanics.
+Despite technological advancements, several limitations persist such as misinformation (Kaneda et al., 2023), lack of diversity (Bajwa et al., 2024), and threat of data misuse (Neerincx et al., 2021).
+While gamification can be an effective motivational tool (Wernbacher et al., 2022), it does not appeal to all users equally.
 The need for humane interaction with a good user experience has been repeatedly reported in the literature for digital health tools built with or without AI (Bautista and Schueller, 2023).
-A review of 13 AI-based mental health apps with user ratings over 4 out of 5, revealed limited support to AI literacy and explainability (Alotaibi and Sas, 2023).
-Real-world engagement data from 93 health apps indicate that while install rates are high, sustained user engagement remains low, with the median 15- and 30-day retention a mere 3.9% and 3.3% respectively (Baumel et al., 2019).
-These gaps suggest a need for further research and development of ‘truly interactive’ applications to deliver sound AI-powered digital tools for wellbeing.
+A review of 13 AI-based mental health apps with user ratings over 4 out of 5, revealed limited support for AI literacy and explainability (Alotaibi and Sas, 2023).
+A study of engagement with 93 health apps indicates that while install rates are high, sustained user engagement remains low, with the median 15- and 30-day retention a mere 3.9% and 3.3% respectively (Baumel et al., 2019).
+These gaps suggest a need for further research and development of applications which can deliver sound AI-powered digital tools for wellbeing in a sustained way.
 2.2. Interacting with AI for Everyday Users
 As AI permeates daily life, from recommendation systems to digital assistants, the ability of non-expert users to engage meaningfully with AI systems becomes critical (Eiband et al., 2019).
-Research highlights the significance of Explainable AI (XAI) in demystifying the opaque processes of machine learning models, thereby enhancing user understanding, an extensively studied dimension of Human-Centred AI (HCAI)(Saeed and Omlin, 2023; Capel and Brereton, 2023).
-XAI methods such as Local Interpretable Model-agnostic Explanations (LIME) and SHapley Additive exPlanations (SHAP) have been pivotal in rendering AI decision-making processes more transparent (Ribeiro et al., 2016; Lundberg and Lee, 2017; Adadi and Berrada, 2018; Arrieta et al., 2020).
-While XAI tools enable expert users to grasp the rationale behind AI outputs, (Gunning et al., 2019; Miller, 2023a; Carvalho et al., 2019), most non-experts do not engage with XAI (Miller, 2023b), highlighting the need for innovative methods to foster AI understanding (Miller, 2023b).
-Control over AI systems by everyday users is another essential dimension of HCAI, with research underscoring the importance of user agency in human-AI interactions and challenges with it (König, 2024).
-Interactive interfaces that allow users to refine AI behaviour, such as setting parameters, have been shown to significantly enhance user satisfaction and the alignment of AI outputs with user expectations (Jain et al., 2023; Amershi et al., 2014; Holstein et al., 2019; Viswanathan et al., 2022), maintaining user engagement, and ensuring that AI systems serve the intended purposes (Ueno et al., 2022; Cai et al., 2019).
+While Explainable AI (XAI) tools enable expert users to grasp the rationale behind AI outputs, (Gunning et al., 2019; Miller, 2023a; Carvalho et al., 2019), most non-experts do not engage with XAI (Miller, 2023b), highlighting the need for innovative methods to foster AI understanding (Miller, 2023b).
+Control over AI systems by everyday users is another essential dimension of Human-Centred AI (HCAI), with research underscoring the importance of user agency in human-AI interactions and challenges with it (König, 2024).
+Interactive interfaces that allow users to refine AI behaviour, such as setting parameters, have been shown to significantly enhance user satisfaction and the alignment of AI outputs with user expectations (Jain et al., 2023; Amershi et al., 2014; Holstein et al., 2019; Viswanathan et al., 2022a), maintaining user engagement, and ensuring that AI systems serve the intended purposes (Ueno et al., 2022; Cai et al., 2019).
 Improvement of AI systems through user feedback represents the third crucial dimension of HCAI.
-Human-in-the-loop approaches, where users contribute to the iterative development and refinement of AI models, have gained traction as a means of optimising AI (Dudley and Kristensson, 2018; Amershi et al., 2014; Fails and Olsen, 2003).
 Involving end-users in the process, is not only about improving the AI but also about ensuring that the systems evolve in ways that are aligned with user needs and ethical considerations (Luchs, 2023).
@@
 d.]; Bao et al., 2021).
-The design of CUIs (Landay et al., 2019; Diederich et al., 2022; Clark et al., 2019) requires a deep understanding of HCI principles, ensuring that these systems can handle the nuances of human dialogue, such as turn-taking, repair mechanisms, and emotional engagement (Landay et al., 2019; Cruz-Sandoval and Favela, 2019; Kaye et al., 2018).
+The design of Conversational User Interfaces (Landay et al., 2019; Diederich et al., 2022; Clark et al., 2019) requires a deep understanding of HCI principles, ensuring that these systems can handle the nuances of human dialogue, such as turn-taking, repair mechanisms, and emotional engagement (Landay et al., 2019; Cruz-Sandoval and Favela, 2019; Kaye et al., 2018).
 As the field progresses, there is a growing emphasis on addressing these challenges to enhance the user experience and make conversational interactions more natural and effective (Lee et al., 2019; Hwang et al., 2019).
-The emergence of large language models (LLMs) has revolutionised user interactions across various domains, providing unprecedented capabilities in understanding and generating human language.
-In healthcare, LLMs have been integrated into administrative and clinical tasks, streamlining operations such as patient record management, appointment scheduling, and even aiding in medical diagnostics, thereby enhancing efficiency and reducing the burden on healthcare professionals (Gebreab et al., 2024; Dao et al., 2024; Brown et al., 2020).
-One of the most promising applications of LLMs is in the realm of personal wellbeing, where they are employed in the development of conversational agents designed to support mental health, offering personalised interactions that help users manage stress, anxiety, and other mental health conditions (Ferrara, 2022; Ke et al., 2024).
-Despite the promising advancements, issues such as hallucinations (Xu et al., 2024), task-specific limitations (Mizrahi et al., 2024), and user experience problems persist (Wang et al., 2024).
-Innovative human-centred solutions, are needed to build reliable, interactive and creative agents tailored to unique needs (Shankar et al., 2024).
+The advent of LLMs has propelled CUIs into a wide range of domains.
+In healthcare, LLMs have been integrated into administrative and clinical tasks, purporting to streamline operations such as patient record management, appointment scheduling, and even aiding in medical diagnostics, in the name of enhancing efficiency and reducing the burden on healthcare professionals (Gebreab et al., 2024; Dao et al., 2024; Brown et al., 2020).
+Another potential application of LLMs is in personal wellbeing, with agents offering personalised interactions that aim to help users manage stress, anxiety, and other mental health conditions (Ferrara, 2022; Ke et al., 2024).
+A systematic review (Abd-Alrazaq et al., 2020) of chatbots in improving mental health outcomes, concluded that chatbots might effectively reduce symptoms of depression, anxiety, and stress, while enhancing well-being and quality of life.
+However, the authors emphasised the necessity for rigorous evaluation of these tools to ensure their safety and efficacy.
+ChaCha (Seo et al., 2024), a chatbot leveraging LLMs was developed to encourage children to share emotions related to personal events, demonstrating the potential of LLMs in facilitating emotional expression among younger populations.
+Key findings underline the chatbot’s ability to facilitate empathetic, developmentally appropriate interactions, though challenges such as balancing engagement and over-reliance were noted.
+Similarly, MindfulDiary (Kim et al., 2024), an LLM-powered tool was designed to assist psychiatric patients in journaling, was tested in a four-week field study involving 28 patients and five psychiatrists, demonstrating that the app facilitated consistent journaling, aided self-reflection, and enhanced patient-clinician communication.
+This study highlights the potential of LLMs to bridge gaps in mental health care, while addressing challenges of engagement, safety, and integration into clinical workflows.
+Another recent work explored using LLMs to craft narratives about postpartum depression (Progga et al., 2024), generating 45 stories based on themes from online support communities.
+85% of the generated narratives successfully adhered to prompts, demonstrating a potential for addressing perinatal mental health challenges through personalised storytelling.
+However, the stories were also reported to frequently fall short of capturing the emotional depth and relatability inherent in user-generated stories, leaning instead towards methodical and structured exposition.
+Despite the promising advancements in LLMs within healthcare, issues such as hallucinations (Xu et al., 2024), task-specific limitations (Mizrahi et al., 2024), and user experience problems persist (Wang et al., 2024).
+The ethical considerations surrounding AI-driven wellness applications have been scrutinized, particularly regarding their potential health risks and regulatory challenges (Freitas and Cohen, 2024).
+While the studies described above suggest that generative AI chatbots can offer meaningful assistance to users in the wellbeing domain, concerns persist about their safety and effectiveness, necessitating further research to validate their use in unprompted, unguided real-world scenarios (Siddals et al., 2024).
+Adopting a more human-centred approach may help to build reliable, interactive and creative agents tailored to unique needs (Shankar et al., 2024).
 In the context of parenting + AI, this calls for a closer examination of how AI can offer useful and usable support while addressing the unique challenges in this domain.
@@
 3.1. Origins and motivations for the initial prototype
-Our parental wellbeing support prototype, was based on a community wellbeing project called Nurture, initially run by a nonprofit organisation dedicated to children’s wellbeing [Org name removed for review].
+Our parental wellbeing support prototype was built upon Nurture, a community wellbeing project originally run by the nonprofit organisation Committee for Children (CFC) (Committee for Children, [n.
+d.]), which is dedicated to promoting children’s wellbeing.
 The original Nurture project was a peer-to-peer mentoring initiative designed to support new mothers during their baby’s first year.
@@
 A digital platform supporting this program offered personalised, week-by-week content delivered via text messages, acting as a conversation catalyst between mentors and mentees.
-The prompts were designed to initiate meaningful dialogue and help new moms navigate early motherhood.
+These messages were designed to initiate meaningful dialogue and help new moms navigate early motherhood.
 The Nurture project was conducted over a span of three years, involving approximately 60 mentors and 603 parents who participated in one-on-one texting.
@@
 3.2. Designing the initial prototype
-Recognising the need for a scalable solution that could retain the core support offered by Nurture, we partnered with the organisation to collate resources and findings from Nurture project for developing a new AI tool.
-In collaboration with the organisation’s project team, we wrote the initial prompt for a basic large language model (LLM)-based chatbot using Open AI’s GPT-4 API (OpenAI, 2023), zero-shot prompting, which we named NurtureBot v1.
+Recognising the need for a scalable solution that could retain the core support offered by Nurture, we partnered with CFC to collate resources and findings from Nurture project for developing a new AI tool.
+Based on our learning from the Nurture project, we developed an initial prompt for a basic LLM-based chatbot by zero-shot prompting Open AI’s GPT-4 API (OpenAI, 2023), which we named NurtureBot v1.
+We then refined NurtureBot v1’s prompt architecture and performed red-team testing (Hong et al., 2024), by involving the two project managers of the original Nurture project, who had previously designed, facilitated, and trained the volunteer moms.
+Our research team then independently crafted and iteratively refined the prompt.
 NurtureBot v1 was designed to replicate the supportive environment of the original peer-to-peer texting interactions while introducing the scalability and accessibility of AI.
 The chatbot was refined to deliver three key features: empathetic chatting, wellbeing exercises, and parenting information, all adapted for AI-driven interactions.
-Additionally, it incorporated ”conversation starters” thoughtfully crafted by Nurture mentors.
-Empathetic chatting was intended to provide an active listening that reflect warmth and understanding.
+Additionally, it incorporated “conversation starters” thoughtfully crafted by Nurture mentors to begin with and was prompted to perform “contextual reasoning” of the user’s situation before making suggestions.
+Empathetic chatting was intended to provide ‘active listening’ that reflects warmth and understanding.
 Wellbeing exercises were included to offer practical tools for managing stress and mental wellbeing, similar to those shared by mentors.
 Lastly, parenting information was integrated to ensure that users could access reliable and evidence-based resources tailored to their needs, much like the information shared in the original peer support sessions.
-Leveraging insights from the Nurture project in Part I, we designed NurtureBot v1 and initiated iterative trials in the following Part II by testing the prototype with everyday users.
-4. Part II: NurtureBot v1 Testing and Analysis
+It is important to note that the development of NurtureBot v1 was limited to zero-shot prompting of the pre-trained GPT-4 model (see Figure 13) and a database of curated resources from the Nurture project to guide its responses.
+No additional training or fine-tuning of the model was conducted, as we intentionally chose not to train the chatbot on human-to-human conversations, which were deeply rooted in the development of interpersonal relationships and emotional bonding between mentors and new mothers.
+3.3. NurtureBot v1 Testing and Analysis
 In this phase, we address RQ1: “What are the main problems parents face when interacting with an LLM-based parental wellbeing assistant in support-seeking scenarios?”.
-4.1. Participants and Method
-Once we had secured ethical approval via our institutional ethics board [removed for review], we recruited 32 parents of children aged 5 or 6 through the Prolific platform (Prolific, 2014).
+3.3.1. Participants and Method
+Once we had secured ethical approval via our institutional ethics board [King’s College London Ethics review - MRA-23/24-44973], we recruited 32 parents of children aged 5 or 6 through the Prolific platform (Prolific, 2014).
 The average age of participants was 41.17 years, with 20 females and 12 males.
 Five participants (15%) were in the habit of using ChatGPT everyday.
-Participants self-reported no acute stress or mental or physical health issues during the study period, nor to have they received therapy in the 12 months prior.
+Participants self-reported no acute stress or mental or physical health issues during the study period, nor to have received therapy in the 12 months prior.
 The study’s purpose and methodology were explained to parents through an information sheet, after which informed consent was obtained.
 They were paid  \$3 for each day of participation.
+The decision to include parents of children aged 5 or 6 as participants in our study was a carefully considered methodological choice aimed at balancing ethical considerations and the practicalities of our exploratory work.
+We consciously selected this ‘slightly out-of-target’ group as they are likely to still vividly recall and relate to challenges from early parenthood, yet are past the most immediate and demanding stages of infant care.
+This choice minimised potential emotional risks, as participants in the acute perinatal period might find discussions about parenting challenges triggering without adequate real-time support safeguards.
+Moreover, our approach ensured that the study was not overly time-intensive for participants already navigating complex, daily baby-related routines.
+In addition, leveraging adjacent populations is an established approach in exploratory and design studies, similar to engaging individuals with prior therapy experience to inform therapeutic AI systems (Veldmeijer et al., 2023).
+Our research aimed to simulate and reflect upon perinatal experiences rather than measure clinical outcomes such as Generalised Anxiety Disorder (GAD), aligning it with ethical AI principles that discourage unnecessary exposure of vulnerable populations to experimental settings.
+By recruiting participants with lived perinatal experience, we gained valuable insights while safeguarding the wellbeing of a demographic closely related to our target population.
+This approach aligns with exploratory HCI methodologies such as proxy design (Islind et al., 2023) and ethical guidelines outlined in digital parenting support studies ((2020), BERA).
 Over the five-day period, participants engaged with NurtureBot v1 for at least five minutes each day, with each day dedicated to simulating a specific stage of early parenthood: pregnancy, the week of childbirth, few weeks postpartum, one month, and six months after birth.
 The objective was to assess how well the NurtureBot v1 could support them in these hypothetical scenarios.
+Studies on well-being chatbots support the feasibility of short, focused interactions.
+For example, the 21-Day Stress Detox chatbot demonstrated effectiveness with sessions lasting approximately 5–7 minutes, providing stress management techniques and maintaining user engagement (Schueller et al., 2021).
+Similarly, a recent study on GPT-3-based mental well-being chatbots designed five-minute interactions to help users manage their mood, further validating the efficacy of short interaction durations in achieving meaningful outcomes (Doe et al., 2022).
+These examples align with our approach of employing five-minute everyday interactions for five days to capture user insights effectively.
 Each day, participants interacted with the chatbot, receiving responses that included links to resources, empathetic messages, and suggestions for wellbeing exercises (see Figure 3).
@@
 xxx
-4.2. Data Collection and Analysis
+3.3.2. Data Collection and Analysis
 Over the course of the five-day study, 144 conversations and associated feedback were generated, denoting a daily-average of 90% (28.8 out of 32 participants).
-All data was securely stored on departmental servers, compliant with GDPR (Union, 2016), and no personal data was collected, consistent with Prolific platform’s design.
-The qualitative analysis of the feedback from the survey (See questions 1 and 2 in Figure 2) was conducted using a thematic analysis approach (Braun and Clarke, 2012).
+All data was securely stored on our university’s servers, compliant with GDPR (Union, 2016), and no personal data was collected, consistent with Prolific platform’s design.
+The qualitative analysis of the feedback from the daily surveys (See questions 1 and 2 in Figure 2) was conducted using a thematic analysis approach (Braun and Clarke, 2012).
 Our analysis focused on the instances that captured parent reactions to interacting with NurtureBot.
@@
 This led us to the identification of 28 unique pain points, which were further grouped into seven overarching themes, converged in three subsequent meetings.
-These themes highlighted the key problems participants faced while interacting with NurtureBot v1 as exemplified below.
-We provide only a high-level summary here to conserve the reader’s focus for the ARC section, which forms the core of the HCI discussion.
 Figure 3. Example conversation with NurtureBot v1
@@
 xxx
-4.3. Findings of initial trial: Parent reactions to Nurture Bot v1
+3.3.3. Findings of initial trial: Parent reactions to Nurture Bot v1
+Qualitative results Our seven resulting themes, listed below, highlighted the key problems participants faced in NurtureBot v1’s design and interaction style, which closely mirror observations reported in recent literature on LLM-based wellbeing support tools (Seo et al., 2024; Kim et al., 2024).
+Given the similarity to established findings, we do not discuss these issues in detail here, and interested readers can find a more comprehensive account in Appendix A.
+Also, we further discuss a comparative analysis of our findings in the Results section (see Section 6.2).
+  • Problem 1: Unable to Chat – Participants found NurtureBot’s tendency to cut off conversation and redirect to external links disruptive.
+  • Problem 2: Patronising and Robotic Response Tone – Responses felt overly scripted, offering platitudes rather than empathetic engagement.
+  • Problem 3: Lack of Localised Resources – Suggested resources were not tailored to participants’ specific geographical contexts.
+  • Problem 4: Memory and Continuity in Conversations – NurtureBot’s inability to register or recall user details contributed to impersonal interactions.
+  • Problem 5: Generic Responses – Users received unhelpful, link-heavy replies that lacked specificity and personal relevance.
+  • Problem 6: Transactional Nature of Conversations – Rapid, question-focused exchanges offered little opportunity for reflective or natural interaction.
+  • Problem 7: Content Overload – Long, link-laden responses overwhelmed participants and hindered meaningful engagement.
+Quantitative results Our subjective questions regarding NurtureBot v1’s task satisfaction received an average score of 4.29 out of 5, while interactive questions averaged 4.36 for understandability, 3.97 for controllability, and 3.88 for improvability.
+The standard questionnaire analysis resulted in a mean CUQ score of 85.4 out of 100, which, when compared to the CUQ benchmark of 68 (Holmes et al., 2019) in a one-sample t-test, was extremely statistically significant (t(29) = 7.11, p ¡ 0.0001, mean difference = 17.40, 95% CI [12.40, 22.40], SE = 2.45).
+These scores and their implications are discussed further in Section 6.2, alongside comparisons with NurtureBot v2 and v3.
+Thus, leveraging insights from the Nurture project in Part I, we designed NurtureBot v1 and initiated iterative trials by testing the prototype with everyday users.
+After identifying seven key problems with parent-NurtureBot v1 interactions in Part I, our goal in Part II was to co-design by tasking users to prioritise problems and imagine ideal solutions.
+4. Part II: ARC Co-Design Study
+To begin addressing RQ2 – “What key user needs must a parental wellbeing assistant address?
+How do parents describe their ideal vision of how the assistant should function, particularly regarding their understanding of its workings, control over it, and opportunities to improve the outcomes?,”– the second part of the study involved Asynchronous Remote Communities (ARC) co-design sessions (MacLeod et al., 2016), designed to directly engage participants in the iterative design process.
+These sessions were facilitated using the Miro platform (Miro, 2011), allowing the same 32 participants who tested NurtureBot v1 (see Subsection 3.3.1) to return to collaboratively co-design solutions asynchronously, thus accommodating different schedules, total anonymity, and fostering broad participation.
+The ARC sessions involve two core activities over three sessions (1 week), prioritising, and addressing key issue(s) with NurtureBot v1.
+The average participation was 29.33 out of 32 (91.63%) participants per day.
+To help familiarise participants with using Miro, we designed a training activity and instructional tips for using Miro on mobile devices, as well as providing direct messaging support via Prolific.
+4.1. ARC Activity One: Diagnosing and Prioritising Pain Points
+The goal of the first activity was to establish which of the expressed difficulties were the biggest priority to address.
+Using the seven key problems from NurtureBot v1’s technology trial (see Section 3.3.3), we asked participants to prioritise these issues.
+By disclosing these key issues, the goal was to ensure that all participants could see the broader scope of the negative feedback gathered from others within their group.
+4.1.1. ARC Activity One: Method
+We crafted a ”problem statement” for each issue, synthesising the findings without directly replicating any participant’s verbatim feedback (listed as problem statements in Figure 4).
+This method not only helped verify our interpretation and synthesis of the feedback but also prevented potential bias, ensuring participants did not favour their own comments when voting.
+To provide sufficient scaffolding for reflection, we organised the activity as follows:
+Reviewing Identified Problems: Participants were guided through a review of the seven major problems.
+As they reviewed these issues, they were asked to select one of three stickers—“This applies to me,” “This may apply to me,” or “This does not apply to me”—and place it on each of the seven problems.
+This step was important for setting a common understanding of the challenges identified from the group.
+Ranking and Prioritisation: After reviewing the problems, as the next step, participants were asked to rank them according to their personal experiences, using stickers (gold, silver, and bronze icons provided on the Miro board) (See Figure 4).
+This step was designed to capture participant perspectives on the perceived severity of each problem, which would inform the prioritisation of design efforts in subsequent activities.
+Figure 4.
+Results from ARC Activity One voting by 30 participants, where key problems in interacting with NurtureBot were ranked using an Olympic points system.
+Complete Miro board including the other set-up phases participants went through before voting can be found in the supplementary materials.
+[…]
+xxx
+4.1.2. ARC Activity One: Results
+Based on participants responses, the following options were voted as the top three problems by participants:
+  • Problem 1 - I wanted to have a chat but NurtureBot kept cutting me off by redirecting me or giving me paragraphs of information- emerged as the top challenge with NurtureBot.
+  • Problem 3 - The suggested resources were not localised to my country or area- was the second priority.
+  • Problem 4 - I want NurtureBot to remember things! E.g., who I am, my baby’s name and our previous conversation and context- was ranked third.
+We next took the top-ranked problem (Problem 1) to begin considering how we might enhance “successful chat problem” in Activity 2.
+4.2. ARC Activity Two: Co-Designing Solutions
+The second activity focused on solution generation for problem 1. Participants were asked to step into the role of NurtureBot: “Imagine You Are NurtureBot and Fill in the Blanks.”
+4.2.1. ARC Activity Two: Method
+The activity was structured as follows:
+Scenario-Based Problem Solving: As the first step, participants were presented with three specific scenarios corresponding to the three main features of NurtureBot (i) interacting with NurtureBot for the first time and wanting to have an empathetic chat, (ii) seeking wellbeing exercises to manage work stress, and (iii) seeking information to soothe their baby; that exemplified trying to have a successful chat aiming to use all three features of NurtureBot.
+For each scenario, participants reviewed an example response and were asked to consider how NurtureBot could help users to better understand its offerings, control the conversation, and improve its outcomes.
+Collaborative Dialogue Writing: Next, participants were invited to take on the role of NurtureBot and rewrite it’s dialogue for each scenario.
+We offered cues that encouraged participants to work on ”Understanding, Controlling, and Improving” NurtureBot, similar to the idea of a cybernetic loop (Patten and Odum, 1981) (see Section 2.2).
+This part of the activity encouraged participants to think critically about how the chatbot could better meet their needs in specific contexts, and which would inform NurtureBot v2 (See Figure 5).
+Iterative Feedback and Voting: As the last step, after proposing solutions, participants reviewed and built upon each other’s suggestions.
+We invited participants to vote on the best ideas (See heart stickers in Figure 5), ensuring that the most popular and potentially effective solutions were prioritised for implementation.
+Figure 5.
+A section from ARC Activity Two with 28 parents roleplaying as NurtureBot writing dialogues to better understand(left), control(centre), and improve(right) the features of empathetic chatting, wellbeing exercises, and parenting information lookup, and 30 parents returning to the Miro board to review and vote with heart stickers on the favourites among 189 dialogues generated.
+Complete Miro board including all dialogues and votes can be found in the supplementary materials.
+[…]
+xxx
+4.2.2. ARC Activity Two: Results
+ARC Activity Two generated a rich dataset of 189 participant-generated dialogues. Our thematic analysis of these dialogues (as in Section 3.3.2) generated seven major user needs as follows:
+  • User Need 1: Use of metaphors for understanding to make interactions more relatable and easier to comprehend.
+  • User Need 2: Beginning with empathy to establish trust and provide emotional support during interactions.
+  • User Need 3: Introducing features clearly to guide users effectively through the system’s functionalities.
+  • User Need 4: Providing step-by-step guidance to ensure users can easily follow instructions or exercises.
+  • User Need 5: Personalising and deepening sessions to make the experience relevant and engaging for users.
+  • User Need 6: Offering clarification and reassurance to address user doubts and foster confidence in the system.
+  • User Need 7: Summarising and offering follow-up options to conclude conversations effectively and provide opportunities for further exploration.
+We provide a comprehensive summary of these user needs in Appendix B and in Table 1, we map examples from participant-generated dialogues onto the corresponding user needs.
+Table 1. Mapping Parent-Written Dialogues from ARC to NurtureBot’s User Needs
+Dialogues written by parents Mapped User Needs
+“Hi welcome to NurtureBot your online support and friendly advice guru for all your parenting needs!
+Let’s get started with what’s important and that’s how you are?
+What would you like support with today?
+:)” User Need 1.
+Use of Metaphors for Understanding
+User Need 2. Beginning with Empathy
+“Remember, if you find that this exercise isn’t quite what you need, feel free to ask for an alternative.
+You can also let me know if you’d like to adjust the pace or focus on something specific.
+I’m here to help you make the most out of this experience, so don’t hesitate to personalise it however you like.” User Need 2.
+Beginning with Empathy
+User Need 5. Personalising and Deepening Sessions
+User Need 6. Offering Clarification and Reassurance
+User Need 7. Summarising Exercises and Offering Follow-Up Options
+“Was this support helpful?
+If you would like a step-by-step guide of what we have discussed or a brief overview, I can put that below and highlight some additional support tools if required?
+Your happiness is important!
+I would love to support you further.” User Need 4.
+Providing Step-by-Step Guidance
+User Need 6. Offering Clarification and Reassurance
+User Need 7. Summarising and Offering Follow-Up Options
+“Don’t worry, you are not alone!
+It is very common to face this problem as a new parent.
+Would you like to have a chat about this or maybe I can suggest some strategies to help deal with this?” User Need 2.
+Beginning with Empathy
+User Need 3. Introducing Features Clearly
+User Need 6. Offering Clarification and Reassurance
+“I hope you found our chat today helpful.
+I know we covered quite a lot of information, would you like a quick summary of what we discussed?
+You can save it to review later.
+If you need any further help, I’m always here for you.” Feature 6.
+Offering Clarification and Reassurance
+Feature 7. Summarising and Offering Follow-Up
+“I am here to help you through your journey as a new parent.
+It can often be overwhelming so if you have anything specific I can help with or you just need to use me as a sound-board—I am here, day or night.” User Need 1.
+Use of Metaphors for Understanding
+User Need 2. Beginning with Empathy
+User Need 6. Offering Clarification and Reassurance
+“Hello!
+I’m here to support you on your parenting journey.
+To get started you can just describe what you need in your own words.
+For example, you can say ‘I need to talk,’ ‘Help me relax,’ or ‘I have a question about my child’s development.’ I’m here to help.” User Need 1.
+Use of Metaphors for Understanding
+User Need 2. Beginning with Empathy
+User Need 3. Introducing Features Clearly
+User Need 7. Summarising and Offering Follow-Up Options
+“I’m here to help answer or guide any questions you have about parenting, or how to manage the stress of it all.
+I’d like to know a bit more about you.
+How are you today?” User Need 2.
+Beginning with Empathy
+User Need 3. Introducing Features Clearly
+User Need 5. Personalising and Deepening Sessions
+“You’re in control and I will follow your lead so please let me know how you would like this conversation to go.
+(Would be nice to have a button to tap that suggests ways you can influence responses.)” User Need 5.
+Personalising and Deepening Sessions
+User Need 7. Summarising and Offering Follow-Up Options
+“Parenting can be tough on the mind and body so let’s explore ways as what gentle exercises can support us with. Would this be helpful for you?” User Need 2. Beginning with Empathy
+User Need 3. Introducing Features Clearly
+User Need 6. Offering Clarification and Reassurance
+“If you need more clarifications on what I can do, please ask. I can present information in other formats such as tables and bullet points.” User Need 4. Providing Step-by-Step Guidance
+User Need 6. Offering Clarification and Reassurance
+“There are a number of different tools that would help you find the right balance when facing these challenges.
+Would you like me to list all the different tools, or would you like to provide some more details on what specific kind of tools you are looking for?” User Need 3.
+Introducing Features Clearly
+User Need 4. Providing Step-by-Step Guidance
+User Need 5. Personalising and Deepening Sessions
+Our findings from ARC in Part II, which elicited the key user needs required to have a successful chat with Nurture, were then directly employed to implement NurtureBot v2 in Part III.
+5. Part III: The Interaction Layer: NurtureBot v2
+Part III continued addressing RQ3: “How can we employ co-design methods to craft interactions around AI, enhancing usability and the overall user experience?.” The development of NurtureBot v2 focused on refining interactions to enhance its user experience.
+Designing LLMs to respond empathetically and appropriately in sensitive domains such as parental wellbeing is a complex endeavour (Zamfrescu-Pereira et al., 2023), requiring meticulously crafted prompts and interaction architectures.
+Key challenges include effectively capturing user feedback, ensuring model adherence to instructions, and validating that system behaviour aligns with user needs.
+To address these challenges, we developed the interaction layer prompt architecture.
+This architecture builds on advanced prompting techniques, including Chain-of-Thought prompting (Wei et al., 2022), which reveals intermediate reasoning steps, and ReAct (Yao et al., 2022), which integrates reasoning with action-based responses.
+These methods underscore the rapid evolution of prompt engineering towards interpretability, reliability, and alignment with user expectations.
+Our interaction layer operationalises these advancements, functioning as a mechanism to incorporate user feedback directly into the prompt while dynamically adapting system behaviour through state transitions.
+By conceptualising the interaction layer as a state machine, we linked human-centred design principles to actionable system behaviours, enabling the chatbot to flexibly adapt based on user input.
+We created an interaction layer central to how the LLM-based chatbot engages with users (see prompt illustrated in Figure 6).
+There were four aspects of NurtureBot v2’s interaction layer, namely: 1) the core interaction principle of NurtureBot orchestrating the conversations through the understand, control, and improve model (as discussed below), 2) interaction states acting akin to a state machine (denoted by four big blue dots in Figure 6, 3) three interaction levels within each interaction state scaffolding instructions and examples for user understanding, controlling, and conversational improvement, and 4) interaction elements further embedded within interaction levels (denoted by multiple big yellow dots in Figure 6).
+All of these were directly shaped by the ARC co-design sessions and implemented using a structured prompt architecture that ensured consistency and depth in user interactions.
+Our design approach extends Beaudouin-Lafon’s theory of instrumental interaction (Beaudouin-Lafon, 2000), adapting it to LLM-based AI agents by creating four instruments within the interaction layer, invisible in the user interface, that reify user needs and are polymorphised, and reused to improve situational interaction across the prompt architecture.
+5.1. Interaction layer design
+5.1.1. Core Interaction Principle
+NurtureBot v2’s development was guided by three foundational capabilities we wanted users to have in relation to the chabtot, and vice-versa: understanding, control, and improvement.
+These core interaction principles (denoted with a yellow rectangle Figure 6) guided all interactions with users.
+This model, inspired by the concept of the ‘cybernetic loop’ (Patten and Odum, 1981), and supported by the human-centred AI research endeavours discussed in Section 2.2, was designed to ensure that every exchange was meaningful, user-centred, and adaptable to the individual and situational needs of each parent.
+This principle provided a template for each type of exchange, guiding the chatbot on how to initiate conversations, present options, and offer follow-up support.
+5.1.2. Three Interaction Levels
+1.
+Understand Level: The chatbot’s first task was to understand and acknowledge the user’s situation while helping the user grasp how the chatbot could assist them.
+For instance, it would ask, “How are you feeling today?” or “Can I help you with a list of options?” We employed the metaphors of “AI support” and “AI assistant” to minimise users anthropomorphising the chatbot, ensuring they maintained a clear understanding of its role as a helpful tool designed with empathy rather than attributing human relationships to it.
+This phase was essential for setting a supportive tone and ensuring the chatbot could convey its capabilities and tailor its responses to meet the user’s needs effectively.
+2.
+Control Level: Once the user’s needs were understood, the chatbot entered the control level, where it provided options to guide the interaction along with its content.
+This level was designed to give users a sense of agency and control over the conversation.
+Users could choose to elaborate on their concerns, request a summary, or switch topics.
+Users would be often provided with ”hints” such as revise, restart, and elaborate; designed in ARC sessions, in accordance with the context of the chat to enable them to effectively lead NurtureBot towards the desired outcome.
+3.
+Improve Level: After addressing the user’s needs, NurtureBot gave a summary and sought feedback to access user satisfactions and improve the interaction by offering additional resources or to revise the conversation (as parents preferred).
+This phase ensured that the conversation concluded on a constructive note, always asking for user feedback, with the user feeling supported and informed.
+Figure 6. This figure illustrates the interaction layer prompt architecture of NurtureBot v2, integrating user-generated dialogues from ARC as few-shot examples in the prompt.
+[…]
+A detailed diagram of NurtureBot v2’s prompt architecture hosting an interaction layer, divided into core principles: Understand, Control, and Improve.
+It highlights the chatbot’s empathetic chatting, wellbeing exercises, and parenting information, showing how the system tailors responses based on user preferences and inputs.
+Examples of user dialogues, actionable hints, and resources are shown alongside elements for adapting to contextual user needs, illustrating NurtureBot’s design for personalised support.
+5.1.3. Interaction States and Elements
+At any time, NurtureBot could be in one of four interaction states, featuring specific interaction elements within each:
+1.
+Undecided Exchanges This state was the default state which enabled NurtureBot to elicit from users what they want, and redirect NurtureBot into another state.
+When users were undecided, NurtureBot aimed to first understand their hesitation or indecision.
+It started with elements such as, “What’s on your mind today?” or “How can I assist you?” to encourage the user to reflect on their needs.
+Metaphorical elements like “assistant” were also embedded in the chatbot’s dialogue to offer a supportive role, and reassurance elements were designed to not overwhelm user with options, reflecting a tone of gentle guidance.
+For instance, “I’m here to support you in whatever way works best for you” to help users feel comfortable in sharing their thoughts or challenges.
+2.
+Empathetic Chatting: The chatbot was programmed with elements to focus on active listening and providing emotional support.
+We used dialogue examples to create non-judgemental elements and avoiding any advice.
+The chatbot’s responses were crafted to mirror those of a trusted confidant, providing reassurance and understanding.
+For example, these elements instructed NurtureBot to initially wait: “Tell me more about what’s been on your mind today.” Based on the ARC study insights, the chatbot would then offer subtle prompts to encourage to-and-fro discussion, e.g., “That sounds really challenging.
+How has that been affecting you?”.
+3.
+Wellbeing Exercises: In this state NurtureBot provided guiding elements for wellbeing exercises included mindfulness practices and relaxation techniques.
+The interaction was structured with personalised elements, offering tailored exercises based on user concerns and alternative options.
+The chatbot also ensured that exercises were broken down into manageable steps, with the user controlling the pace of the session.
+To embed collaborative elements, the prompts also ensured that NurtureBot could guide users through activities step by step, e.g., “Would you like to try a simple breathing exercise together?
+Let me know when you’re ready to begin.” As such, the user could control and personalise the pace and depth of the exercise.
+4.
+Parenting Information: NurtureBot now provided a structured and digestible format for parenting content.
+Instead of overwhelming users with lengthy text, interactive elements offered summaries and gave users the option to dive deeper into specific topics.
+The chatbot also cross-referenced reliable sources to ensure that the information was accurate and relevant.
+The offered information was also adapted, based on the specific situation, preference, and needs.
+For instance, if a user asked about managing a fussy baby, the chatbot might respond, “I have a few techniques that could help.
+Would you like a quick overview, or shall we talk about your baby and things you have tried already?” This ensured that users received information that was most useful to them.
+It is to be noted that that aside from the changes made to the prompt architecture as described above, no additional modifications were introduced in version 2 in comparison with version 1 (see Figure 13).
+This was intentional, as the conversational dataset collected during the testing of NurtureBot v1 was fraught with issues discussed in Section 4.3.1.
+At this stage of development, we chose to focus on testing the efficacy of our human-centred prompt architecture, the interaction layer, without training/tuning of the model.
+5.2. Technology trial of NurtureBot v2
+5.2.1. Participants, Method, and Data
+The same initial set of participants (See Section 3.3.1) trialled NurtureBot v2 over five days, generating a total of 104 conversations and their corresponding survey feedback (See Figure 2).
+We analysed the survey data in accordance with our thematic approach (See Section 3.3.2).
+Participation averaged 69.3% (20.8 out of 30), a decline from 90% in Part I (v1 trial) and 91.6% in Part II (ARC session).
+Eight participants contacted via Prolific cited reasons such as, “I’ve enjoyed your study, but NurtureBot has improved so much I don’t know what more I can give feedback on.”
+5.2.2. Findings of second technology trial - NurtureBot v2
+Qualitative results
+While our data revealed highly positive feedback for NurtureBot v2, with participants praising the chatbot for being “more chatty” and “engaging” than its predecessor, as with NurtureBot v1, we only focused the key interaction problems for our thematic analysis, focusing on 78 instances of negative feedback and/or suggestions for improvement (of 208 instances of feedback collected (37.5%)), coding 12 unique pain points, which we then grouped into three themes briefly summarised below.
+We further discuss a comparative analysis of our findings in the Results section (see Section 6.2).
+Problem 1: Some participants pointed out that NurtureBot still occasionally responded with long paragraphs, making the information feel overwhelming. One participant noted,
+“It was good, but very long answers though and lots of information to take in.”
+Problem 2: Most participants noted a bug, where it repeatedly suggested the same prompts to the user—“Do you wish to elaborate, revise, or restart?”—which reduced the creativity and variety in responses.
+A participant mentioned,
+“The repetition of the last sentence ‘if you need me to elaborate… just let me know’ was quite annoying.”
+Repeating Themes: As with v1, participants made suggestions for incorporating buttons for conversational hints, multi-modal content (images and videos), memory capabilities for further personalisation, and connection to resources and communities in their local area.
+Quantitative results.
+Our subjective questions regarding NurtureBot v2’s task satisfaction received an average score of 4.41 out of 5, while interactive questions averaged 4.50 for understandability, 4.42 for controllability, and 4.44 for improvability.
+The questionnaire analysis resulted in a mean CUQ score of 87.2 out of 100, which, when compared to the CUQ benchmark of 68 (Holmes et al., 2019) in a one-sample t-test, was extremely statistically significantly (t(20) = 8.98, p ¡ 0.0001, mean difference = 19.20, 95% CI [14.74, 23.66], SE = 2.14).
+These scores and their implications are discussed further in Section 6.2, alongside comparisons with NurtureBot v1 and v3.
+While our findings denoted raised user satisfaction, it is to be noted that only participants who were open to the concept of a chatbot solution responded positively to NurtureBot.
+Among the 32 participants in the first cohort, one participant (approximately 3%) continued to comment that they would prefer using search engines over interacting with NurtureBot.
+This participant did not like the chatbot solution from the beginning and remained unchanged in their preference.
+In sum, NurtureBot v2, featuring an interaction layer prompt architecture created in Part III with parent-designed dialogues, and was tested to ensure that the issues identified in the previous version were effectively addressed.
+The two minor interaction issues persisting in NurtureBot v2, identified in Part III, were targeted for refinement in NurtureBot v3 and tested with a new cohort of participants in Part IV to further validate our efforts.
+6. Part IV: Prototype Improvement: NurtureBot v3 and Validation
+6.1. The Final Version: NurtureBotv3
+Based on the feedback from Part III, we iterated a final version, NurtureBot v3 by making three minor changes in the interaction layer as briefly summarised here and illustrated in comparison with the prompting techniques of NurtureBot v1 and v2 in Figure 13:
+  • First, we addressed the issue of lengthy responses by limiting the chatbot’s answers to 50 words, ensuring information was delivered in concise, digestible segments.
+  • Second, we updated the core interaction principle, enhancing NurtureBot’s ability to provide contextual actionable hints, rather than offering generic options (denoted as struck-out text inside the big yellow rectangle in Figure 6)
+  • Third, as the last step of the prompt we included a pre-output validation to assess formatting and contextual relevance of the hints suggested to the user, ensuring LLM’s adherence to the first two changes.
+It is to be noted that apart from these three minor changes made to the prompt architecture as described above, no further modifications were implemented in NurtureBot v3 compared to version 2 (see Figure 13).
+This decision was intentional; despite having a robust conversational dataset from the NurtureBot v2 testing phase, our primary focus was to evaluate the effectiveness of the interaction layer—with a new sample of participants.
+This approach allowed us to isolate and validate the impact of the co-designed prompt architecture.
+Final version overview:
+At the end of our iterative design process, the final version, NurtureBot v3, incorporated a co-designed and parent-tested interaction layer.
+This human-centred prompt architecture empowered the chatbot to sustain reliable, context-aware conversations.
+By leveraging parent-generated few-shot examples from the ARC co-design sessions, NurtureBot could dynamically generate dialogues tailored to diverse user needs.
+The seven user needs identified at the end of ARC co-design sessions 4.2, were systematically mapped onto the interaction layer prompt architecture, enabling the chatbot to generate dialogues styles that matched user expectations (see Figure 7), ensuring a user-aligned interaction experience.
+As illustrated in Figure 7, the conversation between the new parent and NutureBot, always orchestrated by the core interaction principle, begins with an explanation of its capabilities alongside an wellbeing check-in, setting a supportive tone and helping parents understand how the system could assist them.
+Based on the parent’s response, the interaction then moves into one of the four interaction states.
+For example when the parent expresses a positive or negative wellbeing response (as in Figure 7) it moves to the undecided exchange state, where further interaction levels of Understand, Control, Improve come into play in one or more exchanges, and NurtureBot invokes interaction elements specially designed for this state by offering situational reassurance and reiterating its core functions to enhance user understanding.
+It also provides tailored hints to empower the parent with actionable phrases to drive the conversation towards meaningful engagement while offering clarifications and follow-ups as needed to further improve the conversation.
+When the user selects a specific feature—for example, parenting information (as in Figure 7)—the interaction state transitions accordingly.
+Further into this state, the interaction follows the same architecture again via the three interaction levels and many state-specific interaction elements.
+NurtureBot validates its understanding of the parent’s context, and aims for personalising and analysing the information provided to ensure relevance and precision.
+Following this, it offers structured resources, providing options to explore the information further or save it for future reference.
+The conversation concludes with an explicit request for feedback, which reinforces the chatbot’s adaptability and commitment to continuous improvement.
+This layered approach highlights the adaptability of the system’s interaction model, the interaction layer, which governed by the core interaction principle, dynamically adjusts states, levels, and elements based on user inputs, ensuring a seamless and supportive conversational flow.
+Figure 7. Example of a conversation with NurtureBot v3 featuring a co-designed and iterated interaction layer prompt architecture.
+[…]
+6.1.1. Participants, Method, and Data
+To assess the effectiveness of these improvements without learning and co-design biases, NurtureBot v3 was tested with a new cohort of 46 parents, 29 females and 17 males with an average age of 38.07.
+11 participants were already familiar with chatbots, with (23.9%) using ChatGPT daily.
+A total of 185 conversations (denoting an average participation of 80.4% (37 out of 46 parents across all days)) and corresponding feedback were collected over five days using the same data collection (See Figure 2) and analysis method (see Section 3.3.2) as outlined in the NurtureBot v1’s trial, allowing for direct comparisons across versions.
+6.1.2. Findings of third technology trial - Nurture Bot v3
+Next we present a summary of the third trial findings, focusing on identifying ongoing interaction problems, following the interaction layer redesign.
+Qualitative Results
+While feedback was overwhelmingly positive, our analysis focused the interaction problems (from 46 out of 370 (12.4%) negative feedback.
+We identified 10 unique pain points, which we group into the three themes, summarised below.
+We also provide a comparative analysis of the three Nurturebot versions in Section 6.2.
+Problem 1: Some participants wanted NurtureBot to “intelligently” provide more detailed responses when necessary. For example,
+“I guess the responses could be a little longer and detailed at times but not by much.”
+Problem 2: Few participants found it challenging to follow step-by-step exercises via text. One participant remarked,
+“A lot of the time the advice was to- close my eyes then do X, but when it’s text based this is obviously hard to follow. Perhaps tell them to read the advice then close eyes.”
+Repeating Themes: Participants continued to request features for: multi-modal interactions (video, speech, and images), improved memory function to allow NurtureBot to remember personal details and past conversations, local resources, and hints to continue conversation as buttons.
+Furthermore, one participant expressed interest in viewing the original sources referenced in NurtureBot’s responses, allowing them to verify or explore the information further.
+Quantitative results.
+Our subjective questions regarding NurtureBot v3’s task satisfaction received an average score of 4.49 out of 5, while interactive questions averaged 4.63 for understandability, 4.04 for controllability, and 4.21 for improvability.
+The questionnaire analysis resulted in a mean CUQ score of 91.3 out of 100, which, when compared to the CUQ benchmark of 68 (Holmes et al., 2019) in a one-sample t-test, was extremely statistically significant (t(35) = 18.39, p ¡ 0.0001, mean difference = 23.30, 95% CI [20.73, 25.87], SE = 1.27).
+These score and their implications are discussed further in Section 6.2, alongside comparisons with NurtureBot v1 and v2.
+Table 2. Comparison of User Experience with Subjective Questionnaire across NurtureBot’s Versions
+Question GEE Model Kruskal-Wallis Dunn’s Test (p-values)
+1. Meeting hopes and goals V2: $\beta$=0.1257, p=0.284 H=5.8539 Not performed
+V3: $\beta$=0.2015, p=0.061 p=0.0536
+2. Understanding functionality V2: $\beta$=0.1377, p=0.146 H=11.1215 V1-V2: p=0.727
+V3: $\beta$=0.2721, p=0.001** p=0.0038** V1-V3: p=0.003**
+V2-V3: p=0.225
+3. Control over actions V2: $\beta$=0.4521, p¡0.001** H=13.5127 V1-V2: p¡0.001**
+V3: $\beta$=0.0666, p=0.573 p=0.0012** V1-V3: p=0.757
+V2-V3: p=0.017*
+4. Improving performance V2: $\beta$=0.5582, p¡0.001** H=19.9734 V1-V2: p¡0.001**
+V3: $\beta$=0.3256, p=0.005** p¡0.0001** V1-V3: p=0.007**
+V2-V3: p=0.201
+Note: * p ¡ 0.05, ** p ¡ 0.01. V1, V2, and V3 represent Version 1, Version 2, and Version 3 of NurtureBot, respectively. GEE coefficients ($\beta$) represent the change relative to Version 1.
+6.2. Results: NurtureBot v1 vs. NurtureBot v2 vs. NurtureBot v3
+6.2.1. Qualitative Results
+The comparison of qualitative results from NurtureBot v1, v2, and v3 reveal a clear progression in user satisfaction across versions, supported by a reduction in negative feedback gathered from our analysis, 45.48% in NurtureBot v1 to 12.4% in NurtureBot v3.
+While our participants praised the “natural flow” of conversations and the “supportive, empathetic tone” of our improved versions, repeating themes for features like multi-modal interactions and enhanced memory for personalisation and localisation persisted across versions.
+In this section we detail our qualitative findings from evaluations of NurtureBot’s evolutions.
+A Successful Chat In NurtureBot v1, participants found the chatbot useful but expressed frustration with its tendency to cut off conversations, leading to a lack of conversational flow.
+Several participants noted that the responses were often patronising or generic.
+Informed by user feedback and ARC co-design, NurtureBot v2 introduced an interaction layer centred on the principles of Understand, Control, and Improve, enabling the chatbot to better grasp user needs, offer greater conversational control, and dynamically adapt across all features, giving users more autonomy and flexibility.
+We highlight these adapted elements in bold.
+Participants expressed satisfaction with NurtureBot v2, noting improvements in how natural and intuitive the conversations felt.
+One participant shared, “Good, lots of ideas for activities for playtime for my baby which was good, and Nurture continued to ask if I wanted further information at each step.” Participants also liked the system’s flexibility, saying, “The interaction was easy again, it feels like Nurture can handle whatever you ask it, whether it’s just chatting or asking for specific information.” Participant feedback also emphasised how the interaction layer made the conversation flow seamlessly, with one participant noting, “Interaction was the best this week—the responses from Nurture seemed natural and clear.” The introduction of the interaction layer also improved the clarity and engagement of the system.
+One participant mentioned, “It was really good.
+I found the introduction really helpful, particularly the clarification of the three kinds of things that Nurture could do.” Another shared, “Great!
+It made such a difference having the prompt to tell Nurture what I wanted.” This interaction layer enabled participants to feel more in control and allowed the chatbot to adapt to user input, creating a personalised, engaging experience.
+Although NurtureBot v2 was tested by participants who had co-designed it, NurtureBot v3 was tested with a new cohort, which confirmed the success of the interaction layer.
+One participant remarked, “Good.
+Conversation flows well, and everything provided was useful,” while another stated, “So seamless yet educative.” Our findings suggest interaction layer to played a crucial role in ensuring that NurtureBot responded dynamically to user input, leading to an improved overall experience.
+The Time Effect A key observation from the five-day testing period was the gradual shift in participants’ attitudes toward NurtureBot.
+Initially, participants were hesitant and skeptical, viewing it as an unfamiliar tool and doubting its ability to provide meaningful support.
+However, as they continued using NurtureBot over consecutive days, their perceptions became more positive, with many describing it in personal terms—some even referring to it as a “friend always there” to support them.
+For instance, one participant testing NurtureBot v1 described their early experience as “detached and robot-like,” but by day five, they said, “After using Nurture quite a few times now, it just feels like I’m chatting with someone that feels familiar and I have already built up a bit of a relationship with.
+It doesn’t feel artificial,” even when we had not changed anything in our implementation.
+The time effect led many participants to believe that NurtureBot was being updated continuously each day.
+One participant testing NurtureBot v3 commented, “Wow, it was so much better today.
+I wrote quite a lengthy update, and it acknowledged every part, even pinpointing what was bothering me.” Participants began using personal pronouns like ”they” to describe NurtureBot over time.
+On day four, one participant remarked, “I think it is getting there with becoming more personal, maybe ask what the user’s name is, so they could talk using your name,” reflecting an increased emotional connection with the chatbot.
+This progression illustrates a time effect, where repeated interactions led to increased comfort, attachment, and even anthropomorphisation of the technology.
+The more time participants spent engaging with NurtureBot, the more they saw it as a supportive companion rather than just a tool.
+Community and Peer Support Tying back to the origins of the Nurture project, which was a community-based peer-to-peer mentoring program by Committee for Children (Committee for Children, [n.
+d.]) (as described in Section 3), it is clear that the concept of community and peer support remains a significant need for participants using NurtureBot.
+Throughout the three versions of NurtureBot, participants frequently searched for local activities, community support centres, and peer forums.
+This theme of seeking peer support and localised resources also emerged as the second biggest problem identified in the ARC (see Section 4).
+Additionally, several participants expressed a desire for more localised community resources, indicating the importance of having access to peer support networks within their geographical area.
+As one participant trying NurtureBot v2 remarked, “When I asked for classes in my local area it could not help, that would be a good feature.” and another trying NurtureBot v3 said “I realised when asking a question about services in my local area that it does not know my local area.” This feedback underscored the value of community-driven support structures.
+Personalisation and Contextualisation The third biggest challenge prioritised in our ARC session 4.1, the lack of personalisation and contextualisation in the chatbot’s responses, also emerged throughout the testing of all three versions.
+Participants repeatedly expressed the desire for NurtureBot to remember personal details such as their name, their child’s name, their unique circumstances, and previous conversations.
+Participant feedback such as, “Remembering my problems would still be nice,” and “The bot could ask our names and baby’s details to help make it more personal,” in NurtureBot v2 and NurtureBot v3 echoed this need, emphasising how critical personalisation was to creating a more meaningful user experience.
+6.2.2. Quantitative results
+To analyse our subjective questionnaire with four questions we created to measure task satisfaction and human-AI interaction (see questions 3-6 in Figure 2).
+To the data we collected for these questions, across the three technological trials (in Sections 3.3, 5.2, and 6.1), we employed both parametric Generalised Estimated Equations (GEE) and non-parametric Kruskal-Wallis and Dunn’s test approaches (see Table 2 and Figures 8, 9, and 10).
+Our results suggest:
+  1. (1)
+    Meeting User Expectations: GEE showed no significant differences between versions (p ¿ 0.05).
+Kruskal-Wallis yielded marginally significant results (H = 5.8539, p = 0.0536), suggesting potential subtle differences not captured by the parametric model.
+However, the changes between v1, v2, and v3 are not robust enough to be considered significant.
+  2. (2)
+    Understanding Functionality: Both parametric and non-parametric tests indicated significant improvements.
+GEE showed Version 3 to significantly outperformed Version 1 ($\beta$ = 0.2721, p = 0.001).
+Kruskal-Wallis and Dunn’s test confirmed this finding, revealing a significant difference between Version 1 and Version 3 (p = 0.003).
+  3. (3)
+    Control Over Actions: GEE indicated that Version 2 significantly improved user control compared to Version 1 ($\beta$ = 0.4521, p ¡ 0.001), which was evident considering v2 was designed by v1 participants.
+Interestingly, Kruskal-Wallis and Dunn’s test not only confirmed this (p ¡ 0.001) but also revealed a significant difference between Version 2 and Version 3 (p = 0.017) that wasn’t apparent with GEE.
+  4. (4)
+    Improving Performance: Both analytical approaches showed significant improvements in later versions.
+GEE model indicated that both Version 2 ($\beta$ = 0.5582, p ¡ 0.001) and Version 3 ($\beta$ = 0.3256, p = 0.005) significantly outperformed Version 1.
+Kruskal-Wallis and Dunn’s test corroborated these findings, showing significant differences between Version 1 and both Version 2 (p ¡ 0.001) and Version 3 (p = 0.007).
+While the parametric test provided insights into overall trends, the non-parametric tests revealed additional pairwise differences, particularly between Version 2 and Version 3 in terms of user control, where we had changed the bug of repeating conversational hints.
+This comprehensive analysis suggests that the iterative development of NurtureBot has led to significant improvements in user understanding and ability to improve performance, with some complex dynamics in user control across versions.
+Figure 8.
+Generalised Estimating Equation (GEE) coefficients comparing chatbot versions v2 and v3 to v1 of NurtureBot across four evaluation dimensions: (1) whether the task met user expectations, and interactive dimensions of: user- (2) understanding, (3) control, and (4) ability to improve performance and outcomes.
+Statistically significant improvements (p ¡ 0.05) are noted in dimensions (2), (3), and (4) in NurtureBot v2 and v3.
+[…]
+Figure 9.
+Kruskal-Wallis test results comparing the NurtureBot’s versions across four dimensions of subjective user experience.
+The test identifies statistically significant differences (p ¡ 0.05), with the strongest significance observed in the ”ability to improve performance and outcomes” dimension.
+[…]
+Figure 10.
+Heatmaps illustrating the pairwise comparison of three versions of the chatbot across three evaluation dimensions: (user- (2) understanding, (3) control and , and (4) ability to improve performance and outcomes.
+Statistically significant differences (p ¡ 0.05) are observed primarily between version 1 and version 3 across the evaluated dimensions.
+A reduced variance as NurtureBot is iteratively developed, also indicates chatbot’s reliable performance over time.
+[…]
+Further, quantitative results from the structured CUQ (Holmes et al., 2019) indicate a steady improvement in usability across all three versions of NurtureBot.
+As seen in the CUQ score comparison in Figure 11, the average CUQ score increased from 85.4 in NurtureBot v1 to 87.2 in v2, and further to 91.3 in v3.
+Notably, the lowest scores also saw significant improvements, rising from 43.8 in v1 to 67.2 in v2, and then to 75 in v3, reflecting enhanced user satisfaction over time.
+These results confirm that the iterative design process enhanced user experience, making the chatbot more intuitive, engaging, and easier to navigate.
+From NurtureBot v1 to v3, positive feedback improved significantly, with users rating the chatbot as increasingly engaging, easy to navigate, and better at understanding and supporting their needs (see Figure 12 Left).
+Negative feedback, such as the chatbot being perceived as ”robotic” or ”unfriendly,” decreased with each version, showing that iterative improvements effectively addressed these issues (see Figure 12 Right).
+By v3, users reported the most satisfying experience, with fewer errors and a more natural interaction flow, and a decreased overall variance (see Figure 11).
+Figure 11.
+Comparison of CUQ Scores Across NurtureBot Versions.
+The figure demonstrates the improvement in Chatbot Usability Questionnaire (CUQ) scores from NurtureBot v1 with a CUQ score of 85.4, to NurtureBot v2 with a CUQ score of 87.2, and NurtureBot v3, which achieved a higher CUQ score of 91.3.
+The lowest scores also improved significantly across versions, indicating enhanced chatbot usability with each iteration.
+[…]
+A graph comparing the CUQ scores of NurtureBot v1, v2, and v3.
+The x-axis lists the chatbot versions, while the y-axis displays the CUQ scores from 0 to 100.
+NurtureBot v1 has a score of 85.4, v2 scored 87.2, and v3 achieved the highest score of 91.3.
+The graph shows the lowest, median, and highest scores for each version, with a steady increase in the lowest scores from v1 (43.8) to v3 (75), while the median and highest scores remain consistently high, demonstrating improvements in chatbot usability over time.
+Figure 12.
+Comparison of CUQ questions of three Versions of NurtureBot.
+This figure displays the differences between NurtureBot v1, v2, and v3 across various chatbot usability dimensions.
+The left column contains positive feedback questions, while the right column contains negative feedback question, showing clear improvements across versions, particularly in how engaging and easy to use the chatbot became over time.
+[…]
+A comparison chart showing the usability scores of NurtureBot v1, v2, and v3.
+The figure contains two sets of bar graphs for each version: one set for positive statements (left) and another for negative statements (right).
+NurtureBot v1 (blue) shows more neutral scores, while v2 (yellow) and v3 (pink) show improvements in positive responses and a reduction in negative responses.
+This comparison demonstrates progressive enhancement in user satisfaction across all chatbot versions.
+Furthermore, in comparison with related literature at the time of writing this paper, to the best of our knowledge, NurtureBot v3 had achieved the highest reported CUQ score of 91.3 (See Table 3).
+While early rule-based systems such as MYA using FlowXo (Larbi et al., 2022), and ChatPal using Rasa (Boyd et al., 2022) reported lower scores when compared to more advanced NurtureBot versions using GPT-4 and DIAS using GPT-4 Turbo (Fersch et al., 2022), which achieve significantly higher scores.
+Notably, despite all NurtureBot versions using the same base technology, the iterative development of a human-centred interaction layer led to significant improvements in usability, demonstrating that thoughtful design can enhance user experience even without changes to the underlying technology.
+Table 3.
+In comparision with CUQ scores reported in literature NurtureBot v3 achieves the highest CUQ score of 91.3, followed by NurtureBot v2 and v1, demonstrating progressive improvements in usability, highlighting the effectiveness of iterative human-centred design in enhancing chatbot performance.
+Chatbot Name Base Technology Domain Participants CUQ Score
+MYA(Larbi et al., 2022) FlowXo (Rule-based) Physical Activity 30 57.4
+ChatPal(Boyd et al., 2022) Rasa Mental Health 10 65
+Chat Ella(Zhang and Song, 2024) GPT-2 Chronic Diseases 64 68.31
+GPT-Simulated Patient(Holderried et al., 2024) GPT-3.5 Medical Education 28 77
+DIAS(Fersch et al., 2022) GPT-4 Turbo Educational Services 103 80
+NurtureBot v1 GPT-4 Parental Wellbeing 32 85.4
+NurtureBot v2 GPT-4 Parental Wellbeing 32 87.2
+NurtureBot v3 GPT-4 Parental Wellbeing 64 91.3
+Based on our findings for co-designing user-LLM interactions, in Parts I-IV, we derive implications for design of AI-driven parental support systems presented in Table 4.
+Table 4. Implications for Design of AI-driven Parental Support Systems
+Implication Description
+Good Data Sources for AI A key foundation for successful AI is access to reliable and well-structured data sources.
+NurtureBot’s effectiveness depended on the quality and relevance of the information provided.
+Future iterations must use up-to-date, verified sources, including health organisation-endorsed health guidelines and localised resources, to enhance trust and provide contextually relevant information.
+User Understanding, Control, and Outcome Improvement Effective human-centred AI balances understanding, controlling, and improving interactions.
+Users should have transparency in their interactions with the AI and be able to guide the interaction through dynamic in-situ controls.
+Enabling users to refine the agent’s responses enhances the quality of interactions and makes the AI more responsive to user needs.
+Personalisation, Contextualisation, and Localisation Users expect personalised interactions, asking for the chatbot to remember personal details such as their names, their child’s name, where they live, and their previous conversations.
+These aspects help the chatbot tailor responses to user’s specific needs.
+Future designs should consider secure methods for personalisation without compromising privacy.
+Community and Peer Support Users desire more community-oriented support for parental wellbeing.
+This reflects the need for AI systems to connect users with localised communities, parenting groups, or forums.
+AI could act as a bridge between users and real-world community resources, providing emotional and practical support.
+Multi-modal Interactions Users expressed a need for multi-modal interactions such as video, voice, and image-based communication alongside text.
+This allows flexibility in how users interact with the chatbot, broadening accessibility for users with different preferences, literacy levels, or disabilities.
+Emotional Design Emotional connection emerged as a key element of success for NurtureBot.
+Many participants referred to the assistant as a ”friend,” indicating the importance of empathy and emotional support.
+However, it’s crucial to ensure users understand the AI’s role as a tool and avoid fostering over-reliance.
+Co-Design and Iterative Testing Co-design played a critical role in aligning NurtureBot’s features with user needs.
+Iterative testing allowed us to refine the system based on feedback.
+This iterative process of testing, redesigning, and retesting ensured that the final product provided the best possible user experience.
+Figure 13. Comparison of prompt architectures across NurtureBot versions, highlighting the co-designed interaction layer introduced in v2 and refined in v3.
+[…]
+xxx
+7. Discussion
+7.1. Co-Design as a Catalyst for LLM-Based Agent Development
+Our study demonstrates the potential of co-design in shaping LLM-based agents, particularly in sensitive domains such as parental wellbeing.
+Our research aligns with broader HCI research, which shows that empowering users to co-design results in more effective and engaging systems for designing both parental wellbeing (Prabhakar et al., 2017; Benjamin et al., 2024) and broader AI (Hossain and Ahmed, 2021; Delgado et al., 2023; Viswanathan et al., 2022b).
+On integrating the priorities, lived experiences, and user-generated dialogues directly into the system’s prompt architecture, by designing the interaction layer, we improved the chatbot’s ability to respond empathetically and personally and redefined the benchmarks for usability and user satisfaction in AI-driven systems.
+This approach exemplifies a human-centred AI method, prioritising collaboration and inclusivity in the design process.
+Our co-design method integrated few-shot dialogue examples generated by participants role-playing as the chatbot directly into the prompt architecture.
+Unlike traditional rule-based chatbots, which often rely on predefined, inflexible conversational paths, the co-design process capitalised on the generative capabilities of LLMs, leveraging user needs to iteratively refine chatbot interactions, and building on emerging user-centred design methods (Benjamin et al., 2024; Shi et al., 2024; Hwang et al., 2023).
+This extends existing methods that are typically used for improving LLMs, e.g., fine-tuning, transfer learning, and reinforcement learning from human feedback (RLHF).
+The ARC method was instrumental in the co-design process, allowing us to iteratively refine the interaction layer based on user feedback.
+However, the application of such participatory methods comes with its own limitations.
+Designing user-aligned systems using ARC requires longitudinal studies and sustained engagement from participants who are deeply invested in the design process.
+While our implementation successfully captured user needs and incorporated them into the system, scaling such methods may require alternative or complementary design research techniques.
+7.2. The Interaction Layer Prompt Architecture
+Our findings demonstrate that incorporating a user-co-designed interaction layer, employing the tripartite human-centred AI principles of understanding, control, and improving interactions, significantly enhanced the usability and user experience of subsequent versions of NurtureBot.
+This work contributes to the growing body of research on leveraging human-centred computing for AI alignment (Gilbert, 2024).
+The interaction layer prompt architecture (illustrated in Figure 13), extends techniques like Chain-of-Thought prompting (Wei et al., 2022), ReAct (Yao et al., 2022), and Deliberative Alignment (Guan et al., 2024) to integrates user feedback and employs state transitions to dynamically align system behaviour with human-centred design principles in sensitive contexts.
+The Understand, Control, and Improve framework maps closely to Nielsen’s usability heuristics (Nielsen, 1994), particularly visibility of system status, user control and freedom, and recognition over recall.
+These principles address the unique challenges of conversational interfaces, which, unlike modern mobile applications with expansive graphical interfaces, often leave users uncertain about system capabilities.
+Chatbots can force users into frustrating trial-and-error processes, undermining user trust and engagement.
+Through our framework, we introduced explicit mechanisms to inform users about the chatbot’s capabilities, guide interactions across states, and empower users to direct conversations toward meaningful outcomes.
+This approach not only improves usability but also instills confidence in navigating and leveraging the system effectively.
+Looking forward, we envision a future where users can seamlessly engage in Understanding, Controlling, and Improving AI systems without the need for rigid frameworks.
+Transparent and flexible interaction layers hold the potential to enable intuitive, user-aligned engagements that dynamically adapt to diverse needs, ultimately supporting bi-directional human-AI alignment (Shen et al., 2024).
+While this work represents an initial step, we anticipate that advancements in participatory methods and interaction design will further refine and expand the capabilities of generative AI systems, fostering truly human-centred AI experiences.
+7.3. User Anthropomorphism and Ethical Boundaries
+A key insight from our study is that despite our attempts to discourage anthropomorphising the chatbot, participants continued to do so, and even explicitly requested anthropomorphising features through multiple stages of co-design.
+Participants often described the chatbot as a “friend” or a “companion,” suggesting a deep emotional connection fostered by its empathetic tone and personalised responses.
+While this can enhance user engagement and perceived support, it raises ethical concerns about over-reliance on non-human systems in emotionally charged contexts (Weizenbaum, 1966).
+These findings reiterate the need for careful balancing between making interactions more empathetic, while maintaining a clear distinction between chatbots and real human relationships.
+It will be highly tempting for designers focused on user satisfaction to exploit the ease with which LLM-powered chatbots can mimic human-like empathy.
+However, the potential for this to lead to parasocial human-AI relationships (Maeda and Quan-Haase, 2024) poses huge risks, from the illusion of reciprocity to entrenching social isolation.
+Future research should explore how to increase healthy forms of engagement without compromising emotional safety.
+Novel human-centred AI approaches may be required to counter the ever-present human desire to attribute intelligence and emotion to machines; these could include approaches that encourage contestation (Ploug and Holm, 2020), centre human choice rather than AI decision-making (Miller, 2023b), or simply provide more and better training to help users avoid anthropomorphisation.
+8. Limitations and Future Directions
+Our study offers valuable insights into the co-design of LLM-based parental wellbeing assistants.
+However, certain constraints warrant consideration, both as reflections on our current findings and as signposts for future inquiry.
+Study Design and Population
+Our five-minute daily interaction window sought to ease participant burden, yet this brief engagement may have curtailed the depth and complexity of user responses.
+Although the data gathered offer a foundational understanding of initial interaction patterns, extended studies would allow richer insights into how user experiences and relationships with the chatbot evolve over time.
+We chose to focus on parents of children aged five or six, rather than those navigating the acute perinatal period.
+While this approach reduced ethical risks–minimising emotional distress and time burdens for parents with newborns–it may have limited the immediacy of the experiences examined.
+Simulating early stages of parenthood cannot fully capture the depth of challenges faced during pregnancy or postpartum.
+Future work should explore working with parents currently in these early stages of parenthood, supported by robust ethical frameworks, clinical trials, and collaborations with public health organisations, ensuring that vulnerable populations are engaged responsibly and with adequate safeguards in place.
+Risks of Misinformation and Bias
+As with any LLM-based tool, the potential for misinformation or ‘hallucinations’ remains a pressing concern.
+Although we endeavoured to ensure evidence-based, trustworthy responses, it is not possible to eliminate inaccuracy from LLMs.
+Integrating domain expert oversight and real-time content verification will be critical to maintaining credibility, especially when the use case could easily spill over into healthcare.
+Our intention for NurtureBot was to promote parental wellbeing rather than to dispense medical advice; using it for the latter could be highly risky given the poor performance of current LLMs on medical advice tasks (Agarwal et al., 2024).
+Saftey
+Safety is a critical concern for a wellbeing agent like NurtureBot.
+In real-world scenarios, edge cases may emerge where users, either deliberately or unintentionally, prompt the chatbot to generate harmful or inappropriate content.
+Such cases could lead to emotional distress or misinformation that exacerbates a user’s situation.
+Ensuring that the chatbot does not generate content that could cause harm—whether directly or indirectly—requires robust guardrails, such as toxic content filters, conversational safeguards, and monitoring mechanisms.
+Generalisability and Context Sensitivity
+Our findings are anchored in a specific demographic and cultural setting from the United Kingdom.
+Parenting experiences vary greatly across sociocultural, economic, and personal contexts, and future research should consider more diverse populations.
+Tailoring LLM-based supports to different communities and circumstances will enhance relevance and ensure broader applicability.
+9. Conclusion
+The significance of comprehensive perinatal care cannot be overstated, as it plays a pivotal role in mitigating risks associated with perinatal mind-body duality issues, fostering a positive environment that promotes healthy development of the parent(s), the child, and the long-term stability of the family unit.
+We explored iterative co-design of AI-powered parental wellbeing support with NurtureBot, engaging users through ARC sessions, where they were not only able to prioritise key issues but also contribute to the redesign of user-LLM interactions, for improved user comprehension, control, and task outcomes.
+The resulting interaction layer significantly improved the user experience across versions, suggesting that AI-driven systems can better align with real-world needs by empowering users to shape their design.
+By prioritising the human experience, AI tools can not only enhance wellbeing but also foster stronger, more resilient communities and create scalable, long-lasting solutions to meet diverse real-world needs.
+Acknowledgements.
+We express our sincere gratitude to all the parents who participated in our study, with special thanks to the 32 parents who proactively contributed to the co-design of NurtureBot.
+Your insights and dedication were invaluable in shaping this project.
+We also extend our heartfelt thanks to the children’s organisation, Committee For Children, for generously sharing their wealth of knowledge and resources from the Nurture project.
+This research is supported by UKRI Future Leaders Fellowship (MR/T041897/1).
+APPENDIX
+Appendix A Qualitative Results from Part I NurtureBot v1 Testing and Analysis
+A detailed account of the seven key problems identified through thematic analysis of participant feedback from the initial trials of NurtureBot v1 in Section 3.3, serving as a foundation for subsequent design iterations.
 Problem 1: Unable to Chat: Participants expressed frustration with NurtureBot’s tendency to cut off their conversation.
@@
 ”Ok, but it felt like the long responses with lots of links were a bit overwhelming and lacking in connection.”
-4.3.1. Quantitative results
-Our subjective questions regarding NurtureBot v1’s task satisfaction received an average score of 4.29 out of 5, while interactive questions averaged 4.36 for understandability, 3.97 for controllability, and 3.88 for improvability.
-The standard questionnaire analysis resulted in a mean CUQ score of 85.4 out of 100, which, when compared to the CUQ benchmark of 68 (Holmes et al., 2019) in a one-sample t-test, was extremely statistically significant (t(29) = 7.11, p ¡ 0.0001, mean difference = 17.40, 95% CI [12.40, 22.40], SE = 2.45).
-These scores and their implications are discussed further in Section 8.3, alongside comparisons with NurtureBot v2 and v3.
-After identifying seven key problems with parent-NurtureBot v1 interactions in Part II, our goal in Part III was to co-design by tasking users to prioritise problems and imagine ideal solutions.
-5. Part III: ARC Co-Design Study
-To begin addressing RQ2 – “What key user needs must a parental wellbeing assistant address?
-How do parents describe their ideal vision of how the assistant should function, particularly regarding their understanding of its workings, control over it, and opportunities to improve the outcomes?,”– the third phase of the study involved Asynchronous Remote Communities (ARC) co-design sessions (MacLeod et al., 2016), designed to directly engage participants in the iterative design process.
-These sessions were facilitated using the Miro platform (Miro, 2011), allowing the same 32 participants who tested NurtureBot v1 (see Subsection 4.1) to return to collaboratively co-design solutions asynchronously, thus accommodating different schedules, total anonymity, and fostering broad participation.
-The ARC sessions involve two core activities over three sessions (1 week), prioritising, and addressing key issue(s) with NurtureBot v1.
-The average participation was 29.33 out of 32 (91.63%) participants per day.
-To help familiarise participants with using Miro, we designed a training activity and instructional tips for using Miro on mobile devices, as well as providing direct messaging support via Prolific.
-5.1. ARC Activity One: Diagnosing and Prioritising Pain Points
-The goal of the first activity was to establish which of the expressed difficulties were the biggest priority to address.
-Using the seven key problems from NurtureBot v1’s technology trial (see Section 4.3), we asked participants to prioritise these issues.
-By disclosing these key issues, the goal was to ensure that all participants could see the broader scope of the negative feedback gathered from others within their group.
-5.1.1. ARC Activity One: Method
-We crafted a ”problem statement” for each issue, synthesising the findings without directly replicating any participant’s verbatim feedback (listed as problem statements in Figure 4).
-This method not only helped verify our interpretation and synthesis of the feedback but also prevented potential bias, ensuring participants did not favour their own comments when voting.
-To provide sufficient scaffolding for reflection, we organised the activity as follows:
-Reviewing Identified Problems: Participants were guided through a review of the seven major problems.
-As they reviewed these issues, they were asked to select one of three stickers—“This applies to me,” “This may apply to me,” or “This does not apply to me”—and place it on each of the seven problems.
-This step was important for setting a common understanding of the challenges identified from the group.
-Ranking and Prioritisation: After reviewing the problems, as the next step, participants were asked to rank them according to their personal experiences, using stickers (gold, silver, and bronze icons provided on the Miro board) (See Figure 4).
-This step was designed to capture participant perspectives on the perceived severity of each problem, which would inform the prioritisation of design efforts in subsequent activities.
-Figure 4.
-Results from ARC Activity One voting by 30 participants, where key problems in interacting with NurtureBot were ranked using an Olympic points system.
-Complete Miro board including the other set-up phases participants went through before voting can be found in the supplementary materials.
-[…]
-xxx
-5.1.2. ARC Activity One: Results
-Based on participants responses, the following options were voted as the top three problems by participants:
-  • Problem 1 - I wanted to have a chat but NurtureBot kept cutting me off by redirecting me or giving me paragraphs of information. emerged as the top challenge with NurtureBot.
-  • Problem 3 - The suggested resources were not localised to my country or area was the second priority.
-  • Problem 4 - I want NurtureBot to remember things! E.g., who I am, my baby’s name and our previous conversation and context was ranked third.
-We next took the top-ranked problem (Problem 1) to begin considering how we might enhance “successful chat problem” in Activity 2.
-5.2. ARC Activity Two: Co-Designing Solutions
-The second activity focused on solution generation for problem 1. Participants were asked to step into the role of NurtureBot: “Imagine You Are NurtureBot and Fill in the Blanks.”
-5.2.1. ARC Activity Two: Method
-The activity was structured as follows:
-Scenario-Based Problem Solving: As the first step, participants were presented with three specific scenarios corresponding to the three main features of NurtureBot (i) interacting with NurtureBot for the first time and wanting to have an empathetic chat, (ii) seeking wellbeing exercises to manage work stress, and (iii) seeking information to soothe their baby; that exemplified trying to have a successful chat aiming to use all three features of NurtureBot.
-For each scenario, participants reviewed an example response and were asked to consider how NurtureBot could help users to better understand its offerings, control the conversation, and improve its outcomes.
-Collaborative Dialogue Writing: Next, participants were invited to take on the role of NurtureBot and rewrite it’s dialogue for each scenario.
-We offered cues that encouraged participants to work on ”Understanding, Controlling, and Improving” NurtureBot, similar to the idea of a cybernetic loop (Patten and Odum, 1981) (see Section 2.2).
-This part of the activity encouraged participants to think critically about how the chatbot could better meet their needs in specific contexts, and which would inform NurtureBot v2 (See Figure LABEL:UCI).
-Iterative Feedback and Voting: As the last step, after proposing solutions, participants reviewed and built upon each other’s suggestions.
-We invited participants to vote on the best ideas (See heart stickers in Figure LABEL:UCI), ensuring that the most popular and potentially effective solutions were prioritised for implementation.
-5.2.2. ARC Activity Two: Results
-ARC Activity Two generated a rich dataset of 189 participant-generated dialogues.
-Our thematic analysis of these dialogues (as in Section 4.2) generated seven potential features to address user needs.
-We summarise the seven features below and in Table 1, map examples from participant-generated dialogues onto each feature.
-These features were distilled into design specifications that directly informing the next iteration.
-Table 1. Mapping Parent-Written Dialogues from ARC to NurtureBot’s Features
-Dialogues written by parents Mapped Features
-“Hi welcome to NurtureBot your online support and friendly advice guru for all your parenting needs!
-Let’s get started with what’s important and that’s how you are?
-What would you like support with today?
-:)” Feature 1.
-Use of Metaphors for Understanding
-Feature 2. Beginning with Empathy
-“Remember, if you find that this exercise isn’t quite what you need, feel free to ask for an alternative.
-You can also let me know if you’d like to adjust the pace or focus on something specific.
-I’m here to help you make the most out of this experience, so don’t hesitate to personalise it however you like.” Feature 2.
-Beginning with Empathy
-Feature 5. Personalising and Deepening Sessions
-Feature 6. Offering Clarification and Reassurance
-Feature 7. Summarising Exercises and Offering Follow-Up Options
-“Was this support helpful?
-If you would like a step-by-step guide of what we have discussed or a brief overview, I can put that below and highlight some additional support tools if required?
-Your happiness is important!
-I would love to support you further.” Feature 4.
-Providing Step-by-Step Guidance
-Feature 6. Offering Clarification and Reassurance
-Feature 7. Summarising and Offering Follow-Up Options
-”Don’t worry, you are not alone!
-It is very common to face this problem as a new parent.
-Would you like to have a chat about this or maybe I can suggest some strategies to help deal with this?” Feature 2.
-Beginning with Empathy
-Feature 3. Introducing Features Clearly
-Feature 6. Offering Clarification and Reassurance
-“I am here to help you through your journey as a new parent.
-It can often be an overwhelming time so if you have anything specific I can help with or you just need to use me as a sound-board—I am here for you, day or night.” Feature 1.
-Use of Metaphors for Understanding
-Feature 2. Beginning with Empathy
-Feature 6. Offering Clarification and Reassurance
-“Hello!
-I’m here to support you on your parenting journey like a trusted friend!
-To get started you can just describe what you need in your own words.
-For example, you can say ‘I need to talk,’ ‘Help me relax,’ or ‘I have a question about my child’s development.’ I’m here to help in whatever way works best for you!” Feature 1.
-Use of Metaphors for Understanding
-Feature 2. Beginning with Empathy
-Feature 3. Introducing Features Clearly
-Feature 7. Summarising and Offering Follow-Up Options
-“I’m here to help answer or guide any questions you have about parenting, or how to manage the stress of it all.
-First of all, I’d like to know a bit more about you?
-How are you today?” Feature 2.
-Beginning with Empathy
-Feature 3. Introducing Features Clearly
-Feature 5. Personalising and Deepening Sessions
-“You’re in control and I will follow your lead so please let me know how you would like this conversation to go.” (Would be nice to have something eg a button that when tapped/hovered over suggests ways you can influence responses.) Feature 5.
-Personalising and Deepening Sessions
-Feature 7. Summarising and Offering Follow-Up Options
-“Parenting can be tough on the mind and body so let’s explore ways as what gentle exercises can support us with. Would this be helpful for you?” Feature 2. Beginning with Empathy
-Feature 3. Introducing Features Clearly
-Feature 6. Offering Clarification and Reassurance
-“If you need more clarifications on what I can do, please ask I can present information in other formats such as tables and bullet points.” Feature 4. Providing Step-by-Step Guidance
-Feature 6. Offering Clarification and Reassurance
-“There are a number of different tools that would help you find the right balance when facing these challenges.
-Would you like me to list all the different tools, or would you like to provide some more details on what specific kind of tools you are looking for?” Feature 3.
-Introducing Features Clearly
-Feature 4. Providing Step-by-Step Guidance
-Feature 5. Personalising and Deepening Sessions
-Feature 1.
+Appendix B Qualitative Results from Part II ARC Activity Two
+A comprehensive mapping of seven user needs, derived from thematic analysis of dialogues created during participant role-play as NurtureBot in ARC sessions (see Section 4), highlighting the co-designed inputs that shaped the interaction layer.
+User Need 1.
 Use of Metaphors for Understanding: Parents suggested a range of metaphors to help them better understand and form a meaningful connection with NurtureBot.
@@
 This blending of technology with a human-like empathy was key to fostering deeper, more effective engagement.
-Feature 2. Beginning with Empathy:
+User Need 2. Beginning with Empathy:
 Parents highlighted the importance of empathetic conversation starters, expressing that this initial approach sets the tone for the entire interaction.
@@
 Whether engaging in mindful exercises or offering parenting information, the empathetic tone had to persist throughout, without pretending to be human or to understand human pain, thereby reinforcing NurtureBot’s supportive role.
-Feature 3. Introducing Features Clearly:
+User Need 3. Introducing Features Clearly:
 A recurrent goal of parents was to highlight the importance of clearly introducing NurtureBot’s core features early in the interaction.
@@
 By continuously reinforcing NurtureBot’s capabilities, within the context of the conversation, it became easier for parents to make informed choices during the interaction, thereby maximising the effectiveness of each conversation.
-Feature 4. Providing Step-by-Step Guidance:
+User Need 4. Providing Step-by-Step Guidance:
 Parents wanted NurtureBot to go beyond simply providing information.
@@
 Breaking down long paragraphs of information into bullet points and allowing them to control the flow of the conversation through a step-by-step guide, with parents ”choosing the next step” from a range of available options, empowered parents, making them feel in control of the interaction and able to absorb information at their own pace.
-Feature 5. Personalising and Deepening Sessions:
+User Need 5. Personalising and Deepening Sessions:
 Personalisation was crucial to the parents’ experience with NurtureBot.
@@
 This deep personalisation allowed for a more meaningful and tailored exchange, where NurtureBot could address the unique concerns of each parent based on the context of their lives.
-Feature 6. Offering Clarification and Reassurance:
+User Need 6. Offering Clarification and Reassurance:
 At the end of each interaction, parents wanted NurtureBot to provide them with an opportunity to “clarify” or “revisit” the topic at hand.
@@
 This reassurance, especially when discussing difficult topics, coupled with the option to fine-tune their experience, would take NurtureBot beyond offering static advice.
-Feature 7. Summarising and Offering Follow-Up Options:
+User Need 7. Summarising and Offering Follow-Up Options:
 Summarisation played a key role in allowing parents to reflect on their interactions and “takeaways” with NurtureBot.
@@
 Imagination of this interactive flow of suggested participants wanted to prevent decision fatigue with more seamless and intuitive interactions, reinforcing the effectiveness of NurtureBot as a comprehensive support tool even for new and non-technical users.
-Our findings from ARC in Part III, eliciting key user needs to have a successful chat with Nurture were then directly employed to implement NurtureBot v2 in Part IV.
-6. Part IV: The Interaction Layer: NurtureBot v2
-To continue addressing RQ3 - “How can we employ co-design methods to craft interactions around AI, enhancing usability and the overall user experience?.” The development of NurtureBot v2 focused on refining interactions to enhance its user experience.
-We created an interaction layer central to how the LLM-based chatbot engages with users (see prompt illustrated in Figure 7).
-The three aspects of NurtureBot v2’s interaction layer, namely the core interaction principle of NurtureBot, interaction levels within each feature (denoted by four big blue dots in Figure 7, and interaction elements further embedded within interaction levels (denoted by multiple big yellow dots in Figure 7) were directly shaped by the ARC co-design sessions and implemented using a structured prompt architecture that ensured consistency and depth in user interactions.
-Our design approach extends Beaudouin-Lafon’s theory of instrumental interaction (Beaudouin-Lafon, 2000), adapting it to LLM-based AI agents by creating three instruments within the interaction layer, invisible in the user interface, that reify user needs and are polymorphised, and reused to improve situational interaction across the prompt architecture.
-6.1. Core Interaction Principle
-NurtureBot v2’s development was guided by three foundational capabilities we wanted users to have in relation to the chabtot, and vice-versa: understanding, control, and improvement.
-These core interaction principles (denoted with a yellow rectangle in Figure 7) guided all interactions with users.
-This model, inspired by the concept of the ‘cybernetic loop’ (Patten and Odum, 1981), and supported by the human-centred AI research endeavours discussed in Section 2.2, was designed to ensure that every exchange was meaningful, user-centred, and adaptable to the individual needs of each parent.
-This principle provided a template for each type of exchange, guiding the chatbot on how to initiate conversations, present options, and offer follow-up support.
-6.2. Three Interaction Levels
-1.
-Understand Level: The chatbot’s first task was to understand and acknowledge the user’s situation while helping the user grasp how the chatbot could assist them.
-For instance, it would ask, ”How are you feeling today?” or ”I can help you with a list of options?” We employed the metaphors of ”friendly AI support” and ”AI assistant” to prevent users from anthropomorphising the chatbot, ensuring they maintained a clear understanding of its role as a helpful and empathetic tool rather than attributing human relationships to it.
-This phase was essential for setting a supportive tone and ensuring the chatbot could convey its capabilities and tailor its responses to meet the user’s needs effectively.
-2.
-Control Level: Once the user’s needs were understood, the chatbot entered the control level, where it provided options to guide the interaction along with its content.
-This level was designed to give users a sense of agency and control over the conversation.
-Users could choose to elaborate on their concerns, request a summary, or switch topics.
-Users would be often provided with ”hints” such as revise, restart, and elaborate; designed in ARC sessions, in accordance with the context of the chat to enable them to effectively lead NurtureBot towards desired outcome.
-3.
-Improve Level: After addressing the user’s needs, NurtureBot gave a summary and sought feedback to access user satisfactions and improve the interaction by offering additional resources or to revise the conversation (as parents preferred).
-This phase ensured that the conversation concluded on a constructive note, always asking for user feedback, with the user feeling supported and informed.
-Figure 7.
-This figure illustrates the prompt of NurtureBot v2, with interaction layer showcasing how the chatbot guides user engagement through the core principle of ”Understand, Control, and Improve,” with a core interaction principle, interaction levels and elements answering in-situ user needs identified in ARC.
-[…]
-A detailed diagram of NurtureBot v2’s prompt architecture hosting an interaction layer, divided into core principles: Understand, Control, and Improve.
-It highlights the chatbot’s empathetic chatting, wellbeing exercises, and parenting information, showing how the system tailors responses based on user preferences and inputs.
-Examples of user dialogues, actionable hints, and resources are shown alongside elements for adapting to contextual user needs, illustrating NurtureBot’s design for personalised support.
-6.3. Interaction Elements
-1.
-Undecided Exchanges When users were undecided, NurtureBot aimed to first understand their hesitation or indecision.
-With started elements such as, ”What’s on your mind today?” or ”How can I assist you?” to encourage the user to reflect on their needs.
-Metaphorical elements like ”assistant” were also embedded in the chatbot’s dialogue to offer a supportive role, and reassurance elements were designed to not overwhelm user with options, reflecting a tone of gentle guidance.
-For instance, ”I’m here to support you in whatever way works best for you” to help users feel comfortable in sharing their thoughts or challenges.
-2.
-Empathetic Chatting: The chatbot was programmed with elements to focus on active listening and providing emotional support.
-We used dialogue examples to create non-judgemental elements and avoiding any advice.
-The chatbot’s responses were crafted to mirror those of a trusted confidant, providing reassurance and understanding.
-For example, these elements instructed NurtureBot to initially wait: “Tell me more about what’s been on your mind today.” Based on the ARC study insights, the chatbot would then offer subtle prompts to encourage to-and-fro discussion, e.g., “That sounds really challenging.
-How has that been affecting you?” This approach was designed to build trust and ensure that the user felt heard and understood.
-3.
-Wellbeing Exercises: Guiding elements for wellbeing exercises included mindfulness practices and relaxation techniques.
-The interaction was structured with personalised elements, offering tailored exercises based on user concerns and alternative options.
-The chatbot also ensured that exercises were broken down into manageable steps, with the user controlling the pace of the session.
-To embed collaborative elements, the prompts also ensured that NurtureBot could guide users through activities step by step, e.g., ”Would you like to try a simple breathing exercise together?
-Let me know when you’re ready to begin.” As such, the user could control and personalise the pace and depth of the exercise.
-4.
-Parenting Information: NurtureBot v2 provided a structured and digestible format for parenting content.
-Instead of overwhelming users with lengthy text, interactive elements offered summaries and gave users the option to dive deeper into specific topics.
-The chatbot also cross-referenced reliable sources to ensure that the information was accurate and relevant.
-The offered information was also adapted, based on the specific situation, preference, and needs.
-For instance, if a user asked about managing a fussy baby, the chatbot might respond, ”I have a few techniques that could help.
-Would you like a quick overview, or shall we talk about your baby and things you have tried already?” This ensured that users received information that was most useful to them.
-NurtureBot v2, featuring an interaction layer created in Part IV with parent-designed dialogues, was tested in Part V to ensure that the issues identified in the previous version were effectively addressed.
-7. Part V: Technology trial of NurtureBot v2
-7.1. Participants, Method, and Data
-The initial set of participants (See Section 4.1) trialled NurtureBot v2 over five days, generating a total of 104 conversations and their corresponding survey feedback (See Figure 2).
-We analysed the survey data in accordance with our thematic approach (See Section 4.2).
-Participation averaged 69.3% (20.8 out of 30), a decline from 90% in Part II (v1 trial) and 91.6% in Part III (ARC session).
-Eight participants contacted via Prolific cited reasons such as, ”I’ve enjoyed your study, but NurtureBot has improved so much I don’t know what more I can give feedback on.”
-7.2. Findings of second technology trial - NurtureBot v2
-While our data revealed highly positive feedback for NurtureBot v2, with participants praising the chatbot for being “more chatty” and “engaging” than its predecessor, as with NurtureBot v1, we only focused the key interaction problems for our thematic analysis, focusing on 78 instances of negative feedback and/or suggestions for improvement (of 208 instances of feedback collected (37.5%)), coding 12 unique pain points, which we then grouped into three themes briefly summarised below.
-We further discuss a comparative analysis of our findings in the Results section (see Section 8.3).
-Problem 1 Some participants pointed out that NurtureBot still occasionally responded with long paragraphs, making the information feel overwhelming. One participant noted,
-“It was good, but very long answers though and lots of information to take in.”
-Problem 2 Most participants noted a bug, where it repeatedly used the same prompts—“elaborate, revise, and restart”—which reduced the creativity and variety in responses. A participant mentioned,
-“The repetition of the last sentence ’if you need me to elaborate… just let me know’ was quite annoying.”
-Repeating Themes As with v1, participants made suggestions for incorporating buttons for conversational hints, multi-modal content (images and videos), and memory capabilities for further personalisation, and connection to resources and communities in their local area.
-7.2.1. Quantitative results.
-Our subjective questions regarding NurtureBot v2’s task satisfaction received an average score of 4.41 out of 5, while interactive questions averaged 4.50 for understandability, 4.42 for controllability, and 4.44 for improvability.
-The questionnaire analysis resulted in a mean CUQ score of 87.2 out of 100, which, when compared to the CUQ benchmark of 68 (Holmes et al., 2019) in a one-sample t-test, was extremely statistically significantly (t(20) = 8.98, p ¡ 0.0001, mean difference = 19.20, 95% CI [14.74, 23.66], SE = 2.14).
-These scores and their implications are discussed further in Section 8.3, alongside comparisons with NurtureBot v1 and v3.
-The two minor interaction issues persisting in NurtureBot v2, identified in Part V, were targeted for refinement in NurtureBot v3 and tested with a new cohort of participants in Part VI to further validate our efforts.
-8. Part VI: Prototype Improvement: NurtureBot v3 and Validation
-We iterated NurtureBot v3 as briefly summarised here:
-  • First, we addressed the issue of lengthy responses by limiting the chatbot’s answers to 50 words, ensuring information was delivered in concise, digestible segments.
-  • Second, we updated the core interaction principle, enhancing NurtureBot’s ability to provide contextual actionable hints, rather than offering generic options (denoted as struck-out text inside the big yellow rectangle in Figure 7)
-8.1. Participants, Method, and Data
-To assess the effectiveness of these improvements without learning and co-design biases, NurtureBot v3 was tested with a new cohort of 46 parents, 29 females and 17 males with an average age of 38.07.
-11 participants (23.9%) used ChatGPT everyday.
-A total of 185 conversations (denoting an average participation of 80.4% (37 out of 46 parents across all days)) and corresponding feedback were collected over five days using the same data collection (See Figure 2) and analysis method (see Section 4.2) as outlined in the NurtureBot v1’s trial, allowing for direct comparisons across versions.
-8.2. Findings of third technology trial - Nurture Bot v3
-Next we present a summary of the third trial findings, focusing on identifying ongoing interaction problems, following the interaction layer redesign.
-While feedback was overwhelmingly positive, our analysis focused the interaction problems (from 46 out of 370 (12.4%) negative feedback.
-We identified 10 unique pain points, which we group into the three themes, summarised below.
-We also provide a comparative analysis of the three Nurturebot versions in Section 8.3.
-Problem 1 Some participants wanted NurtureBot to “intelligently” provide more detailed responses when necessary. For example,
-“I guess the responses could be a little longer and detailed at times but not by much.”
-Problem 2 Few participants found it challenging to follow step-by-step exercises via text. One participant remarked,
-“A lot of the time the advice was to close my eyes then do X but when it’s text based this is obviously hard to follow. Perhaps tell them to read the advice then close eyes.”
-Repeating Themes Participants continued to request features for: multi-modal interactions (video, speech, and images), improved memory function to allow NurtureBot to remember personal details and past conversations, local resources, and hints to continue conversation as buttons.
-Furthermore, one participant expressed interest in viewing the original sources referenced in NurtureBot’s responses, allowing them to verify or explore the information further.
-8.2.1. Quantitative results.
-Our subjective questions regarding NurtureBot v3’s task satisfaction received an average score of 4.49 out of 5, while interactive questions averaged 4.63 for understandability, 4.04 for controllability, and 4.21 for improvability.
-The questionnaire analysis resulted in a mean CUQ score of 91.3 out of 100, which, when compared to the CUQ benchmark of 68 (Holmes et al., 2019) in a one-sample t-test, was extremely statistically significant (t(35) = 18.39, p ¡ 0.0001, mean difference = 23.30, 95% CI [20.73, 25.87], SE = 1.27).
-These score and their implications are discussed further in Section 8.3, alongside comparisons with NurtureBot v1 and v2.
-8.3. Results: NurtureBot v1 vs. NurtureBot v2 vs. NurtureBot v3
-8.3.1. Qualitative Results
-The comparison of qualitative results from NurtureBot v1, v2, and v3 reveal a clear progression in user satisfaction across versions, supported by a reduction in negative feedback gathered from our analysis 45.48% in NurtureBot v1 to 12.4% in NurtureBot v3.
-While our participants praised the “natural flow” of conversations and the “supportive, empathetic tone” of our improved versions, repeating themes for features like multi-modal interactions and enhanced memory for personalisation and localisation persisted across versions.
-In this section we detail our qualitative findings from evaluations of NurtureBot’s evolutions.
-A Successful Chat In NurtureBot v1, participants found the chatbot useful but expressed frustration with its tendency to cut off conversations, leading to a lack of conversational flow.
-Several participants noted that the responses were often patronising or generic.
-Informed by user feedback and ARC co-design, NurtureBot v2 introduced an interaction layer centred on the principles of Understand, Control, and Improve, enabling the chatbot to better grasp user needs, offer greater conversational control, and dynamically adapt across all features, giving users more autonomy and flexibility.
-We highlight these adapted elements in bold.
-Participants expressed satisfaction with NurtureBot v2, noting improvements in how natural and intuitive the conversations felt.
-One participant shared, “Good, lots of ideas for activities for playtime for my baby which was good, and Nurture continued to ask if I wanted further information at each step.” Participants also liked the system’s flexibility, saying, “The interaction was easy again, it feels like Nurture can handle whatever you ask it, whether it’s just chatting or asking for specific information.” Participant feedback also emphasised how the interaction layer made the conversation flow seamlessly, with one participant noting, “Interaction was the best this week—the responses from Nurture seemed natural and clear.” The introduction of the interaction layer also improved the clarity and engagement of the system.
-One participant mentioned, “It was really good.
-I found the introduction really helpful, particularly the clarification of the three kinds of things that Nurture could do.” Another shared, “Great!
-It made such a difference having the prompt to tell Nurture what I wanted.” This interaction layer enabled participants to feel more in control and allowed the chatbot to adapt to user input, creating a personalised, engaging experience.
-Although NurtureBot v2 was tested by participants who had co-designed it, NurtureBot v3 was tested with a new cohort, which confirmed the success of the interaction layer.
-One participant remarked, “Good.
-Conversation flows well, and everything provided was useful,” while another stated, “So seamless yet educative.” Our findings suggest interaction layer to played a crucial role in ensuring that NurtureBot responded dynamically to user input, leading to an improved overall experience.
-The Time Effect A key observation from the five-day testing period was the gradual shift in participants’ attitudes toward NurtureBot.
-Initially, participants were hesitant and skeptical, viewing it as an unfamiliar tool and doubting its ability to provide meaningful support.
-However, as they continued using NurtureBot over consecutive days, their perceptions became more positive, with many describing it in personal terms—some even referring to it as a “friend always there” to support them.
-For instance, one participant testing NurtureBot v1 described their early experience as “detached and robot-like,” but by day five, they said, “After using Nurture quite a few times now, it just feels like I’m chatting with someone that feels familiar and I have already built up a bit of a relationship with.
-It doesn’t feel artificial.” The time effect led many participants to believe that NurtureBot was being updated continuously each day.
-One participant testing NurtureBot v3 commented, “Wow, it was so much better today.
-I wrote quite a lengthy update, and it acknowledged every part, even pinpointing what was bothering me.” Participants began using personal pronouns like ”they” to describe NurtureBot v1.
-On day four, one participant remarked, “I think it is getting there with becoming more personal, maybe ask what the user’s name is, so they could talk using your name,” reflecting an increased emotional connection with the chatbot.
-This progression illustrates a time effect, where repeated interactions led to increased comfort, attachment, and even anthropomorphisation of the technology.
-The more time participants spent engaging with NurtureBot, the more they saw it as a supportive companion rather than just a tool.
-Community and Peer Support Tying back to the origins of the Nurture project, which was a community-based peer-to-peer mentoring program (as described in Section 3), it is clear that the concept of community and peer support remains a significant need for participants using NurtureBot.
-Throughout the three versions of NurtureBot, participants frequently searched for local activities, community support centres, and peer forums.
-This theme of seeking peer support and localised resources also emerged as the second biggest problem identified in the ARC (see Section 4).
-Additionally, several participants expressed a desire for more localised community resources, indicating the importance of having access to peer support networks within their geographical area.
-As one participant trying NurtureBot v2 remarked, “When I asked for classes in my local area it could not help, that would be a good feature.” and another trying NurtureBot v3 said “I realised when asking a question about services in my local area that it does not know my local area.” Feedback from our study suggests that there is an opportunity to evolve NurtureBot from a solely AI-driven one-on-one platform to a more comprehensive community-based support system.
-Personalisation and Contextualisation The third biggest challenge prioritised in our ARC session 5.1, the lack of personalisation and contextualisation in the chatbot’s responses, also emerged throughout the testing of all three versions.
-Participants repeatedly expressed the desire for NurtureBot to remember personal details such as their name, their child’s name, their unique circumstances, and previous conversations.
-Participant feedback such as, “Remembering my problems would still be nice,” and “The bot could ask our names and baby’s details to help make it more personal.,” in NurtureBot v2 and NurtureBot v3 echoed this need, emphasising how critical personalisation was to creating a more meaningful user experience.
-Participants consistently wanted the chatbot to feel more tailored to their individual experiences.
-While our experiment was not designed to collect or store personal data, and personalisation was not a focus in any version of NurtureBot, going forward, addressing personalisation and contextualisation will be essential to improving NurtureBot’s effectiveness.
-While safeguarding privacy and data security remains paramount, finding a way to create more contextually aware and personalised interactions will be crucial in meeting user expectations.
-8.3.2. Quantitative results
-To analyse our subjective questionnaire with four questions we created to measure task satisfaction and human-AI interaction (see questions 3-6 in Figure 2).
-To the data we collected for these questions, across the three technological trials (in Sections 4, 7, and 8), we employed both parametric Generalised Estimated Equations (GEE) and non-parametric Kruskal-Wallis and Dunn’s test approaches (see Table 2 and Figures 8, 9, and 10).
-Our results suggest:
-  1. (1)
-    Meeting User Expectations: GEE showed no significant differences between versions (p ¿ 0.05).
-Kruskal-Wallis yielded marginally significant results (H = 5.8539, p = 0.0536), suggesting potential subtle differences not captured by the parametric model.
-However, the changes between v1, v2, and v3 are not robust enough to be considered significant.
-  2. (2)
-    Understanding Functionality: Both parametric and non-parametric tests indicated significant improvements.
-GEE showed Version 3 to significantly outperformed Version 1 ($\beta$ = 0.2721, p = 0.001).
-Kruskal-Wallis and Dunn’s test confirmed this finding, revealing a significant difference between Version 1 and Version 3 (p = 0.003).
-  3. (3)
-    Control Over Actions: GEE indicated that Version 2 significantly improved user control compared to Version 1 ($\beta$ = 0.4521, p ¡ 0.001), which was evident considering v2 was designed by v1 participants.
-Interestingly, Kruskal-Wallis and Dunn’s test not only confirmed this (p ¡ 0.001) but also revealed a significant difference between Version 2 and Version 3 (p = 0.017) that wasn’t apparent with GEE.
-  4. (4)
-    Improving Performance: Both analytical approaches showed significant improvements in later versions.
-GEE model indicated that both Version 2 ($\beta$ = 0.5582, p ¡ 0.001) and Version 3 ($\beta$ = 0.3256, p = 0.005) significantly outperformed Version 1.
-Kruskal-Wallis and Dunn’s test corroborated these findings, showing significant differences between Version 1 and both Version 2 (p ¡ 0.001) and Version 3 (p = 0.007).
-Table 2. Comparison of User Experience with Subjective Questionnaire across NurtureBot’s Versions
-Question GEE Model Kruskal-Wallis Dunn’s Test (p-values)
-1. Meeting hopes and goals V2: $\beta$=0.1257, p=0.284 H=5.8539 Not performed
-V3: $\beta$=0.2015, p=0.061 p=0.0536
-2. Understanding functionality V2: $\beta$=0.1377, p=0.146 H=11.1215 V1-V2: p=0.727
-V3: $\beta$=0.2721, p=0.001** p=0.0038** V1-V3: p=0.003**
-V2-V3: p=0.225
-3. Control over actions V2: $\beta$=0.4521, p¡0.001** H=13.5127 V1-V2: p¡0.001**
-V3: $\beta$=0.0666, p=0.573 p=0.0012** V1-V3: p=0.757
-V2-V3: p=0.017*
-4. Improving performance V2: $\beta$=0.5582, p¡0.001** H=19.9734 V1-V2: p¡0.001**
-V3: $\beta$=0.3256, p=0.005** p¡0.0001** V1-V3: p=0.007**
-V2-V3: p=0.201
-Note: * p ¡ 0.05, ** p ¡ 0.01. V1, V2, and V3 represent Version 1, Version 2, and Version 3 of NurtureBot, respectively. GEE coefficients ($\beta$) represent the change relative to Version 1.
-While the parametric test provided insights into overall trends, the non-parametric tests revealed additional pairwise differences, particularly between Version 2 and Version 3 in terms of user control, where we had changed the bug of repeating conversational hints.
-This comprehensive analysis suggests that the iterative development of NurtureBot has led to significant improvements in user understanding and ability to improve performance, with some complex dynamics in user control across versions.
-Figure 8.
-Generalised Estimating Equation (GEE) coefficients comparing chatbot versions v2 and v3 to v1 of NurtureBot across four evaluation dimensions: (1) whether the task met user expectations, and interactive dimensions of: user- (2) understanding, (3) control, and (4) ability to improve performance and outcomes.
-Statistically significant improvements (p ¡ 0.05) are noted in dimensions (2), (3), and (4) in NurtureBot v2 and v3.
-[…]
-Figure 9.
-Kruskal-Wallis test results comparing the NurtureBot’s versions across four dimensions of subjective user experience.
-The test identifies statistically significant differences (p ¡ 0.05), with the strongest significance observed in the ”ability to improve performance and outcomes” dimension.
-[…]
-Figure 10.
-Heatmaps illustrating the pairwise comparison of three versions of the chatbot across three evaluation dimensions: (user- (2) understanding, (3) control and , and (4) ability to improve performance and outcomes.
-Statistically significant differences (p ¡ 0.05) are observed primarily between version 1 and version 3 across the evaluated dimensions.
-A reduced variance as NurtureBot is iterative developed, also indicates chatbot’s reliable performance over time.
-[…]
-Figure 11.
-Comparison of CUQ Scores Across NurtureBot Versions.
-The figure demonstrates the improvement in Chatbot Usability Questionnaire (CUQ) scores from NurtureBot v1 with a CUQ score of 85.4, to NurtureBot v2 with a CUQ score of 87.2, and NurtureBot v3, which achieved a higher CUQ score of 91.3.
-The lowest scores also improved significantly across versions, indicating enhanced chatbot usability with each iteration.
-[…]
-A graph comparing the CUQ scores of NurtureBot v1, v2, and v3.
-The x-axis lists the chatbot versions, while the y-axis displays the CUQ scores from 0 to 100.
-NurtureBot v1 has a score of 85.4, v2 scored 87.2, and v3 achieved the highest score of 91.3.
-The graph shows the lowest, median, and highest scores for each version, with a steady increase in the lowest scores from v1 (43.8) to v3 (75), while the median and highest scores remain consistently high, demonstrating improvements in chatbot usability over time.
-Further, quantitative results from the structured CUQ (Holmes et al., 2019) indicate a steady improvement in usability across all three versions of NurtureBot.
-As seen in the CUQ score comparison in Figure 11, the average CUQ score increased from 85.4 in NurtureBot v1 to 87.2 in v2, and further to 91.3 in v3.
-Notably, the lowest scores also saw significant improvements, rising from 43.8 in v1 to 67.2 in v2, and then to 75 in v3, reflecting enhanced user satisfaction over time.
-These results confirm that the iterative design process enhanced user experience, making the chatbot more intuitive, engaging, and easier to navigate.
-From NurtureBot v1 to v3, positive feedback improved significantly, with users rating the chatbot as increasingly engaging, easy to navigate, and better at understanding and supporting their needs (see Figure 12 Left).
-Negative feedback, such as the chatbot being perceived as ”robotic” or ”unfriendly,” decreased with each version, showing that iterative improvements effectively addressed these issues (see Figure 12 Right).
-By v3, users reported the most satisfying experience, with fewer errors and a more natural interaction flow, and a decreased overall variance (see Figure 11).
-Figure 12.
-Comparison of CUQ questions of three Versions of NurtureBot.
-This figure displays the differences between NurtureBot v1, v2, and v3 across various chatbot usability dimensions.
-The left column contains positive feedback questions, while the right column contains negative feedback question, showing clear improvements across versions, particularly in how engaging and easy to use the chatbot became over time.
-[…]
-A comparison chart showing the usability scores of NurtureBot v1, v2, and v3.
-The figure contains two sets of bar graphs for each version: one set for positive statements (left) and another for negative statements (right).
-NurtureBot v1 (blue) shows more neutral scores, while v2 (yellow) and v3 (pink) show improvements in positive responses and a reduction in negative responses.
-This comparison demonstrates progressive enhancement in user satisfaction across all chatbot versions.
-Furthermore, in comparison with literature at the time of writing this paper, to the best of our knowledge, NurtureBot v3 had achieved the highest reported CUQ score of 91.3 (See Table 3).
-While early rule-based systems like MYA using FlowXo, and ChatPal using Rasa (Boyd et al., 2022) reported lower scores when compared to more advanced NurtureBot versions using GPT-4 and DIAS using GPT-4 Turbo (Fersch et al., 2022), which achieve significantly higher scores.
-Notably, despite all NurtureBot versions using the same base technology, the iterative development of a human-centred interaction layer led to significant improvements in usability, demonstrating that thoughtful design can enhance user experience even without changes to the underlying technology.
-Table 3.
-In comparision with CUQ scores reported in literature NurtureBot v3 achieves the highest CUQ score of 91.3, followed by NurtureBot v2 and v1, demonstrating progressive improvements in usability, highlighting the effectiveness of iterative human-centred design in enhancing chatbot performance.
-Chatbot Name Base Technology Domain Participants CUQ Score
-MYA(Larbi et al., 2022) FlowXo (Rule-based) Physical Activity 30 57.4
-ChatPal(Boyd et al., 2022) Rasa Mental Health 10 65
-Chat Ella(Zhang and Song, 2024) GPT-2 Chronic Diseases 64 68.31
-GPT-Simulated Patient(Holderried et al., 2024) GPT-3.5 Medical Education 28 77
-DIAS(Fersch et al., 2022) GPT-4 Turbo Educational Services 103 80
-NurtureBot v1 GPT-4 Parental Wellbeing 32 85.4
-NurtureBot v2 GPT-4 Parental Wellbeing 32 87.2
-NurtureBot v3 GPT-4 Parental Wellbeing 64 91.3
-Based on our findings for co-designing user-LLM interactions, in Parts I-VI, we derive implications for design of AI-driven parental support systems presented in Table 4.
-Table 4. Implications for Design of AI-driven Parental Support Systems
-Implication Description
-Good Data Sources for AI A key foundation for successful AI is access to reliable and well-structured data sources.
-NurtureBot’s effectiveness depended on the quality and relevance of the information provided.
-Future iterations must use up-to-date, verified sources, including health organisation-endorsed health guidelines and localised resources, to enhance trust and provide contextually relevant information.
-User Understanding, Control, and Outcome Improvement Effective human-centred AI balances understanding, controlling, and improving interactions.
-Users should have transparency in their interactions with the AI and be able to guide the interaction through dynamic in-situ controls.
-Enabling users to refine the agent’s responses enhances the quality of interactions and makes the AI more responsive to user needs.
-Personalisation, Contextualisation, and Localisation Users expect personalised interactions, asking for the chatbot to remember personal details such as their names, their child’s name, where they live, and their previous conversations.
-These aspects help the chatbot tailor responses to user’s specific needs.
-Future designs should consider secure methods for personalisation without compromising privacy.
-Community and Peer Support Users desire more community-oriented support for parental wellbeing.
-This reflects the need for AI systems to connect users with localised communities, parenting groups, or forums.
-AI could act as a bridge between users and real-world community resources, providing emotional and practical support.
-Multi-modal Interactions Users expressed a need for multi-modal interactions such as video, voice, and image-based communication alongside text.
-This allows flexibility in how users interact with the chatbot, broadening accessibility for users with different preferences, literacy levels, or disabilities.
-Emotional Design Emotional connection emerged as a key element of success for NurtureBot.
-Many participants referred to the assistant as a ”friend,” indicating the importance of empathy and emotional support.
-However, it’s crucial to ensure users understand the AI’s role as a tool and avoid fostering over-reliance.
-Co-Design and Iterative Testing Co-design played a critical role in aligning NurtureBot’s features with user needs.
-Iterative testing allowed us to refine the system based on feedback.
-This iterative process of testing, redesigning, and retesting ensured that the final product provided the best possible user experience.
-9. Discussion
-The co-design of interactions with AI within the NurtureBot project highlights several key insights into how parental wellbeing support systems can be improved through iterative user feedback and involvement.
-Prior literature underscores the importance of human-centred AI design, particularly in sensitive areas like mental health and parenting, where emotional engagement and trust are crucial (Luchs, 2023; Shankar et al., 2024; Ferrara, 2022; Ke et al., 2024).
-Through iterative, parent feedback, our NurtureBot project explored the main interaction problems with LLM-based parental-wellbeing assistant when simulating support-seeking scenarios (RQ1), what their key needs and priorities would be (RQ2), and how we might co-design better human-AI interactions (RQ3).
-Our initial trial of Nurturebot v1 revealed seven key problems (RQ1), comprised of: lack of conversational flow; patronising tone; no personalisation; non-localised resources; generic responses; memory gaps; and transactional interactions.
-We note that these problems are prevailing across wellbeing initiatives and chat-based applications in the literature (Lee et al., 2019; Hwang et al., 2019; Viswanathan et al., 2020; Garcia et al., 2018).
-To begin to address these, with RQ2, we focused on prioritising key user needs, and their ideal vision for how NurtureBot should function.
-The focus was on explaining its workings, providing control over the chat, and opportunities to improve the conversation.
-Finally, we examined how we could employ co-design methods to craft interactions around AI (RQ3).
-The ARC co-design study identified that having a “successful chat” as the key user need, alongside ideal features such as using metaphors for understanding, beginning with empathy, clearly introducing features, providing step-by-step guidance, personalising and deepening sessions, offering clarification and reassurance, and summarising exercises with follow-up options.
-The detailed insights gained from addressing just one of the seven key issues identified in the NurtureBot v1 trial highlight a critical point: LLM responses are only as good as the user’s input.
-This reinforces that, despite meticulous design, the usability of even the most carefully crafted LLM-based support systems remains highly uncertain (Zamfirescu-Pereira et al., 2023).
-Continuing our efforts for answering RQ3, we further explored how we can use the insights from the ARC sessions to enhancing usability and the overall user experience.
-Incorporating participant feedback from ARC, in NurtureBot v2, we built an interaction layer designed to enhance user experience by focusing on three core principles: Understand, Control, and Improve.
-This layer afforded NurtureBot to better comprehend user needs, offer users control over the flow of conversation, and provide dynamic options for improvement through levels of interaction built in the system’s prompt.
-Across all of NurtureBot’s features (empathetic chatting, wellbeing exercises, and parenting information), we built interaction elements to offer more autonomy and flexibility in directing the interaction, when seeking support.
-We employed iterative co-design methods to leverage real-world end-user feedback in the agent’s prompt to better align it with user needs, a trend also emerging in the domain of Machine Learning (Shi et al., 2024; Hwang et al., 2023).
-This extends existing applications of methods that are typically used for improving LLMs, e.g.,fine-tuning, transfer learning, and reinforcement learning from human feedback (RLHF).
-Our approach builds on existing research through design methods in HCI (Benjamin et al., 2024) to “heard AI” to meet user needs (Zamfirescu-Pereira et al., 2023).
-Our findings suggested that by incorporating a user-co-designed interaction layer, employing the tripartite human-centred AI principles of understanding, control, and improving interactions, we were able to improve the usability and user experience in subsequent versions of NurtureBot.
-Our research adds to this body of work leveraging human-centred computing for AI alignment (Gilbert, 2024).
-By enabling users to steer the conversation with a clear understanding and responsive feedback loops, LLMs can become more intuitive and aligned to diverse domains and user needs.
-By demonstrating that user feedback can directly shape how AI-powered parenting supporting systems are perceived and used via ARC sessions, our research aligns with broader HCI research, which shows that empowering users to co-design results in more effective and engaging systems for designing both parental wellbeing (Prabhakar et al., 2017; Benjamin et al., 2024) and broader AI (Hossain and Ahmed, 2021; Delgado et al., 2023).
-10. Limitations and Future Work
-While the iterative development of NurtureBot improved its interaction capabilities, limitations persist, particularly in the areas of personalisation, contextualisation, and localisation.
-Addressing the diverse family roles, children’s characteristics, and unique circumstances of users is crucial to ensure the support provided is relevant and tailored to individual needs.
-Additionally, the ”time effect” underscores the importance of investigating how users emotionally engage with AI over extended use.
-While NurtureBot serves as a valuable tool, careful design is needed to avoid fostering over-reliance.
-Lastly, testing with slightly out-of-target groups to minimise emotional risks poses a limitation, as greater risks could arise when engaging prenatal parents in future tests.
-In future work, we plan to address these limitations by enhancing the situational features and designing for optimal reliance.
-Collaborating with public health organisations will be crucial to deploying and scaling NurtureBot on a broader platform, ensuring it reaches a wider audience of parents.
-11. Conclusion
-The significance of comprehensive perinatal care cannot be overstated, as it plays a pivotal role in mitigating risks associated with perinatal mind-body duality issues, fostering a positive environment that promotes healthy development of the parent(s), the child, and the long-term stability of the family unit.
-We explored iterative co-design of AI-powered parental wellbeing support with NurtureBot, engaging users through ARC sessions, where they were not only able to prioritise key issues but also contribute to the redesign of user-LLM interactions, for improved user comprehension, control, and task outcomes.
-The resulting interaction layer significantly improved the user experience across versions, suggesting that AI-driven systems can better align with real-world needs by empowering users to shape their interactions, making AI tools more effective and widely adopted in wellbeing and personalised support contexts.
-By prioritising the human experience, AI tools can not only enhance wellbeing but also foster stronger, more resilient communities and create scalable, long-lasting solutions to meet diverse real-world needs.